### v1

# From attention to profit: quantitative trading strategy based on transformer

Abstract
In traditional quantitative trading practice, navigating the complicated and dynamic financial market presents a persistent challenge. Former machine learning approaches have struggled to fully capture various market variables, often ignore long-term information and fail to catch up with essential signals that may lead the profit. This paper introduces an enhanced transformer architecture and designs a novel factor based on the model. By transfer learning from sentiment analysis, the proposed model not only exploits its original inherent advantages in capturing long-range dependencies and modelling complex data relationships but is also able to solve tasks with numerical inputs and accurately forecast future returns over a period. This work collects more than 5,000,000 rolling data of 4,601 stocks in the Chinese capital market from 2010 to 2019. The results of this study demonstrated the model’s superior performance in predicting stock trends compared with other 100 factor-based quantitative strategies with lower turnover rates and a more robust half-life period. Notably, the model’s innovative use transformer to establish factors, in conjunction with market sentiment information, has been shown to enhance the accuracy of trading signals significantly, thereby offering promising implications for the future of quantitative trading strategies.
keywords:
Quantitative finance, Transformer, Stock selection, Market sentiment
1 Introduction
The goal of stock trading is to optimize the return on the investment in the capital market according to the process of buying or selling one or more companies’ shares. Traders obtain profit when the difference is generated by the fluctuation of stock spice. However, stocks are influenced by a large number of factors, which constitute a complex system and increase the difficulty for people to make a profit. The assessment of a stock’s evolving trend is inherently challenging due to the highly volatile and interconnected nature of the market, which sets it apart from typical time series modeling [1]. So, many strategies and tools were built, which corresponded to the development of the capital market and quantitative strategies have played important roles among them.
Figure 1: The Quantitative Trading Architecture with Factors
Some traditional quantitative tools, such as Markowitz Portfolio [2] and Capital Asset Pricing Model(CAPM) [3], focuses mainly on the fundamental analysis. In other words, these strategies aim to make a profit by simple calculation and analysis. However, depending on the development of computer science, more quantitative methods and tools were introduced. Within these methods, strategy with factors has attracted people’s attention. In 1993, Fama and French [4] introduced their Fama French Three Factor Model (FF3), which is one of the most influential models in quantitative trading. In 2015, Fama and French renewed their model, the Five-Factor Asset Pricing Model (FF5) [5] was introduced. Besides this classical theory, other trading strategies have been published for decades.
Quantitative trading with factors typically manifests in two primary strategies, which are shown in Fg.1. The first approach involves the computation of stock factor values. Based on these calculated values, stocks are ranked to establish a pool. Once this pool is established, assets are held for a predetermined period. Adjustments to the portfolio are then made at specific time intervals, ensuring alignment with evolving market conditions and factor readings. The second method employs a fixed pool of stocks, wherein factors guide the derivation of long/short signals. Traders can execute corresponding actions when they receive the signals from factors, allowing for a dynamic response to market fluctuations based on factor insights.
In recent years, Machine Learning (ML) stands out as an instrumental tool in enhancing trading algorithms and decision-making processes. Its core principle, enabling systems to learn from and make decisions based on data, lends itself particularly well to the vast and dynamic landscapes of stock markets. Some traditional literature exploits machine learning algorithms based on manual indicators [6], but the hypothesis about the random walk may be less useful in the task of catching non-stationary series. With the rise of different architectures, quantitative researchers found some novel approaches to identifying intricate patterns and factors in trading data, which opened a new field in quantitative trading [7].
Although there are several previous experiments about ML factors in the quantitative finance field, the research still faces two concerns. Firstly, in the field of NLP, it is necessary to convert words in text to word vectors through word embeddings to serve as inputs. However, in the financial dataset, there are categorical data such as industry types, as well as numerical data like price fluctuation, turnover rate, and financial indicators. If the input comprises only categorical data, the time series can be treated as a sentence [8]. In most cases, the input will involve numerical data, which cannot be transformed via word embeddings.
For the second concern, in NLP tasks, most of them can be transformed into sequence-to-sequence (seq2seq) problems, such as in machine translation, dialogue systems, and speech recognition. To utilize existing outputs, decoders in NLP tasks sequentially output samples and use masking operations to handle input sequences during training. However, in stock prediction, where the aim is often to accurately forecast future returns over a period, the transformer model is rarely used for such tasks.
To address these concerns, we propose an approach that combines time series stock data with a numerical-adapted transformer to establish a factor and put it into back-tests. This approach allows us to use transformer model embedding numerical input directly, which refers to a method similar to sentiment analysis. The structure of the paper is designed as follows. Section 2 discussed the previous quantitative financial works mainly based on machine learning. Section 3 and 4 introduce the establishment of the model. A factor based on the model will be trained and back-tested. For the practical test, we collected data from more than 4,600 stocks in the previous 14 years (from 2010 to 2023) from the data-collect platform. To comprehensively test the ability of the factor, we divided of data by different frequencies (monthly, weekly, and daily). Finally, the result of the back-tests including the comparison between the transformer-factor and others and the insights gained from such comparative analysis will be discussed in the section 4 and 5.
Figure 2: Example for stock prediction
2 Preliminaries
This section briefly introduces the related works about the paper and the development of quantitative financial trading methods, especially based on Machine Learning (ML) methods.
2.1 Stock data with market sentiment
Stock data can be considered that can influence the future trend of this stock in some cases [9]. Previous works have discussed the correlation between investor sentiment and stock market metrics like turnover rate, indicating that investor feelings, both individual and institutional, significantly influence market behavior, affecting aspects such as volatility and returns [10, 11]. This relationship isn’t uniform but varies with changes in investor sentiment. Research also highlights that investor psychology, especially during periods of high market stress can drastically impact market behavior and investor decisions. This link between sentiment and market performance is further affirmed by studies focusing on factors like the turnover ratio, which are seen as reflective of market liquidity and investor behavior [12]. These findings underscore the mutual relationship between investor sentiment and market performance, demonstrating how psychological factors can drive market dynamics.
2.2 Applications for ML methods in trading
Machine learning exploits a range of algorithms and statistical techniques used for tasks such as regression, clustering, and classification. In past decades, there have been methods that have been significantly increasingly applied in stock due to their ability to process vast amounts of data and make predictions, offering a potential advantage in the financial markets. At the same time, the finance sector, particularly quantitative trading, has started to be aware of the potential of deep learning models for predicting stock movements, portfolio optimization, and risks.
2.2.1 Support Vector Machines (SVM)
SVM was widely used in quantitative trading, which can manage high-dimensional spaces and intricacies found in datasets. In 2003, SVM tried to forecast financial time series [13], especially with the lens focused on the South Korean stock market. This work exhibited the robustness of SVM against traditional statistical methods and some other computational techniques, setting a precedent for its adoption in financial forecasting. Similar to the previous one, [14] showcased SVM’s prowess in anticipating the movement direction of Japanese stock markets, especially when determining bullish or bearish market sentiments. [15], found SVM’s adaptability and precision in financial tasks like portfolio and quantitative trading.
However, SVM also shows some disadvantages in quantitative trading. In 2009, SVM was underscored for the accuracy of overfitting [16]. When applied to financial datasets, the model’s propensity to fit too closely to the training data might lead to reduced generalization capabilities, causing the strategy ineffective on unseen data. In high-frequency trading, though the model displayed competency in specific scenarios, researchers pinpointed potential limitations when grappling with certain volatile market dynamics, suggesting the need for adaptive [17].
2.2.2 Long Short-Term Memory (LSTM)
LSTM was designed to capture long-term dependencies in sequential data according to their unique architecture [18]. [19] deployed LSTM with stacked autoencoders to forecast financial time series. LSTM shows the ability to accentuate and effectively model volatile financial markets, showcasing significant improvements over traditional time series models. In 2018, [20] tested the efficacy of LSTM for price predictions. They observed LSTMs’ prowess in noisy data and their superiority in prediction accuracy over some conventional models. LSTM’s capability to extract features from time series was explored by [21]. They combined LSTM with other architectures and found that these can enhance the feature extraction process, thereby improving prediction accuracy. For predicting stock prices, [22] harnessed the power of LSTM to unearth multi-frequency trading patterns. Their research underlined the adaptability of LSTMs in modeling complex situations inherent in stock prices.
2.2.3 GRU
Gated Recurrent Units (GRU) [23] have emerged as a useful method in the quest for accurate financial forecasting. GRU can capture long-term dependencies in time series data and has seen increasing popularity in the world of quantitative trading. [19] introduced a comprehensive deep learning framework for financial time series, which Incorporated GRUs alongside LSTM and stacked autoencoders. Their model, when applied to diverse datasets, performs better than traditional time series models, accentuating the adaptability and robustness of GRUs. [24] centralized on LSTM networks, provided comparative insights into GRUs. However, in their analysis, LSTMs showed a slight edge in prediction accuracy, indicating that while GRUs are powerful, selecting between them and LSTMs may boil down to specific use cases and computational constraints. What’s more, [25] presented a model fusing GRUs with Convolutional Neural Networks (CNN) [26]. The strategy combined CNNs and GRUs, also received advantages among them, creating a formidable solution for stock price prediction.
2.3 Transformer in quantitative trading
2.3.1 Introduction for transformer
Transformer [27], which was introduced in 2017, has received high attention in the NLP field. According to its special encoder and decoder stacks with self-attention blocks, the transformer shows its advantage in stability, speed, and long-term memory compared with other traditional models. The transformer-based models like GPT [28] and BERT [29] are the basis of the specific field of application. In the application field, transformer-based models like ChatGPT [30], FinBERT [31, 32] and BloomberGPT [33] show advantages and ability in solving different types of problems in different filed.
Similar to other neutral sequence transduction models, transformer also has an encoder-decoder structure. The model takes input as a sequence of context which could be natural language sentences and generates output with texts depending on the input, like a translation according to the input or a predicted sentence. Transformer using stacks which are called self-attention and point-wise, which connect layers in the architecture for encoder and decoder. For these encoder and decoder layers, the self-attention method plays an important role in matrix calculation. The linear and softmax functions are used in the dot-product attention module. In the vanilla transformer model, there are multi-head attention layers are used and the result from each layer will be outputted after going through an additional &\&& normalize layer. The above-mentioned steps covered most of the progress in the vanilla transformer.
2.3.2 Previous transformer-based strategy
Transformer has achieved superior performance on long-term dependencies modeling. In the quantitative trading domain, some previous researchers tried to test the potential of transformer. [34] firstly tried to exploit Transformer on trading sequences to classify stock price movements.
Besides using a traditional transformer, other innovative methods of transformer are worth mentioning. [35] devoted to optimizing the efficiency of time complexity and memory usage of Transformer on extremely long time series by informer. [36] leveraged the advantages of CNNs and Transformers to model short-term and long-term dependencies in financial time series. They showed the advantage of the proposed approach on intraday stock price prediction of S&P 500 constituents.
Similar to previous work, [1] innovated the structure of a Transformer, which is Adaptive Long-Short Pattern Transformer (ALSP-TF). Their model is structurally innovated for hierarchical representation and interaction of stock price series at different context scales. With the help of a learnable function, they make the self-attention aware of the weighted time intervals between patterns, to adaptively adjust their dependencies beyond similarity matching. Finally, they obtained more than 10% of annual return on average.
However, there are some disadvantages of transformer. [37] mentioned that the global self-attention module focuses on point-wise token similarities without contextual insights. As fluctuations of stocks are conditioned on composite signals over manifold periods, lacking pattern-wise interaction hinders the adequate discrimination of stock tendency and is susceptible to noise points. On the other hand, [38] claimed that the basic query-key matching paradigm is position agnostic. Although position embedding is inserted into the sequential inputs, it may not be optimal because of the inability to reveal precise distances.
3 Methodology
This section introduces the framework of the work and the steps of the establishment of the model, including the data processing, model enhancement and output processing.
3.1 Framework overview
Fig. 3 illustrates the overview of the structure. There are three major parts:
(i) Data Initialization to align the time series of stocks into a regular matrix;
(ii) Embedding the training dataset and training them by transformer which replaces the word-embedding layer to linear layer and deletes the Mask(opt.) layer;
(iii) Lastly, using the trained model to predict the possible trend of stocks.
Figure 3: Overview of the work
3.2 Problem formulation
Inputs
Given the candidate stock set $S^{t}$ for N stocks on the trading timestamp $t$:
$S^{t}=\{s_{1}^{t},s_{2}^{t}\ldots s_{N}^{t}\}$
For a stock $s_{k}^{t}$ in stock set $S^{t}$, where $k\in N$, it details a sequence for a historical 1-time step (such as one month), and the dimension is 2:
${\chi}^{t}_{k}=[x_{k}^{t-20},x_{k}^{t-19}\ldots x_{k}^{t}]\in\mathbb{R}^{20\times 2}$
In each $x_{k}^{t}$, there are two features. The first one is the accumulated daily profit rate during the time step, where $p_{i}$ and $p_{i-1}$ is the close price of the $i^{th}$ and ${i-1}^{th}$ day, respectively.
$r^{t}=\sum r_{i}=\sum\frac{p_{i}-p_{i-1}}{p_{i-1}},\quad\sigma^{t}=\sum\sigma_{i}$ (1)
Another is the accumulated daily turnover rate $\sigma^{t}$ during the time step and the calculation method of $\sigma^{t}$ is shown above. So each $\chi^{t}_{k}$ can be represented as:
$\chi^{t}_{k}=\left[\begin{matrix}r^{t-20}&r^{t-19}&\ldots&r^{t}\\ \sigma^{t-20}&\sigma^{t-19}&\ldots&\sigma^{t}\\ \end{matrix}\right]\in\mathbb{R}^{20\times 2}$
Outputs
For a stock set $S^{t}$ on the trading timestamp $t$, each stock $S^{t}_{i}$ has a profit rate $r_{i}^{t+1}$ on the trading timestamp $t+1$, where the calculation method is same as the method of $r^{t}$ in the sequence of inputs shown in Eq.(1). Then the value in the list of next-time stamp’s profit for $N$ stocks in the timestamps $t+1$ can be represented as:
$P^{t+1}=\{r_{1}^{t+1},r_{2}^{t+1}\ldots r_{N}^{t+1}\}$
And the list will be arranged as $n$ parts and each part has an equal proportion $\frac{1}{n}$ ($n$ should be an odd number, such as 3,5,7). Take $n=5$ as an example. For those ranked values in the first, middle and last $\frac{1}{5}=20\%$, the output of the stock $s_{k}$ records as below, respectively:
$y^{t}_{k}=[1,0,0]^{T},\quad y^{t}_{k}=[0,1,0]^{T},\quad y^{t}_{k}=[0,0,1]^{T}$
3.3 Embedding initialization
Normalization
The next step is the normalize the input data for each time step by zero-mean normalization (Z-Score), the normalization equation is shown in Eq.(2).
$\tilde{x}^{t}_{k,i}$ represents the normalized value at $i^{th}$ day of stock $k$ in time $t$.
$E[x^{t}_{i}]$ and $std[x^{t}_{i}]$ is the mean and standard of all the value of $x^{t}_{i}$, respectively.
$\tilde{x}^{t}_{k,i}=\frac{x^{t}_{k,i}-E[x^{t}_{i}]}{std[x^{t}_{i}]}$ (2)
In this way, the input consequences towards zero mean and unit variance, which aims to reduce the influence from different time points and allow different features to be comparable with each other [39].
3.4 Transformer encoder
Then a transformer encoder structure will be built in this subsection. The representation hierarchy consists of $L$ blocks of multi-head self-attention layers. Taken initialized stock embedding sequences $\textbf{P}\in\mathbb{R}^{N\times 20\times 2}$ as input, and the canonical self-attention in [27] performs information exchange between every time points for each stock $s_{i}$.
To process both categorical and numerical data, the word embedding layer is replaced by a standard linear layer, utilizing linear transformations to substitute the process of word embedding. On the other side, in stock prediction, we aim to accurately forecast the return for a future period, thus the model’s output is generally a single value or the probability of price increase or decrease. Therefore, the decoder is simplified by removing the autoregressive prediction mechanism (in the decoder) and the masking operations.
For each $\textbf{P}_{i}\in\mathbb{R}^{20\times 2}$, transforms into query, key, and value metrics, respectively:
$\textbf{Q}_{i,h}=\textbf{E}_{i}\textbf{W}^{Q}_{h},\quad\textbf{K}_{i,h}=\textbf{E}_{i}\textbf{W}^{K}_{h},\quad\textbf{V}_{i,h}=\textbf{E}_{i}\textbf{W}^{V}_{h}$ (3)
where $h=1,\ldots,H$ is the head index, as well as $\textbf{W}^{Q}_{h}$, $\textbf{W}^{Q}_{h}$, $\textbf{W}^{Q}_{h}$ $\in\mathbb{R}^{2\times d_{f}}$. Lastly, the final layer is represented by the concatenation of all attention heads.
$\textbf{F}_{i}=Multihead(\textbf{Q}_{i,h},\textbf{K}_{i,h},\textbf{V}_{i,h})=||^{h=H}_{h=1}Softmax(\textbf{Q}_{i,h},\textbf{K}_{i,h}^{T}/\sqrt{d_{f}})\textbf{V}_{i,h})$ (4)
Just like Eq.(4) shows, ||||| | represents the concatenation operator, and $d_{f}$ is the dimension of projected feature space. In the end, all input consequences are represented in the form of:
$\textbf{F}=[\textbf{F}_{1};\textbf{F}_{2};\ldots\textbf{F}_{N}]\in\mathbb{R}^{N\times 20\times Hd_{f}}$
which is followed by feed-forward layers.
According to Sec.3.2, the embedding output is denoted as:
$\textbf{Y}^{t}=\{y^{t}_{1},y^{t}_{2},\ldots y^{t}_{N}\}\in\mathbb{R}^{N\times 1\times n}$
3.5 Output process
Eventually, the output is represented:
$\textbf{Z}^{t}=\{z^{t}_{1},z^{t}_{2},\ldots z^{t}_{N}\}\in\mathbb{R}^{N\times 1\times n}$ (5)
The output will be activated by the $softmax$ function. And for each $z^{t}_{i}$, there are:
$z^{t}_{i}=[z^{t}_{i,1},\ldots z^{t}_{i,m},\ldots z^{t}_{i,n}]\in\mathbb{R}^{1\times n}$
the output of probability can be represented as (m is the middle number from 1 to n):
$Y^{t}_{o}=[y^{t}_{o,1},\ldots y^{t}_{o,m},\ldots y^{t}_{o,n}]^{T}=[\frac{e^{z_{1}}}{\sum_{i=1}^{n}e^{z_{i}}},\dots\frac{e^{z_{m}}}{\sum_{i=1}^{n}e^{z_{i}}},\ldots\frac{e^{z_{n}}}{\sum_{i=1}e^{z_{i}}}]$ (6)
In Eq.(6), $y^{t}_{o,i}$ are larger than 0 and smaller than 1 as well as the sum of all the items in $Y^{t}_{o}$ is 1. So the output means the probability of the stock’s raising, flatting, or downing in the next timestamp.
3.6 Prediction
The Mean Squared Error Loss (MSELoss) will be used to quantify the loss. MSE Loss is a widely used metric for assessing the discrepancy between a model’s predictions and the actual values. It calculates the average of the squares of the differences between the predicted value and the actual prices. This is formally represented as:
$MSELoss=\frac{1}{n}\sum^{n}_{i=1}(y_{i}-\hat{y}_{i})^{2}$ (7)
In Eq.(7), $y_{i}$ is the actual value of the $i^{th}$ equation, $\hat{y}_{i}$ is the model’s prediction for the same, and n represents the total number of samples in the inputs. During each epoch, the MSE Loss is computed to guide the optimization process, to minimize this value over successive iterations. A lower MSE indicates a closer alignment between the model’s predictions and the actual stock prices, signifying an improvement in the model’s learning and its ability to generalize from the training data. This loss function is particularly useful in regression tasks, as it not only penalizes the model for inaccurate predictions but also emphasizes reducing larger errors, thereby ensuring a robust learning process.
4 Experiments
Based on the methodology in Section 3, in this section, the set of experiments is introduced, including the data resource, implementation details, trading strategy and metrics.
4.1 Dataset
The training data of the transformer factor comes from the Chinese exchange market. The dataset is collected from the Shanghai Stock Exchange (SHSE) and Shenzhen Stock Exchange (SZSE), which contains 4601 stocks that are listed or had been listed from January 2010 to May 2023. And all these data come from AKShare [^1] https://akshare.akfamily.xyz/ and Tushare [^2] https://www.tushare.pro/, which is one of the quantitative finance terminals. The train period is from January 2010 to December 2019 and the test period starts from January 2020.
Adjustment closing price is used for the stock training price [40]. Adjustment closing price is essential in stock market analysis, particularly when analyzing historical data for long-term trends and patterns. This price takes into account factors like dividends, stock splits, and other corporate actions that can affect a stock’s price over time [41]. By adjusting for these events, the adjusted close price provides a more accurate picture of a stock’s value and performance, which is more persuadable in quantitative financial back-test.
4.2 Timestamp
To test the performance of the factor better, three kinds of frequency are chosen, which are stock data monthly, weekly, and daily, respectively.
In the first experiments, a timestamp was set as one month. So for a stock $s_{i}$ from stock set $S^{t}$, it details a feature sequence in which each item in the sequence contains the accumulated profit and accumulated turnover rate in a month. Both the accumulated profit and accumulated turnover rate are the sum of the trading day’s data in that month. In some cases, if one stock was first listed on the market or resumed trading in one month, the available data may not cover whole trading days, the data of the stock from that from would still be recorded. However, if there is a whole month that the stock did not have the trading record (perhaps stopped trading or not listed in that month), this month’s data will be recorded as ”NaN”. In the sequence $\chi^{t}_{k}$, the sequence will not be used to train the model when there are missing values in the sequence.
Within the first experiments, three sub-experiments were set. For the sub-experiments, the input of them are same. For output data, outputs with different dimensions and different lengths but the same dimension are used. $n=5$ and $n=7$ are used in this experiments. In the group of $n=5$, the first one deleted the data with output like $[0,0,0]^{T}$, and the second one used the origin dataset (zero-output).
Most of the accumulated profit values in different years are in the range $(-10\%,10\%)$. There are some values point outside the range $(-100\%,100\%)$ that are shown in it because these values are too large (small), such as the stock ”002961.SZ” first listed at SZSE on September $4^{th}$, 2019 and its accumulated profit rate on that month is 184.66%. These stocks received unusual performance mainly because they were first listed on the market, resumed trading, or reported great positive/negative news.
Most of the accumulated turnover rate values are in the range $(0,100\%)$. There are some values outside the range $(0,200\%)$ that are shown in it because these values are too large, such as the accumulated turnover rate in July 2015 of ”002596.SZ” is $777.67\%$, this was because the actual controller abundantly increased his holdings of the company and the company announced that they received findings by shareholders. Similar to the example, the situation may be caused by great positive or negative information.
Outliers of both accumulated profit and accumulated turnover rate will not be removed from the dataset these situations may happen in the future and they are expected to be predicted, though these incidents rarely happen.
In the second and third experiments, the inputs are in weekly and daily frequency, respectively. The sub-experiments under different groups are similar to the first one. All the accumulated number of parameters, the number of trained sections (such as 100 months from 2010 to 2019), and dimension are shown in the Table 1.
Table 1: Detailed information of experiments
Strategy Frequency Output Dim Training samples Section zero-output
Month_1 Monthly 3 85,490 100 w/o
Month_2 Monthly 3 142,409 100 w/
Month_3 Monthly 5 142,409 100 w/
Week_1 Weekly 3 455,157 466 w/o
Week_2 Weekly 3 758,300 466 w/
Week_3 Weekly 5 758,300 466 w/
Day_1 Daily 3 3,586,435 2,420 w/o
Day_2 Daily 3 5,140,279 2,420 w/
Day_3 Daily 5 5,140,279 2,420 w/
4.3 Implementation Details
The training process and model are implemented with PyTorch [42]. The parameter of the model is searched by gird search, which is reliable in low dimensional spaces and simple to implement [43]. As a result, in the situation of input dimension is 2, the dimension of hidden feature space $d_{f}$ is 16. The number of multi-head attention modules is 16, and the number of layers of encoder and decoder is 6. The model was trained on an NVIDIA GeForce RTX 2080 GPU and NVIDIA A100 Tensor Core GPU by Adam optimizer [44] for 50 epochs, the learning rate is 0.001, and the batch size is 64.
4.4 Trading strategy
A trading strategy is used for the model and Algorithm (1) shows the algorithm of strategy. Before the first trade date of the timestamp $t$, all sequence $\chi^{t}_{i,k}$ from stock set $S^{t}$ will be put in the model and obtain the list of the output. Then the stocks will be ranked according to the first element of the output and the first $\frac{1}{n}\%$ stocks will be added to the stock pool. If the stock has already been in the stock pool on the last timestamp, it will keep holding; if the stock is in the predicted pool but not in the previous pool, it will be bought in with the same proportion of the whole account. Stocks that are not in the predicted pool will be sold out. The same method will run repeated during the rest period.
The backtest starts from January 2020, in other words, the result of the sequences from May 2018 to December 2019 will be used as the first stock pool to trade.
Algorithm 1 Trading Strategy
1:The sequence $x^{t}_{i,k}$ for each stock $s_{i}$ in the set $S^{t}$ and the amount of cash.
2:run_monthly(trade, monthday=1, time=‘open’)
3:function obtain_stockpool($x^{t}_{i,k}$):
4:     stockpool $=$ sort(model($x^{t}_{k}$), key=lambda x : x[1], reverse=True)
5:     return stockpool
6:end function
7:function trade
8:     stockpool $=$ obtain_stockpool($x^{t}_{i,k}$)
9:     for stock in previous_stockpool do
10:         if stock not in stockpool then
11:              order(stock, 0) $\triangleright$ sellout
12:         end if
13:     end for
14:     for stock in stockpool do
15:         if stock not in previous_stockpool then
16:              order(stock, amount / len(stockpool)) $\triangleright$ buy
17:         end if
18:     end for
19:end function
4.5 Metrics
Formally, the Sharpe Ratio (SR) and the $\alpha$ ratio will be used to test the performance of the strategy. The SR [3] is a measure of risk-adjusted return which can describe the additional earnings an investor receives per unit of increase, which is shown in Eq.(8), where $R_{p}$ is the return of the portfolio, $R_{f}$ is the risk-free rate:
$SR=\frac{E[R_{p}]-R_{f}}{std[R_{p}]}$ (8)
Alpha represents the excess return of an investment portfolio relative to its benchmark. It reflects the stock selection skills of an investment. A positive alpha indicates that the investment portfolio has achieved a higher return than its benchmark after risk adjustment. Specifically, alpha is the excess return of the actual portfolio return over its expected theoretical return. Eq.(9) shows the calculation method of alpha rate, where $E(R_{m})$ is the market return and $\beta$ is the correlation between the portfolio and systemic risk.
$Alpha=(R_{p}-r_{f})-\beta(E[R_{m}]-r_{f})$ (9)
Besides the SR ratio and alpha rate, the annual return (AR) of the stock, the annual excess return (AER), the average turnover rate (TR) of the portfolio, and the win rate (WR) will also shown. The excess returns are returns achieved above and beyond the return of a proxy and the CSI 300 index is used as the basis when calculating the excess returns. The turnover rate is the average percentage of the portfolio adjusted daily in terms of market. The win rate is the percentage of trade days where the portfolio encounters a positive return overall trade days.
5 Result and discussion
Building upon the analysis of individual strategies, this section presents an evaluation and discussion of the experimental outcomes. The discussion focuses on the comparative effectiveness of monthly, weekly, and daily strategies. The results underscore the nuanced relationship between trading frequency and market performance, offering valuable insights for the development of robust quantitative trading strategies.
5.1 Overall performance
With the starting amount of RMB 5,000,000, the results of different strategies are obtained. The result of the experiment is shown in Table 2. For the monthly group, Month_1 shows an annual return of 17.35% and an annual excess return of 19.43%, indicating a robust performance over that period. Conversely, Month_3 shows a diminished annual return and annual excess return of 7.37% and 9.91% respectively. The turnover rate is notably high in the second month at 51.69%. The win rate ranges from 43.1% to 57.3%, with the first month showing the highest win rate. The Sharpe Ratio and Alpha, vary across the strategies, with some periods showing negative values, suggesting underperformance relative to the risk taken, while others show positive values.
Table 2: Result of experiments
Strategy AR AER TR WR SR Alpha
Month_1 17.35% 19.43% 26.09% 57.8% 0.915 0.162
Month_2 9.91% 13.86% 51.69% 49.3% 0.289 0.102
Month_3 7.37% 9.91% 32.33% 51.6% 0.246 0.064
Week_1 -0.83% 1.31% 7.13% 46.4% -0.236 -0.03
Week_2 7.49% 10.81% 1.18% 49.2% 0.160 0.085
Week_3 12.3% 12.73% 1.39% 54.4% 0.372 0.116
Day_1 7.89% 11.4% 6.71% 43.1% 0.181 0.090
Day_2 10.23% 10.94% 6.51% 44.4% 0.279 0.097
Day_3 9.81% 10.03% 5.57% 44.6% 0.281 0.092
CSI300 1.77% \ \ \ \ \
In the weekly strategy results, Week_3 showed a remarkable annual return of 12.3% and an annual excess return of 12.73%, while Week_1 displayed a negative annual return. For the daily strategies, Day_2 had a return of 10.23% and an annual excess return of 10.94%. The turnover rate across the daily strategies showed lesser variation than the monthly or weekly, indicating a more consistent trading frequency. The win rate for the daily strategies remained fairly stable, hovering around the mid-40% mark. The Sharpe Ratio and Alpha for weekly and daily strategies show a mix of positive and negative values, reflecting the fluctuating nature of shorter-term trading efficacy.
To demonstrate the advantages of the Transformer factor over traditional factors, 100 price-volume type factors from JoinQuant  [^3] https://www.joinquant.com/ (one of Chinese largest financial quantitative platforms) were selected for backtesting under the same trading strategy (Algorithm 1). Price-volume factors were used because these are calculated based on stocks’ prices, volumes, and turnover rates, which are similar to the training data used by Transformer factors (also used stock turnover rates and price data). The detailed performance of factors and benchmark is shown in Appendix A. On average, these factors achieved an annual return of -3.78%, an average excess return of -2.15%, and a Sharpe ratio of -0.36, where the benchmark return is -1.77%. The average maximum drawdown of the strategy was 44.88%, which means the potential worst-case scenario or the most extreme possible loss. The average volatility of the strategy ($\sigma_{p}$) was 0.26, with the benchmark at 0.197.
Figure 4: Backtest results based on different factors
Among these factor-based strategies, the Transformer factor performed well. Except for the win rate, which was slightly lower than a few strategies, it excelled in terms of annual return, annual excess return, $\sigma_{p}$, and Sharpe Ratio, ranking the best among the 101 factors. Figure 4 illustrates the return curves of the Transformer (Trans_Month_1) alongside the benchmark and some traditional factors. ”25_week_rank” represents the current price’s position over the past 25 weeks; ”EMAC20” and ”MAC20” are the 20-day index moving average and the stock’s 20-day moving average respectively; ”DAVOL20” and ”VOL20” represent the ratio of the 20-day average turnover rate to the 120-day average turnover rate, and the mean of the stock’s 20-day turnover rate; ”ROC20” is the price rate of change over 20 days; ”MAWVAD” is calculated as the product of the difference between the closing price and the opening price, divided by the range of the highest and lowest prices, all multiplied by the volume, accumulated over six days; ”Variance20” is the variance of the stock’s 20-day annualized returns. Most of these factors are compared based on a window of 20 timestamps, as the Transformer factor was also trained on 20 timestamps. In comparison with these factors, the Transformer factor (blue line) demonstrated a significantly better performance in terms of returns than the benchmark (orange line) and other factor strategies, expressing the improvement of the Transformer factor over traditional price-volume factors.
5.2 Half-life period test
To test the half-life period of the factor, another three portfolios based on $Month\_3$ were conducted. The first one is the stock pool used in the next month; the second is using it in the month after the month; the last one is using it in the third month after the last feature of inputs. For instance, in the first strategy, the trade will start in January 2020; in the second, it will start in February 2020; in the second, it will start in March 2020. For the period from January 2020 to February 2022, the strategy achieved a total return of 65.29% and an annual return of 17.87%. The excess return over this period was 66.39%, indicating a strong performance above a benchmark or risk-free rate. The Sharpe Ratio for this period was 0.863, which is a respectable figure that suggests a favorable risk-adjusted return. Additionally, the strategy generated an alpha of 0.164, demonstrating its ability to outperform the expected return on a risk-adjusted basis.
Table 3: Result of the same monthly strategy with different start month
Strategy Total return Annual return Excess return Sharpe ratio Alpha
2020/01-2022/02 65.29% 17.87% 66.39% 0.863 0.164
2020/02-2022/03 67.52% 18.21% 65.58% 0.940 0.161
2020/03-2022/04 56.41% 15.63% 52.96% 0.739 0.135
In the subsequent period from February 2020 to March 2023, the strategy yielded a slightly higher total return of 67.52% and an annual return of 18.21%. The excess return was 65.58%, with an improved Sharpe Ratio of 0.940, reflecting a better risk-adjusted return compared to the previous period. The alpha for this interval was 0.161, indicating the strategy’s consistent performance in delivering additional returns.
Lastly, Table 3 shows the result of the backtest. From March 2020 to April 2023, the strategy’s total return decreased to 56.41%, with an annual return of 15.63% and an excess return of 52.96%. The Sharpe Ratio also declined to 0.739, but it still represents a relatively efficient investment compared to the volatility of returns. The alpha reduced to 0.135, which, while lower than the other periods, still indicates that the strategy managed to provide returns above those predicted by its risk profile.
5.3 Discussion
Comparing the results between monthly, weekly, and daily strategies provides further insights. The monthly strategies can deliver the highest returns, suggesting that longer-term signals better capture short-term trends. The weekly strategies underperform monthly ones but outperform the daily frequency overall. This indicates weekly data strikes a balance between filtering noise and responding swiftly to emerging patterns.
The daily frequency strategies perform lower annual returns compared to others. This highlights the challenge of making profitable trades at higher frequencies - noise and volatility make reliable signals more difficult. The distribution of returns also shows increased dispersion at the daily level. However, daily strategies achieve more consistent turnover rates around 5-7%. This allows for portfolio adjustments while avoiding excessive trading costs. In contrast, monthly strategies see a turnover rate above 50% in some cases. There is likely an optimal point between 1 week to 1 month where matches market regularly.
6 Conclusion
This study proposed an approach to exploit transformer models for quantitative stock prediction and trading. We addressed the need for handling numerical input data rather than text and adapting the model for forecasting tasks rather than sequence-to-sequence problems common in NLP. To enable direct processing of numerical time series data, we replaced the word embedding layer with a standard linear layer and removed the output masking operations. We also simplified the decoder to produce a probability distribution over future price movements rather than autoregressively generating token sequences.
Eventually, we tested the factor in the reality back test of the financial market and compared the factor with other 100 price-volume factors. Our experimental results demonstrate the promise of this approach. The model-based trading strategies were able to deliver substantial excess returns over the benchmark under a longer time period, with Sharpe Ratios indicating sound risk-adjusted performance. The positive alphas affirm the strategy’s ability to outperform expected returns after accounting for risk factors.
There is further scope to enhance the framework, for instance by incorporating additional signals like news and fundamentals as auxiliary inputs. The self-attention mechanism could likely be improved to encode greater temporal context. Or use transformer-based models such as GPT-4 [45] and Claude3 [46] to establish the strategy as these models are fine-tuned and may perform better in quantitative finance tasks.
Overall, our work illustrates the viability of transformers for financial data modeling, provides a flexible framework to model the market from multiple perspectives of financial time series and develops profitable trading strategies that consider different variables. With further research to solve the above limitations, we believe such ML-based quantitative methods hold rich potential. The framework provides a flexible foundation for better understanding markets and developing profitable trading strategies.
7 Acknowledgements and declaration
The authors have no competing interests to declare that are relevant to the content of this article.
Appendix A Detailed backtest results between 100 factors and the transformer factor
All the data of factors comes from JoinQuant [^4] https://www.joinquant.com/, which is one of the most common quantitative finance platforms. Detailed descriptions of the following factors are at JoinQuant Factors.
Seven metrics are used to evaluate the performance of each factor:
  • [label=•]
  • Annual Return (AR) and Annual Excess Return (AER) are two metrics to reflect the return of the portfolio directly. The excess returns are returns achieved above and beyond the return of a proxy and the CSI300 index is used as the basis when calculating the excess returns.
  • Win Rate (WR) is the metric to evaluate the fraction between the action that makes a profit and all the actions.
  • Sharpe Ratio (SR) is a measure of risk-adjusted return that can describe the additional earnings an investor receives per unit of increase, which is shown in Eq.(8).
  • The Alpha measures the portfolio in generating returns above the market benchmark. A positive alpha indicates that the investment portfolio has achieved a higher return than its benchmark after risk adjustment.
  • Beta is the correlation between the portfolio and systemic risk, reflecting the sensitivity of the strategy to the change in the market.
  • Max Drawdown (MD) means the potential worst-case scenario or the most extreme possible loss.
  • Portfolio volatility ($\sigma_{p}$) is used to measure the risk of a strategy, where is positively correlated to the risk of a strategy.
Factor AR (%) AER (%) WR (%) SR Alpha Beta MD (%) $\sigma_{p}$
ARBR -1.47 0.32 54.7 -0.251 -0.002 0.914 34.46 0.221
AR -5.87 -4.37 53.7 -0.406 -0.048 0.981 47.84 0.260
ATR14 -2.12 -0.37 62.7 -0.250 0.006 1.160 46.55 0.250
ATR6 -3.94 -2.31 61.2 -0.328 -0.014 1.175 50.11 0.253
BBIC 1.17 3.14 50.9 -0.109 0.033 1.038 32.64 0.258
BR -6.91 -5.47 53.6 -0.467 -0.064 0.934 51.72 0.254
CCI10 -18.65 -17.99 43.3 -1.402 -0.261 0.812 72.04 0.220
CCI15 -17.70 -16.98 44.2 -1.293 -0.240 0.832 71.43 0.223
CCI20 -15.52 -14.66 46.3 -1.051 -0.196 0.868 69.12 0.235
CR20 -4.27 -2.67 54.8 -0.329 -0.032 0.927 49.92 0.263
DAVOL10 -7.44 -6.04 54.1 -0.504 -0.071 0.920 45.45 0.248
DAVOL20 -10.95 -9.78 54.4 -0.702 -0.119 0.919 48.96 0.246
DAVOL5 -9.54 -8.28 51.8 -0.612 -0.097 0.950 46.41 0.250
EMA5 2.77 4.85 48.3 -0.047 0.049 1.041 30.70 0.259
EMAC10 0.27 2.17 50.6 -0.144 0.024 1.037 31.42 0.258
EMAC12 1.58 3.57 50.6 -0.093 0.037 1.032 31.76 0.258
EMAC20 6.01 8.29 54.1 0.069 0.079 1.044 32.43 0.260
EMAC26 3.83 5.97 54.1 -0.008 0.059 1.046 34.87 0.261
EMAC120 1.74 3.75 48.3 -0.088 0.034 0.960 35.81 0.254
Kurtosis20 -2.96 -1.27 52.5 -0.395 -0.027 0.772 27.71 0.182
Kurtosis60 -0.28 1.59 53.9 -0.240 -0.001 0.716 22.07 0.179
Kurtosis120 -4.12 -2.51 49.0 -0.468 -0.043 0.708 28.28 0.182
MAC5 0.54 2.46 46.9 -0.134 0.026 1.035 30.03 0.257
MAC10 -1.20 0.61 49.6 -0.206 0.008 1.039 34.15 0.256
MAC20 2.47 4.52 53.5 -0.059 0.046 1.043 34.58 0.257
MAC60 11.10 13.72 53.8 0.243 0.120 0.987 28.61 0.254
MAC120 3.11 5.21 47.7 -0.035 0.049 0.987 35.40 0.259
MACDC -11.12 -9.97 52.4 -0.674 -0.117 1.003 46.48 0.261
MASS -10.55 -9.35 51.0 -0.673 -0.115 0.886 48.34 0.249
MAWVAD 7.29 9.67 55.9 0.119 0.082 0.891 36.54 0.248
MFI14 -6.21 -4.73 51.5 -0.456 -0.056 0.918 45.65 0.241
PLRC6 -14.83 -13.92 47.8 -0.905 -0.178 0.960 66.24 0.259
PLRC12 -7.28 -5.87 53.3 -0.452 -0.064 1.012 52.30 0.273
PLRC24 -5.77 -4.26 58.1 -0.361 -0.044 1.034 53.28 0.289
PSY -9.11 -7.83 50.6 -0.660 -0.096 0.880 49.79 0.223
Price1M -13.55 -12.55 53.0 -0.760 -0.156 0.977 64.85 0.280
Price3M -6.69 -5.25 58.4 -0.380 -0.056 1.023 59.14 0.305
Price1Y 1.55 3.54 66.8 -0.075 0.041 1.116 60.25 0.326
ROC6 -16.13 -15.30 47.4 -0.982 -0.200 0.978 68.14 0.262
ROC20 -11.63 -10.51 54.7 -0.628 -0.123 1.014 60.28 0.291
ROC60 -6.86 -5.43 59.6 -0.382 -0.056 1.062 62.16 0.309
ROC120 2.36 4.41 66.9 -0.051 0.049 1.117 56.36 0.319
Skewness20 -5.91 -4.41 50.8 -0.538 -0.058 0.820 40.65 0.197
Skewness60 -6.47 -5.01 51.9 -0.595 -0.068 0.765 38.02 0.190
Skewness120 -4.37 -2.77 51.1 -0.459 0.044 0.743 36.01 0.192
TRIX5 -10.72 -9.54 52.5 -0.607 -0.110 1.009 56.07 0.280
TRIX10 -3.51 -1.85 58.4 -0.268 -0.019 1.008 52.92 0.292
TVMA20 -1.92 -0.16 64.9 -0.209 0.014 1.268 44.66 0.288
TVMA6 -0.75 1.09 62.4 -0.166 0.026 1.259 39.46 0.288
TVSTD20 -1.83 -0.06 58.5 -0.214 0.012 1.218 35.89 0.277
TVSTD6 -3.77 -2.13 53.4 -0.297 -0.010 1.202 46.57 0.272
VDEA -3.05 -1.37 57.6 -0.305 -0.023 0.857 37.01 0.239
VDIFF -7.93 -6.57 0.2 -0.533 -0.079 0.895 39.79 0.247
VEMA5 -0.41 1.46 57.9 -0.186 0.006 0.850 37.31 0.237
VEMA10 0.95 2.91 58.9 -0.128 0.019 0.844 35.51 0.236
VEMA12 1.01 2.96 59.3 -0.126 0.020 0.840 35.17 0.235
VEMA26 2.41 4.46 59.3 -0.068 0.033 0.838 31.03 0.231
VMACD -1.03 0.79 50.5 -0.216 0.001 0.875 33.96 0.235
VOL5 -8.12 -6.76 55.9 -0.452 -0.066 1.155 53.86 0.296
VOL10 -7.07 -5.64 58.3 -0.404 -0.053 1.145 50.65 0.298
VOL20 -5.43 -3.89 58.6 -0.337 -0.032 1.156 50.19 0.298
VOL60 0.42 2.34 58.1 -0.125 0.032 1.145 38.35 0.286
VOL120 -1.88 -0.12 53.0 -0.213 0.007 1.136 42.09 0.281
VOL240 -5.51 -3.98 47.0 -0.382 -0.038 1.075 43.53 0.265
VOSC -6.35 -4.88 51.2 -0.453 -0.056 0.944 40.40 0.246
VROC6 -16.30 -15.49 42.7 -1.179 -0.208 0.897 69.03 0.221
VROC12 -11.47 -10.34 46.7 -0.793 -0.127 0.915 51.62 0.228
VR -8.43 -7.10 52.8 -0.566 -0.083 0.932 48.57 0.244
VSTD10 -0.77 1.07 56.3 -0.205 0.001 0.843 37.75 0.234
VSTD20 -1.60 0.19 53.7 -0.245 -0.007 0.844 37.31 0.232
Variance20 -1.94 -0.18 59.9 -0.201 0.005 1.124 46.38 0.302
Variance60 9.96 12.51 65.2 0.168 0.122 1.178 30.63 0.313
Variance120 5.40 7.65 60.8 0.040 0.083 1.193 39.54 0.312
Volume1M -15.54 -14.68 20.7 -0.867 -0.189 0.991 63.68 0.285
WVAD -0.71 1.13 52.3 -0.199 0.002 0.836 42.38 0.238
arron_down_25
-7.35 -5.94 48.3 -0.593 -0.072 0.880 42.63 0.209
arron_up_25 -15.51 -14.65 46.4 -1.049 -0.195 0.870 66.73 0.235
bear_power -13.31 -12.30 50.8 -0.800 -0.154 0.934 59.58 0.261
beta 10.30 12.87 62.4 0.168 0.134 1.344 35.96 0.328
boll-up 2.19 4.23 54.5 -0.068 0.044 1.062 33.09 0.265
book_to_price 1.87 3.86 58.5 -0.117 0.031 0.889 26.73 0.181
bull_power -14.81 -13.91 51.1 -0.844 -0.176 0.984 65.99 0.278
earnings_yield -2.41 -0.68 52.6 -0.348 -0.029 0.627 25.87 0.188
growth 2.90 4.99 59.9 -0.045 0.052 1.068 37.35 0.244
leverage 2.05 4.07 53.1 -0.089 0.023 0.719 22.57 0.217
liquidity -2.71 -1.00 56.7 -0.243 -0.002 1.137 41.28 0.284
momentum 10.93 13.54 65.8 0.195 0.129 1.159 43.19 0.310
money_flow_20 -1.72 0.05 65.5 -0.202 0.016 1.269 44.10 0.288
price_no_fq 6.83 9.17 76.1 0.095 0.092 1.132 44.86 0.267
pull-down -2.28 -0.54 47.6 -0.317 -0.015 0.831 34.46 0.203
residual_volatility
4.13 6.30 62.3 0.003 0.062 1.048 41.34 0.269
sharpe_ratio_20
-11.88 -10.78 54.0 -0.663 -0.129 0.979 57.66 0.282
sharpe_ratio_60
-7.42 -6.02 58.8 -0.417 -0.065 1.029 65.19 0.300
sharpe_ratio_120
-0.32 1.55 65.7 -0.143 0.019 1.062 60.52 0.303
size -6.36 -4.89 51.8 -0.571 -0.058 0.906 36.00 0.196
turnover_volatility
-10.80 -9.62 51.8 -0.624 -0.108 1.075 57.05 0.274
1day_VPT -16.45 -15.65 44.8 -1.062 -0.211 0.904 70.43 0.248
1day_VPT_6
-6.36 -4.89 53.0 -0.416 -0.053 0.995 47.06 0.269
1day_VPT_12
-0.07 1.82 57.9 -0.146 0.020 1.033 46.51 0.278
25week_close
-11.51 -10.39 51.0 -0.859 -0.133 0.817 47.79 0.211
Trans_Month_1
17.35 19.43 57.8 0.915 0.162 0.588 18.35 0.161
Average -3.78 -2.15 53.71 -0.36 -0.03 0.98 44.88 0.26
CSI300 1.77 \ \ \ \ \ \ 0.197

### v2

@@
-From attention to profit: quantitative trading strategy based on transformer
+Quantformer: from attention to profit with a quantitative transformer trading strategy
 Abstract
 In traditional quantitative trading practice, navigating the complicated and dynamic financial market presents a persistent challenge.
-Former machine learning approaches have struggled to fully capture various market variables, often ignore long-term information and fail to catch up with essential signals that may lead the profit.
-This paper introduces an enhanced transformer architecture and designs a novel factor based on the model.
-By transfer learning from sentiment analysis, the proposed model not only exploits its original inherent advantages in capturing long-range dependencies and modelling complex data relationships but is also able to solve tasks with numerical inputs and accurately forecast future returns over a period.
+Fully capturing various market variables, including long-term information, as well as essential signals that may lead to profit remains a difficult task for learning algorithms.
+In order to tackle this challenge, this paper introduces quantformer, an enhanced neural network architecture based on transformers, to build investment factors.
+By transfer learning from sentiment analysis, quantformer not only exploits its original inherent advantages in capturing long-range dependencies and modeling complex data relationships, but is also able to solve tasks with numerical inputs and accurately forecast future returns over a given period.
 This work collects more than 5,000,000 rolling data of 4,601 stocks in the Chinese capital market from 2010 to 2019.
-The results of this study demonstrated the model’s superior performance in predicting stock trends compared with other 100 factor-based quantitative strategies with lower turnover rates and a more robust half-life period.
-Notably, the model’s innovative use transformer to establish factors, in conjunction with market sentiment information, has been shown to enhance the accuracy of trading signals significantly, thereby offering promising implications for the future of quantitative trading strategies.
+The results of this study demonstrated the model’s superior performance in predicting stock trends compared with other 100 factor-based quantitative strategies.
+Notably, the model’s innovative use of transformer-liked model to establish factors, in conjunction with market sentiment information, has been shown to enhance the accuracy of trading signals significantly, thereby offering promising implications for the future of quantitative trading strategies.
+The implementation details and code is available on https://github.com/QuantFormer.
 keywords:
-Quantitative finance, Transformer, Stock selection, Market sentiment
+quantformer, transformer, quantitative finance, stock selection, market sentiment
 1 Introduction
-The goal of stock trading is to optimize the return on the investment in the capital market according to the process of buying or selling one or more companies’ shares.
-Traders obtain profit when the difference is generated by the fluctuation of stock spice.
-However, stocks are influenced by a large number of factors, which constitute a complex system and increase the difficulty for people to make a profit.
+The goal of stock trading is to optimize the return on investment in the capital market according to the process of buying or selling one or more companies’ shares.
+Traders obtain profit when a positive difference is generated by the fluctuation of stock price.
+However, stocks are influenced by a large number of factors, which constitute a complex system and make it difficult for people to make a profit.
 The assessment of a stock’s evolving trend is inherently challenging due to the highly volatile and interconnected nature of the market, which sets it apart from typical time series modeling [1].
-So, many strategies and tools were built, which corresponded to the development of the capital market and quantitative strategies have played important roles among them.
-Figure 1: The Quantitative Trading Architecture with Factors
-Some traditional quantitative tools, such as Markowitz Portfolio [2] and Capital Asset Pricing Model(CAPM) [3], focuses mainly on the fundamental analysis.
+As a result, many strategies and tools have been built, in parallel with the development of capital markets, and quantitative strategies have been playing an important role among them.
+Figure 1: Quantitative trading architecture with factors
+Some traditional quantitative tools, such as the Markowitz portfolio theory [2] and the Capital Asset Pricing Model (CAPM) [3], focus mainly on static fundamental analysis.
 In other words, these strategies aim to make a profit by simple calculation and analysis.
-However, depending on the development of computer science, more quantitative methods and tools were introduced.
-Within these methods, strategy with factors has attracted people’s attention.
-In 1993, Fama and French [4] introduced their Fama French Three Factor Model (FF3), which is one of the most influential models in quantitative trading.
-In 2015, Fama and French renewed their model, the Five-Factor Asset Pricing Model (FF5) [5] was introduced.
-Besides this classical theory, other trading strategies have been published for decades.
-Quantitative trading with factors typically manifests in two primary strategies, which are shown in Fg.1.
+Since then, along with the development of computer science, more quantitative methods and tools have been introduced.
+Within these methods, factor-based strategies have attracted much attention.
+In 1993, [4] introduced their Fama-French Three Factor Model (FF3), which has become an influential model in quantitative trading.
+In 2015, Fama and French revised their model with a Five-Factor Asset Pricing Model (FF5) [5].
+Besides this classical theory, numerous trading strategies have been published for decades.
+[6] built a optimal portfolio to catch future investment opportunities (FIO) by multi-factor models.
+Quantitative trading with factors typically follows two primary approaches, which are shown in Figure 1.
 The first approach involves the computation of stock factor values.
@@
 Traders can execute corresponding actions when they receive the signals from factors, allowing for a dynamic response to market fluctuations based on factor insights.
-In recent years, Machine Learning (ML) stands out as an instrumental tool in enhancing trading algorithms and decision-making processes.
+In recent years, Machine Learning (ML) has become an instrumental tool for enhancing trading algorithms and decision-making processes.
 Its core principle, enabling systems to learn from and make decisions based on data, lends itself particularly well to the vast and dynamic landscapes of stock markets.
-Some traditional literature exploits machine learning algorithms based on manual indicators [6], but the hypothesis about the random walk may be less useful in the task of catching non-stationary series.
-With the rise of different architectures, quantitative researchers found some novel approaches to identifying intricate patterns and factors in trading data, which opened a new field in quantitative trading [7].
-Although there are several previous experiments about ML factors in the quantitative finance field, the research still faces two concerns.
-Firstly, in the field of NLP, it is necessary to convert words in text to word vectors through word embeddings to serve as inputs.
-However, in the financial dataset, there are categorical data such as industry types, as well as numerical data like price fluctuation, turnover rate, and financial indicators.
+[7] exploits machine learning algorithms based on manual indicators, but the underlying random walk hypothesis may hamper the task of understanding inherently non-stationary series.
+With the rise of different architectures, [8] finds some novel approaches to catch the sentiment of the market in trading data, which opened a new field in quantitative trading.
+Although there exist several previous experiments with ML factors that attempted to fetch market sentiment in the quantitative finance field, this research still faces two difficulties.
+Firstly, in the field of sentiment analysis, which is a branch of Natural Language Processing (NLP, a field of computer science that aims to understand, interpret, and generate human language), models are used to convert words in text to word vectors through word embeddings to serve as inputs.
+However, financial datasets contain categorical data instead of words such as industry types, as well as quantitative data such as price fluctuation, turnover rate, and financial indicators.
 If the input comprises only categorical data, the time series can be treated as a sentence [9].
 In most cases, the input will involve numerical data, which cannot be transformed via word embeddings.
-For the second concern, in NLP tasks, most of them can be transformed into sequence-to-sequence (seq2seq) problems, such as in machine translation, dialogue systems, and speech recognition.
-To utilize existing outputs, decoders in NLP tasks sequentially output samples and use masking operations to handle input sequences during training.
+Secondly, most NLP tasks can be transformed into sequence-to-sequence (seq2seq) problems, such as in machine translation, dialogue systems, and speech recognition.
+As an example, the transformer architecture is based on the seq2seq architecture [10].
+To utilize existing outputs, decoders in transformer sequentially output samples and use masking operations to handle input sequences during training.
 However, in stock prediction, where the aim is often to accurately forecast future returns over a period, the transformer model is rarely used for such tasks.
-To address these concerns, we propose an approach that combines time series stock data with a numerical-adapted transformer to establish a factor and put it into back-tests.
-This approach allows us to use transformer model embedding numerical input directly, which refers to a method similar to sentiment analysis.
-The structure of the paper is designed as follows.
+To address these problems, we propose quantformer, which is a modified transformer architecture adapted to quantitative data, and used as an investment factor.
+Quantformer is able to input numerical data directly, which refers to a method similar to sentiment analysis.
+The paper is structured as follows.
 Section 2 discussed the previous quantitative financial works mainly based on machine learning.
-Section 3 and 4 introduce the establishment of the model.
-A factor based on the model will be trained and back-tested.
-For the practical test, we collected data from more than 4,600 stocks in the previous 14 years (from 2010 to 2023) from the data-collect platform.
-To comprehensively test the ability of the factor, we divided of data by different frequencies (monthly, weekly, and daily).
-Finally, the result of the back-tests including the comparison between the transformer-factor and others and the insights gained from such comparative analysis will be discussed in the section 4 and 5.
-Figure 2: Example for stock prediction
+Sections 3 and 4 introduce quantformers.
+A factor based on quantformer will be trained and back-tested.
+For the practical backtest, we collected data from more than 4,600 stocks in the past 14 years (from 2010 to 2023) from the data-collect platform.
+To comprehensively test the ability of the factor, we divided the data by different frequencies (Section 4.2) and trained under different training scales (Section 5.3).
+Finally, the result of the back-tests including the comparison between the quantformer-factor and other 100 factors as well as the insights gained from such comparative analysis will be discussed in Sections 4 and 5.
+Figure 2: Example of stock prediction
 2 Preliminaries
-This section briefly introduces the related works about the paper and the development of quantitative financial trading methods, especially based on Machine Learning (ML) methods.
-2.1 Stock data with market sentiment
-Stock data can be considered that can influence the future trend of this stock in some cases [9].
-Previous works have discussed the correlation between investor sentiment and stock market metrics like turnover rate, indicating that investor feelings, both individual and institutional, significantly influence market behavior, affecting aspects such as volatility and returns [10, 11].
-This relationship isn’t uniform but varies with changes in investor sentiment.
-Research also highlights that investor psychology, especially during periods of high market stress can drastically impact market behavior and investor decisions.
-This link between sentiment and market performance is further affirmed by studies focusing on factors like the turnover ratio, which are seen as reflective of market liquidity and investor behavior [12].
+This section briefly introduces the related works about stock prediction with market sentiment and the development of quantitative financial trading methods, especially based on Machine Learning (ML) methods.
+2.1 Stock data and market sentiment
+It is known that, in some cases, stock data can be deemed to influence the future trend of a stock price [11].
+[12, 13] have discussed the correlation between investor sentiment and stock market metrics such as turnover rate, indicating that investor feelings, both individual and institutional, significantly influence market behavior, affecting aspects such as volatility and returns.
+This relationship is not uniform but varies with changes in investor sentiment.
+It also highlights that investor psychology, especially during periods of high market stress can drastically impact market behavior and investor decisions.
+This link between sentiment and market performance is further affirmed by studies focusing on factors such as the turnover ratio, which are seen as reflective of market liquidity and investor behavior [14].
+[15] built a new factor based on sentiment analysis with the average of trading signals from technical trading strategies to benchmark stocks of the S&P 500 index and DJIA.
+The sentiment factor shows correlation with stock returns.
 These findings underscore the mutual relationship between investor sentiment and market performance, demonstrating how psychological factors can drive market dynamics.
-2.2 Applications for ML methods in trading
+2.2 Applications of ML methods for trading
 Machine learning exploits a range of algorithms and statistical techniques used for tasks such as regression, clustering, and classification.
-In past decades, there have been methods that have been significantly increasingly applied in stock due to their ability to process vast amounts of data and make predictions, offering a potential advantage in the financial markets.
+In past decades, there have been methods that have been increasingly applied to stock data due to their ability to process vast amounts of data and make predictions based on them, offering a potential advantage in the financial markets.
 At the same time, the finance sector, particularly quantitative trading, has started to be aware of the potential of deep learning models for predicting stock movements, portfolio optimization, and risks.
 2.2.1 Support Vector Machines (SVM)
-SVM was widely used in quantitative trading, which can manage high-dimensional spaces and intricacies found in datasets.
-In 2003, SVM tried to forecast financial time series [13], especially with the lens focused on the South Korean stock market.
-This work exhibited the robustness of SVM against traditional statistical methods and some other computational techniques, setting a precedent for its adoption in financial forecasting.
-Similar to the previous one, [14] showcased SVM’s prowess in anticipating the movement direction of Japanese stock markets, especially when determining bullish or bearish market sentiments.
-[15], found SVM’s adaptability and precision in financial tasks like portfolio and quantitative trading.
-However, SVM also shows some disadvantages in quantitative trading.
-In 2009, SVM was underscored for the accuracy of overfitting [16].
-When applied to financial datasets, the model’s propensity to fit too closely to the training data might lead to reduced generalization capabilities, causing the strategy ineffective on unseen data.
-In high-frequency trading, though the model displayed competency in specific scenarios, researchers pinpointed potential limitations when grappling with certain volatile market dynamics, suggesting the need for adaptive [17].
+SVM has been used in quantitative trading, as it can manage high-dimensional spaces and intricacies found in datasets.
+[16] used SVMs to forecast financial time series in 2015, with a focus on the South Korean stock market, exhibiting the robustness of SVM against traditional statistical methods and some other computational techniques, setting a precedent for its adoption in financial forecasting.
+Similarly, [17] showcased SVMs’ prowess in anticipating the movement direction of Japanese stock markets, especially when determining bullish or bearish market sentiments.
+[18] assessed SVMs’ adaptability and precision in financial tasks such as portfolio and quantitative trading.
+However, SVMs can also show some disadvantages in quantitative trading.
+In [19], SVM was underscored for the accuracy of overfitting.
+When applied to financial datasets, the model’s propensity to fit too closely to the training data might lead to reduced generalization capabilities, making the strategy ineffective on unseen data.
+In high-frequency trading, though the model displayed competency in specific scenarios, researchers pinpointed potential limitations when grappling with certain volatile market dynamics, suggesting the need for adaptations [20].
 2.2.2 Long Short-Term Memory (LSTM)
-LSTM was designed to capture long-term dependencies in sequential data according to their unique architecture [18].
+LSTM is a neural network architecture designed to capture long-term dependencies in sequential data [21].
+LSTM neural networks can accentuate and effectively model volatile financial markets, showcasing significant improvements over traditional time series models.
+[22] combined LSTM with linear Kalman filter and set experiments on different volatility stocks to build strategies.
 [23] deployed LSTM with stacked autoencoders to forecast financial time series.
-LSTM shows the ability to accentuate and effectively model volatile financial markets, showcasing significant improvements over traditional time series models.
-In 2018, [20] tested the efficacy of LSTM for price predictions.
-They observed LSTMs’ prowess in noisy data and their superiority in prediction accuracy over some conventional models.
-LSTM’s capability to extract features from time series was explored by [21].
-They combined LSTM with other architectures and found that these can enhance the feature extraction process, thereby improving prediction accuracy.
+[24] tested the efficacy of LSTM for price predictions they observed LSTMs’ prowess in noisy data and their superiority in prediction accuracy over other conventional models.
+[25] combined LSTM with other architectures and found that these can enhance the feature extraction process, thereby improving prediction accuracy.
 For predicting stock prices, [26] harnessed the power of LSTM to unearth multi-frequency trading patterns.
-Their research underlined the adaptability of LSTMs in modeling complex situations inherent in stock prices.
-2.2.3 GRU
-Gated Recurrent Units (GRU) [23] have emerged as a useful method in the quest for accurate financial forecasting.
-GRU can capture long-term dependencies in time series data and has seen increasing popularity in the world of quantitative trading.
-[19] introduced a comprehensive deep learning framework for financial time series, which Incorporated GRUs alongside LSTM and stacked autoencoders.
+Their method underlined the adaptability of LSTMs in modeling complex situations inherent in stock prices.
+2.2.3 Gated Recurrent Units (GRUs)
+GRUs [27] are another neural network architecture which has emerged as a method in the quest for accurate financial forecasting.
+GRUs can capture long-term dependencies in time series data and have seen increasing popularity in the world of quantitative trading.
+[23] introduced a comprehensive deep-learning framework for financial time series, which incorporated GRUs alongside LSTM and stacked autoencoders.
 Their model, when applied to diverse datasets, performs better than traditional time series models, accentuating the adaptability and robustness of GRUs.
@@
 However, in their analysis, LSTMs showed a slight edge in prediction accuracy, indicating that while GRUs are powerful, selecting between them and LSTMs may boil down to specific use cases and computational constraints.
-What’s more, [25] presented a model fusing GRUs with Convolutional Neural Networks (CNN) [26].
-The strategy combined CNNs and GRUs, also received advantages among them, creating a formidable solution for stock price prediction.
+Beyond that, [29] presented a model fusing GRUs with Convolutional Neural Networks (CNN) [30].
+The strategy to combine CNNs and GRUs also exhibits advantages for stock price prediction.
+[31] introduced the Attention-GRU model which combined attention and GRU to establish a new factor based on CVaR [32, 33] portfolio.
+The Attention-GRU model fitted market return, stocks return from 28 Dow Jones Industrial Average index (DJIA) stocks, which achieved better performance than other models in 8 metrics such as annual return, standard deviation and information ratio.
 2.3 Transformer in quantitative trading
-2.3.1 Introduction for transformer
-Transformer [27], which was introduced in 2017, has received high attention in the NLP field.
+Transformer [10], which was introduced in 2017, has received considerable attention in the NLP field.
 According to its special encoder and decoder stacks with self-attention blocks, the transformer shows its advantage in stability, speed, and long-term memory compared with other traditional models.
-The transformer-based models like GPT [28] and BERT [29] are the basis of the specific field of application.
-In the application field, transformer-based models like ChatGPT [30], FinBERT [31, 32] and BloomberGPT [33] show advantages and ability in solving different types of problems in different filed.
-Similar to other neutral sequence transduction models, transformer also has an encoder-decoder structure.
-The model takes input as a sequence of context which could be natural language sentences and generates output with texts depending on the input, like a translation according to the input or a predicted sentence.
+The transformer-based models such as GPT [34] and BERT [35] are the basis of the specific field of application.
+In the application field, transformer-based models such as ChatGPT [36], FinBERT [37, 38] and BloomberGPT [39] show advantages and ability in solving different types of problems in different fields.
+[40] utilized transformer to to predict the stock market index (including CSI 300, S&P 500, Hang Seng Index, and Nikkei 225) and concluded that the model could better catch the rules of stock market dynamics.
+Similar to other neural sequence transduction models, transformer also has an encoder-decoder structure.
+The model takes as input a sequence of context which could be natural language sentences and generates output with texts depending on the input, similar to a translation according to the input or a predicted sentence.
 Transformer using stacks which are called self-attention and point-wise, which connect layers in the architecture for encoder and decoder.
@@
 The linear and softmax functions are used in the dot-product attention module.
-In the vanilla transformer model, there are multi-head attention layers are used and the result from each layer will be outputted after going through an additional &\&& normalize layer.
-The above-mentioned steps covered most of the progress in the vanilla transformer.
-2.3.2 Previous transformer-based strategy
-Transformer has achieved superior performance on long-term dependencies modeling.
+In the vanilla transformer model, multi-head attention layers are used and the result from each layer will be outputted after going through an additional normalization layer.
+The above-mentioned steps cover most of the progress in the vanilla transformer.
 In the quantitative trading domain, some previous researchers tried to test the potential of transformer.
@@
 [42] devoted to optimizing the efficiency of time complexity and memory usage of transformer on extremely long time series by informer.
-[36] leveraged the advantages of CNNs and Transformers to model short-term and long-term dependencies in financial time series.
+[43] takes the advantage of CNNs and transformers to model short-term and long-term dependencies in financial time series.
 They showed the advantage of the proposed approach on intraday stock price prediction of S&P 500 constituents.
-Similar to previous work, [1] innovated the structure of a Transformer, which is Adaptive Long-Short Pattern Transformer (ALSP-TF).
+[44] proposed a new ConvLSTM model by combining convolutional transformer and LSTM.
+The model was trained by stocks from S&P 500 from 2004-2021 and performs better than other models such as VaR and ARIMA.
+Similar to previous work, [1] innovated the structure of a transformer, in the form of Adaptive Long-Short Pattern Transformer (ALSP-TF).
 Their model is structurally innovated for hierarchical representation and interaction of stock price series at different context scales.
 With the help of a learnable function, they make the self-attention aware of the weighted time intervals between patterns, to adaptively adjust their dependencies beyond similarity matching.
-Finally, they obtained more than 10% of annual return on average.
-However, there are some disadvantages of transformer.
+In the end, they obtained more than 10% of annual return on average.
+However, transformer has some disadvantages.
 [45] mentioned that the global self-attention module focuses on point-wise token similarities without contextual insights.
@@
 3 Methodology
-This section introduces the framework of the work and the steps of the establishment of the model, including the data processing, model enhancement and output processing.
+This section introduces the framework of the work and the steps of the establishment of the model, including the data processing, quantformer construction and prediction.
 3.1 Framework overview
-Fig. 3 illustrates the overview of the structure. There are three major parts:
+Figure 3 gives an overview of the structure. There are three major parts:
 (i) Data Initialization to align the time series of stocks into a regular matrix;
-(ii) Embedding the training dataset and training them by transformer which replaces the word-embedding layer to linear layer and deletes the Mask(opt.) layer;
+(ii) Embedding the training dataset and training them by a quantformer with word-embedding layer replaced by linear layer and with no Mask(opt.) layer;
 (iii) Lastly, using the trained model to predict the possible trend of stocks.
@@
 $S^{t}=\{s_{1}^{t},s_{2}^{t}\ldots s_{N}^{t}\}$
-For a stock $s_{k}^{t}$ in stock set $S^{t}$, where $k\in N$, it details a sequence for a historical 1-time step (such as one month), and the dimension is 2:
+For a stock $s_{k}^{t}$ in the stock set $S^{t}$, where $k\in\{1,\ldots,N\}$, consider the two-dimensional historical 1-time step sequence (such as one month):
 ${\chi}^{t}_{k}=[x_{k}^{t-20},x_{k}^{t-19}\ldots x_{k}^{t}]\in\mathbb{R}^{20\times 2}$
-In each $x_{k}^{t}$, there are two features.
-The first one is the accumulated daily profit rate during the time step, where $p_{i}$ and $p_{i-1}$ is the close price of the $i^{th}$ and ${i-1}^{th}$ day, respectively.
+Each $x_{k}^{t}$ contains two features.
+The first one is the accumulated daily profit rate during the time step, where $p_{i}$ and $p_{i-1}$ are the close price of the $i^{\mathrm{th}}$ and ${(i-1)}^{\mathrm{th}}$ day, respectively.
 $r^{t}=\sum r_{i}=\sum\frac{p_{i}-p_{i-1}}{p_{i-1}},\quad\sigma^{t}=\sum\sigma_{i}$ (1)
-Another is the accumulated daily turnover rate $\sigma^{t}$ during the time step and the calculation method of $\sigma^{t}$ is shown above. So each $\chi^{t}_{k}$ can be represented as:
+The other feature is the accumulated daily turnover rate $\sigma^{t}$ during the time step and the calculation method of $\sigma^{t}$ is shown above. So each $\chi^{t}_{k}$ can be represented as:
 $\chi^{t}_{k}=\left[\begin{matrix}r^{t-20}&r^{t-19}&\ldots&r^{t}\\ \sigma^{t-20}&\sigma^{t-19}&\ldots&\sigma^{t}\\ \end{matrix}\right]\in\mathbb{R}^{20\times 2}$
 Outputs
-For a stock set $S^{t}$ on the trading timestamp $t$, each stock $S^{t}_{i}$ has a profit rate $r_{i}^{t+1}$ on the trading timestamp $t+1$, where the calculation method is same as the method of $r^{t}$ in the sequence of inputs shown in Eq.(1).
+For a stock set $S^{t}$ on the trading timestamp $t$, each stock $s^{t}_{i}$ has a profit rate $r_{i}^{t+1}$ on the trading timestamp $t+1$, where the calculation method is the same as for $r^{t}$ in the sequence of inputs shown in equation (1).
 Then the value in the list of next-time stamp’s profit for $N$ stocks in the timestamps $t+1$ can be represented as:
 $P^{t+1}=\{r_{1}^{t+1},r_{2}^{t+1}\ldots r_{N}^{t+1}\}$
-And the list will be arranged as $n$ parts and each part has an equal proportion $\frac{1}{n}$ ($n$ should be an odd number, such as 3,5,7).
-Take $n=5$ as an example.
-For those ranked values in the first, middle and last $\frac{1}{5}=20\%$, the output of the stock $s_{k}$ records as below, respectively:
-$y^{t}_{k}=[1,0,0]^{T},\quad y^{t}_{k}=[0,1,0]^{T},\quad y^{t}_{k}=[0,0,1]^{T}$
+This list is then sorted and partitioned into $q$ equal parts, where it is set to 3 or more.
+For example, set $q=5$; for the stocks $s_{k}$ whose profit values ranked in the bottom $20\%$, middle $20\%$, and top $20\%$ respectively.
+And the dimension of the outputs $d_{f}$ is the number of selected quantiles in training.
+The value of $d_{f}$ should be larger than 3 and not larger than $q$.
+Figure 4: Training outputs remarked between different $d_{f}$
+Both the dimensions of training and predicted outputs are the same, which is the $d_{f}$ set at the beginning.
+For example, set $q=5$; for the stocks $s_{k}$ whose profit values ranked in the bottom $20\%$, middle $20\%$, and top $20\%$ respectively, and the corresponding output of $d_{f}=3$ is recorded, respectively, as
+$y^{t}_{k}=[1,0,0]^{T},\quad y^{t}_{k}=[0,1,0]^{T},\quad y^{t}_{k}=[0,0,1]^{T},\quad y^{t}_{k}=[0,0,1]^{T}$
+The dimension of the output also can be 5, and in this case, the outputs of $d_{f}=5$ are:
+$y^{t}_{k}=[1,0,0,0,0]^{T},\quad y^{t}_{k}=[0,0,1,0,0]^{T},\quad y^{t}_{k}=[0,0,0,0,1]^{T}$
+In this example, when $d_{f}=3$, the output dimension will be 3 while the dimension will be 5 as $d_{f}$ is set as 5.
 3.3 Embedding initialization
 Normalization
-The next step is the normalize the input data for each time step by zero-mean normalization (Z-Score), the normalization equation is shown in Eq.(2).
-$\tilde{x}^{t}_{k,i}$ represents the normalized value at $i^{th}$ day of stock $k$ in time $t$.
-$E[x^{t}_{i}]$ and $std[x^{t}_{i}]$ is the mean and standard of all the value of $x^{t}_{i}$, respectively.
-$\tilde{x}^{t}_{k,i}=\frac{x^{t}_{k,i}-E[x^{t}_{i}]}{std[x^{t}_{i}]}$ (2)
-In this way, the input consequences towards zero mean and unit variance, which aims to reduce the influence from different time points and allow different features to be comparable with each other [39].
-3.4 Transformer encoder
-Then a transformer encoder structure will be built in this subsection.
+The next step is to normalize the input data for each time step by zero-mean, unit-variance normalization (Z-score), the normalization equation is shown in equation (2).
+$\tilde{x}^{t}_{k,i}$ represents the normalized value at $i^{\mathrm{th}}$ day of stock $k$ at time $t$.
+$\mathbb{E}[x^{t}_{i}]$ and $\mathrm{std}[x^{t}_{i}]$ are the mean and standard deviation of the variables $x^{t}_{i}$, respectively.
+$\tilde{x}^{t}_{k,i}=\frac{x^{t}_{k,i}-\mathbb{E}[x^{t}_{i}]}{\mathrm{std}[x^{t}_{i}]}$ (2)
+In this way, the input sequences are normalized with zero mean and unit variance, which aims to reduce the influence from outlying time points and allow different features to be comparable with each other [47].
+3.4 Quantformer encoder
+Then a quantformer encoder structure will be built in this subsection.
 The representation hierarchy consists of $L$ blocks of multi-head self-attention layers.
-Taken initialized stock embedding sequences $\textbf{P}\in\mathbb{R}^{N\times 20\times 2}$ as input, and the canonical self-attention in [27] performs information exchange between every time points for each stock $s_{i}$.
+Taking initialized stock embedding sequences $\textbf{P}\in\mathbb{R}^{N\times 20\times 2}$ as input, the canonical self-attention in [10] can perform information exchange between every time points for each stock $s_{i}$.
 To process both categorical and numerical data, the word embedding layer is replaced by a standard linear layer, utilizing linear transformations to substitute the process of word embedding.
-On the other side, in stock prediction, we aim to accurately forecast the return for a future period, thus the model’s output is generally a single value or the probability of price increase or decrease.
+On the other side, in stock prediction, we aim to accurately forecast the return for a future period, thus the model’s output is generally a single value representing the probability of price increase or decrease.
 Therefore, the decoder is simplified by removing the autoregressive prediction mechanism (in the decoder) and the masking operations.
-For each $\textbf{P}_{i}\in\mathbb{R}^{20\times 2}$, transforms into query, key, and value metrics, respectively:
-$\textbf{Q}_{i,h}=\textbf{E}_{i}\textbf{W}^{Q}_{h},\quad\textbf{K}_{i,h}=\textbf{E}_{i}\textbf{W}^{K}_{h},\quad\textbf{V}_{i,h}=\textbf{E}_{i}\textbf{W}^{V}_{h}$ (3)
-where $h=1,\ldots,H$ is the head index, as well as $\textbf{W}^{Q}_{h}$, $\textbf{W}^{Q}_{h}$, $\textbf{W}^{Q}_{h}$ $\in\mathbb{R}^{2\times d_{f}}$.
-Lastly, the final layer is represented by the concatenation of all attention heads.
-$\textbf{F}_{i}=Multihead(\textbf{Q}_{i,h},\textbf{K}_{i,h},\textbf{V}_{i,h})=||^{h=H}_{h=1}Softmax(\textbf{Q}_{i,h},\textbf{K}_{i,h}^{T}/\sqrt{d_{f}})\textbf{V}_{i,h})$ (4)
-Just like Eq.(4) shows, ||||| | represents the concatenation operator, and $d_{f}$ is the dimension of projected feature space. In the end, all input consequences are represented in the form of:
+More specifically, each $\textbf{P}_{i}\in\mathbb{R}^{20\times 2}$ is transformed into query, key, and value metrics, respectively:
+$\textbf{Q}_{i,h}=\textbf{P}_{i}\textbf{W}^{Q}_{h},\quad\textbf{K}_{i,h}=\textbf{P}_{i}\textbf{W}^{K}_{h},\quad\textbf{V}_{i,h}=\textbf{P}_{i}\textbf{W}^{V}_{h}$ (3)
+where $h=1,\ldots,H$ is the head index, as well as trainable weights $\textbf{W}^{Q}_{h}$, $\textbf{W}^{Q}_{h}$, $\textbf{W}^{Q}_{h}$ $\in\mathbb{R}^{2\times d_{f}}$.
+Then, the final layer is represented by the concatenation of all attention heads.
+$\textbf{F}_{i}=\mathrm{Multihead}(\textbf{Q}_{i,h},\textbf{K}_{i,h},\textbf{V}_{i,h})=\big{(}||^{h=H}_{h=1}\mathrm{Attention}(\textbf{Q}_{i,h},\textbf{K}_{i,h},\textbf{V}_{i,h})\big{)}\textbf{W}^{O}$ (4)
+where ||||| | represents the concatenation operator and $d_{f}$ is the dimension of the projected feature space.
+After the attention computation is completed, the multi-head attention output is divided into different ”heads”.
+These outputs from the heads need to be recombined back into the original input dimension.
+$\textbf{W}^{O}$ $\in\mathbb{R}^{Hd_{f}\times 2}$ in equation 5 plays the role of combining these separate heads back into the original dimension.
+The attention head is shown below:
+$\mathrm{Attention}(\textbf{Q}_{i,h},\textbf{K}_{i,h},\textbf{V}_{i,h})=\mathrm{softmax}(\textbf{Q}_{i,h},\textbf{K}_{i,h}^{T}/\sqrt{d_{f}})\textbf{V}_{i,h}$ (5)
+In the end, all input sequences are represented in the form of:
 $\textbf{F}=[\textbf{F}_{1};\textbf{F}_{2};\ldots\textbf{F}_{N}]\in\mathbb{R}^{N\times 20\times Hd_{f}}$
-which is followed by feed-forward layers.
-According to Sec.3.2, the embedding output is denoted as:
-$\textbf{Y}^{t}=\{y^{t}_{1},y^{t}_{2},\ldots y^{t}_{N}\}\in\mathbb{R}^{N\times 1\times n}$
+which are fed into feed-forward layers. From Section 3.2, the embedding output is denoted as:
+$\textbf{Y}^{t}=\{y^{t}_{1},y^{t}_{2},\ldots y^{t}_{N}\}\in\mathbb{R}^{N\times 1\times d_{f}}$
 3.5 Output process
-Eventually, the output is represented:
-$\textbf{Z}^{t}=\{z^{t}_{1},z^{t}_{2},\ldots z^{t}_{N}\}\in\mathbb{R}^{N\times 1\times n}$ (5)
-The output will be activated by the $softmax$ function. And for each $z^{t}_{i}$, there are:
-$z^{t}_{i}=[z^{t}_{i,1},\ldots z^{t}_{i,m},\ldots z^{t}_{i,n}]\in\mathbb{R}^{1\times n}$
-the output of probability can be represented as (m is the middle number from 1 to n):
-$Y^{t}_{o}=[y^{t}_{o,1},\ldots y^{t}_{o,m},\ldots y^{t}_{o,n}]^{T}=[\frac{e^{z_{1}}}{\sum_{i=1}^{n}e^{z_{i}}},\dots\frac{e^{z_{m}}}{\sum_{i=1}^{n}e^{z_{i}}},\ldots\frac{e^{z_{n}}}{\sum_{i=1}e^{z_{i}}}]$ (6)
-In Eq.(6), $y^{t}_{o,i}$ are larger than 0 and smaller than 1 as well as the sum of all the items in $Y^{t}_{o}$ is 1.
-So the output means the probability of the stock’s raising, flatting, or downing in the next timestamp.
+Eventually, the output is represented as:
+$\textbf{Z}^{t}=\{z^{t}_{1},z^{t}_{2},\ldots z^{t}_{N}\}\in\mathbb{R}^{N\times 1\times d_{f}}$ (6)
+The output layer uses the softmax activation function, and for each $z^{t}_{i}$ can be written as:
+$z^{t}_{i}=[z^{t}_{i,1},\ldots z^{t}_{i,m},\ldots z^{t}_{i,d_{f}}]\in\mathbb{R}^{1\times d_{f}}$
+($m=1,2,\ldots,n$ is the index number between $1$ and $n$). The output probability can be represented as:
+$Y^{t}_{o}=[y^{t}_{o,1},\ldots y^{t}_{o,m},\ldots y^{t}_{o,d_{f}}]^{T}=\left[\frac{e^{z_{1}}}{\sum_{i=1}^{n}e^{z_{i}}},\dots\frac{e^{z_{m}}}{\sum_{i=1}^{n}e^{z_{i}}},\ldots\frac{e^{z_{n}}}{\sum_{i=1}e^{z_{d_{f}}}}\right]$ (7)
+In equation (7), the $y^{t}_{o,i}$ take values between 00 and $1$ and sum to $1$, which means that they can be interpreted as the probability of the stock’s performance in the slice $m$ in the next timestamp
 3.6 Prediction
@@
 MSE Loss is a widely used metric for assessing the discrepancy between a model’s predictions and the actual values.
-It calculates the average of the squares of the differences between the predicted value and the actual prices.
-This is formally represented as:
-$MSELoss=\frac{1}{n}\sum^{n}_{i=1}(y_{i}-\hat{y}_{i})^{2}$ (7)
-In Eq.(7), $y_{i}$ is the actual value of the $i^{th}$ equation, $\hat{y}_{i}$ is the model’s prediction for the same, and n represents the total number of samples in the inputs.
-During each epoch, the MSE Loss is computed to guide the optimization process, to minimize this value over successive iterations.
+It is defined as the average of the squares of the differences between the predicted value and the actual prices:
+$\mathrm{MSELoss}=\frac{1}{N}\sum^{N}_{i=1}(y_{i}-\hat{y}_{i})^{2}$ (8)
+In equation (8), $y_{i}$ is the actual value of the $i^{\mathrm{th}}$ observation, $\hat{y}_{i}$ is the corresponding model prediction, and $N$ is the total number of input samples.
+During each epoch, the MSE Loss is computed to guide the optimization process, to minimize this loss over successive iterations.
 A lower MSE indicates a closer alignment between the model’s predictions and the actual stock prices, signifying an improvement in the model’s learning and its ability to generalize from the training data.
-This loss function is particularly useful in regression tasks, as it not only penalizes the model for inaccurate predictions but also emphasizes reducing larger errors, thereby ensuring a robust learning process.
 4 Experiments
@@
 4.1 Dataset
-The training data of the transformer factor comes from the Chinese exchange market.
+The training data of the quantformer comes from the Chinese exchange market.
 The dataset is collected from the Shanghai Stock Exchange (SHSE) and Shenzhen Stock Exchange (SZSE), which contains 4601 stocks that are listed or had been listed from January 2010 to May 2023.
-And all these data come from AKShare [^1] https://akshare.akfamily.xyz/ and Tushare [^2] https://www.tushare.pro/, which is one of the quantitative finance terminals.
-The train period is from January 2010 to December 2019 and the test period starts from January 2020.
-Adjustment closing price is used for the stock training price [40].
-Adjustment closing price is essential in stock market analysis, particularly when analyzing historical data for long-term trends and patterns.
-This price takes into account factors like dividends, stock splits, and other corporate actions that can affect a stock’s price over time [41].
-By adjusting for these events, the adjusted close price provides a more accurate picture of a stock’s value and performance, which is more persuadable in quantitative financial back-test.
+The data comes from AKShare [^1] https://akshare.akfamily.xyz/ and Tushare [^2] https://www.tushare.pro/, which are quantitative finance terminals.
+The training period is from January 2010 to December 2019 and the testing period starts from January 2020.
+Closing price adjustments are used for the stock training price [48].
+Closing price adjustments are essential in stock market analysis, particularly when analyzing historical data for long-term trends and patterns.
+This price takes into account factors such as dividends, stock splits, and other corporate actions that can affect a stock’s price over time [49].
+By adjusting for these events, the adjusted closing price provides a more accurate picture of a stock’s value and performance, which is more appropriate for financial backtesting.
 4.2 Timestamp
-To test the performance of the factor better, three kinds of frequency are chosen, which are stock data monthly, weekly, and daily, respectively.
+Data frequency concerns the number of data points within a specific unit of time, which may reflect different market characteristics [50].
+To better test the performance of the factor, three frequencies are considered: monthly, weekly, and daily stock data.
 In the first experiments, a timestamp was set as one month.
-So for a stock $s_{i}$ from stock set $S^{t}$, it details a feature sequence in which each item in the sequence contains the accumulated profit and accumulated turnover rate in a month.
+So for a stock $s_{i}$ from the stock set $S^{t}$, it details a feature sequence in which each item in the sequence contains the accumulated profit and accumulated turnover rate in a month.
 Both the accumulated profit and accumulated turnover rate are the sum of the trading day’s data in that month.
 In some cases, if one stock was first listed on the market or resumed trading in one month, the available data may not cover whole trading days, the data of the stock from that from would still be recorded.
-However, if there is a whole month that the stock did not have the trading record (perhaps stopped trading or not listed in that month), this month’s data will be recorded as ”NaN”.
+However, if there is a whole month during which the stock did not have any trading (perhaps due to stopped trading or not being listed in that month), this month’s data will be recorded as “NaN”.
 In the sequence $\chi^{t}_{k}$, the sequence will not be used to train the model when there are missing values in the sequence.
 Within the first experiments, three sub-experiments were set.
-For the sub-experiments, the input of them are same.
+The sub-experiments share the same inputs.
 For output data, outputs with different dimensions and different lengths but the same dimension are used.
 $n=5$ and $n=7$ are used in this experiments.
-In the group of $n=5$, the first one deleted the data with output like $[0,0,0]^{T}$, and the second one used the origin dataset (zero-output).
-Most of the accumulated profit values in different years are in the range $(-10\%,10\%)$.
-There are some values point outside the range $(-100\%,100\%)$ that are shown in it because these values are too large (small), such as the stock ”002961.SZ” first listed at SZSE on September $4^{th}$, 2019 and its accumulated profit rate on that month is 184.66%.
-These stocks received unusual performance mainly because they were first listed on the market, resumed trading, or reported great positive/negative news.
-Most of the accumulated turnover rate values are in the range $(0,100\%)$.
-There are some values outside the range $(0,200\%)$ that are shown in it because these values are too large, such as the accumulated turnover rate in July 2015 of ”002596.SZ” is $777.67\%$, this was because the actual controller abundantly increased his holdings of the company and the company announced that they received findings by shareholders.
-Similar to the example, the situation may be caused by great positive or negative information.
+In the group of $n=5$, the first one deleted the data with output such as $[0,0,0]^{T}$, and the second one used the origin dataset (zero-output).
 Outliers of both accumulated profit and accumulated turnover rate will not be removed from the dataset these situations may happen in the future and they are expected to be predicted, though these incidents rarely happen.
@@
 The sub-experiments under different groups are similar to the first one.
-All the accumulated number of parameters, the number of trained sections (such as 100 months from 2010 to 2019), and dimension are shown in the Table 1.
+All the accumulated number of parameters, the number of trained sections (such as 100 months from 2010 to 2019), and dimension are shown in Table 1.
 Table 1: Detailed information of experiments
@@
 The training process and model are implemented with PyTorch [51].
-The parameter of the model is searched by gird search, which is reliable in low dimensional spaces and simple to implement [43].
-As a result, in the situation of input dimension is 2, the dimension of hidden feature space $d_{f}$ is 16.
+The parameters of the model are optimized by grid search, which is reliable in low dimensional spaces and simple to implement [52].
+Here the input dimension is 2, and the dimension of the hidden feature space $d_{f}$ is 16.
 The number of multi-head attention modules is 16, and the number of layers of encoder and decoder is 6.
-The model was trained on an NVIDIA GeForce RTX 2080 GPU and NVIDIA A100 Tensor Core GPU by Adam optimizer [44] for 50 epochs, the learning rate is 0.001, and the batch size is 64.
+The model was trained on an NVIDIA GeForce RTX 2080 GPU and NVIDIA A100 Tensor Core GPU using the Adam optimizer [53] for 50 epochs, the learning rate is 0.001, and the batch size is 64.
 4.4 Trading strategy
-A trading strategy is used for the model and Algorithm (1) shows the algorithm of strategy.
-Before the first trade date of the timestamp $t$, all sequence $\chi^{t}_{i,k}$ from stock set $S^{t}$ will be put in the model and obtain the list of the output.
-Then the stocks will be ranked according to the first element of the output and the first $\frac{1}{n}\%$ stocks will be added to the stock pool.
-If the stock has already been in the stock pool on the last timestamp, it will keep holding; if the stock is in the predicted pool but not in the previous pool, it will be bought in with the same proportion of the whole account.
+A trading strategy is used for the model and Algorithm 1 shows the pseudocode of the strategy.
+Before the first trade date of the timestamp $t$, all sequence $\chi^{t}_{i,k}$ from the stock set $S^{t}$ will be put in the model and obtain the list of outputs.
+Then, the stocks will be ranked according to the first element of the output and the first $\frac{1}{q}\%$ stocks will be added to the stock pool.
+If the stock already was in the stock pool on the last timestamp, it will be held; if the stock is in the predicted pool but not in the previous pool, it will be bought in with the same proportion of the whole account.
 Stocks that are not in the predicted pool will be sold out.
-The same method will run repeated during the rest period.
+The same method is run repeatedly during the subsequent periods.
 The backtest starts from January 2020, in other words, the result of the sequences from May 2018 to December 2019 will be used as the first stock pool to trade.
@@
 Formally, the Sharpe Ratio (SR) and the $\alpha$ ratio will be used to test the performance of the strategy.
-The SR [3] is a measure of risk-adjusted return which can describe the additional earnings an investor receives per unit of increase, which is shown in Eq.(8), where $R_{p}$ is the return of the portfolio, $R_{f}$ is the risk-free rate:
-$SR=\frac{E[R_{p}]-R_{f}}{std[R_{p}]}$ (8)
+The SR [3] is a measure of risk-adjusted return that describes the additional earnings an investor receives for each standard deviation unit increase, which is shown in equation (9), where $R_{p}$ is the return of the portfolio and $R_{f}$ is the risk-free rate is the rate of London Interbank Offered Rate (LIBOR) which calculated averaged from estimates submitted by banks in London.
+The risk-free rate in the backtest will be calculated under the average LIBOR rates during the test period.
+$\mathrm{SR}=\frac{\mathbb{E}[R_{p}]-R_{f}}{\mathrm{std}[R_{p}]}$ (9)
 Alpha represents the excess return of an investment portfolio relative to its benchmark.
@@
 Specifically, alpha is the excess return of the actual portfolio return over its expected theoretical return.
-Eq.(9) shows the calculation method of alpha rate, where $E(R_{m})$ is the market return and $\beta$ is the correlation between the portfolio and systemic risk.
+Equation (10) shows the calculation method of the alpha rate, where $E(R_{m})$ is the market return and $\beta$ is the correlation between the portfolio and systemic risk.
 $Alpha=(R_{p}-R_{f})-\beta(E[R_{m}]-R_{f})$ (10)
@@
 The win rate is the percentage of trade days where the portfolio encounters a positive return overall trade days.
-5 Result and discussion
+$VaR=\mu+\sigma N^{-1}(X)$ (11)
+[…]
+5 Results and discussion
 Building upon the analysis of individual strategies, this section presents an evaluation and discussion of the experimental outcomes.
-The discussion focuses on the comparative effectiveness of monthly, weekly, and daily strategies.
+The discussion focuses on the comparative effectiveness of different timestamps (monthly, weekly, and daily) and different training scales.
 The results underscore the nuanced relationship between trading frequency and market performance, offering valuable insights for the development of robust quantitative trading strategies.
 5.1 Overall performance
-With the starting amount of RMB 5,000,000, the results of different strategies are obtained.
 The result of the experiment is shown in Table 2.
@@
 The Sharpe Ratio and Alpha, vary across the strategies, with some periods showing negative values, suggesting underperformance relative to the risk taken, while others show positive values.
+In the comparison of 99% VaR, the strategies in monthly frequency perform better than others, where Month_1 and Month_3 are lower than 3%, which means the portfolio maximum loss 2.81% and 2.3% of return a day in 99% confidence.
 Table 2: Result of experiments
-Strategy AR AER TR WR SR Alpha
-Month_1 17.35% 19.43% 26.09% 57.8% 0.915 0.162
-Month_2 9.91% 13.86% 51.69% 49.3% 0.289 0.102
-Month_3 7.37% 9.91% 32.33% 51.6% 0.246 0.064
-Week_1 -0.83% 1.31% 7.13% 46.4% -0.236 -0.03
-Week_2 7.49% 10.81% 1.18% 49.2% 0.160 0.085
-Week_3 12.3% 12.73% 1.39% 54.4% 0.372 0.116
-Day_1 7.89% 11.4% 6.71% 43.1% 0.181 0.090
-Day_2 10.23% 10.94% 6.51% 44.4% 0.279 0.097
-Day_3 9.81% 10.03% 5.57% 44.6% 0.281 0.092
-CSI300 1.77% \ \ \ \ \
+Strategy AR AER TR WR SR Alpha VaR
+Month_1 17.35% 19.43% 26.09% 57.8% 0.915 0.162 2.81
+Month_2  9.91% 13.86% 51.69% 49.3% 0.289 0.102 3.61
+Month_3  7.37%  9.91% 32.33% 51.6% 0.246 0.064 2.3
+Week_1 -0.83%  1.31% 7.13% 46.4% -0.236 -0.030 3.05
+Week_2  7.49% 10.81% 1.18% 49.2% 0.160 0.085 3.73
+Week_3 12.3% 12.73% 1.39% 54.4% 0.372 0.116 3.77
+Day_1  7.89% 11.4% 6.71% 43.1% 0.181 0.090 3.92
+Day_2 10.23% 10.94% 6.51% 44.4% 0.279 0.097 3.91
+Day_3  9.81% 10.03% 5.57% 44.6% 0.281 0.092 4.02
+CSI300 1.77% \ \ \ -0.015 \ 3.19%
 In the weekly strategy results, Week_3 showed a remarkable annual return of 12.3% and an annual excess return of 12.73%, while Week_1 displayed a negative annual return.
@@
 The Sharpe Ratio and Alpha for weekly and daily strategies show a mix of positive and negative values, reflecting the fluctuating nature of shorter-term trading efficacy.
-To demonstrate the advantages of the Transformer factor over traditional factors, 100 price-volume type factors from JoinQuant  [^3] https://www.joinquant.com/ (one of Chinese largest financial quantitative platforms) were selected for backtesting under the same trading strategy (Algorithm 1).
-Price-volume factors were used because these are calculated based on stocks’ prices, volumes, and turnover rates, which are similar to the training data used by Transformer factors (also used stock turnover rates and price data).
+To demonstrate the advantages of the quantformer factor over traditional factors, 100 price-volume type factors from JoinQuant  [^3] https://www.joinquant.com/ (one of China’s largest financial quantitative platforms) were selected for backtesting under the same trading strategy (Algorithm 1).
+Price-volume factors were used because these are calculated based on stocks’ prices, volumes, and turnover rates, which are similar to the training data used by quantformer factors (also used stock turnover rates and price data).
 The detailed performance of factors and benchmark is shown in Appendix A.
@@
 Figure 5: Backtest results based on different factors
-Among these factor-based strategies, the Transformer factor performed well.
-Except for the win rate, which was slightly lower than a few strategies, it excelled in terms of annual return, annual excess return, $\sigma_{p}$, and Sharpe Ratio, ranking the best among the 101 factors.
-Figure 4 illustrates the return curves of the Transformer (Trans_Month_1) alongside the benchmark and some traditional factors.
-”25_week_rank” represents the current price’s position over the past 25 weeks; ”EMAC20” and ”MAC20” are the 20-day index moving average and the stock’s 20-day moving average respectively; ”DAVOL20” and ”VOL20” represent the ratio of the 20-day average turnover rate to the 120-day average turnover rate, and the mean of the stock’s 20-day turnover rate; ”ROC20” is the price rate of change over 20 days; ”MAWVAD” is calculated as the product of the difference between the closing price and the opening price, divided by the range of the highest and lowest prices, all multiplied by the volume, accumulated over six days; ”Variance20” is the variance of the stock’s 20-day annualized returns.
-Most of these factors are compared based on a window of 20 timestamps, as the Transformer factor was also trained on 20 timestamps.
-In comparison with these factors, the Transformer factor (blue line) demonstrated a significantly better performance in terms of returns than the benchmark (orange line) and other factor strategies, expressing the improvement of the Transformer factor over traditional price-volume factors.
+Among these factor-based strategies, the quantformer factor performed well.
+Except for the win rate, which was slightly lower than a few strategies, it excelled in terms of annual return, annual excess return, $\sigma_{p}$, Sharpe Ratio and sortino ratio ranking the best among the 101 factors.
+The 99% VaR of quantformer factor is in the quantile of the first 10% among all factors.
+Figure 5 illustrates the return curves of the quantformer (QF_Month_1) alongside the benchmark and some traditional factors.
+“25_week_rank” represents the current price’s position over the past 25 weeks; “EMAC20” and “MAC20” are the 20-day index moving average and the stock’s 20-day moving average respectively; “DAVOL20” and “VOL20” represent the ratio of the 20-day average turnover rate to the 120-day average turnover rate, and the mean of the stock’s 20-day turnover rate; “ROC20” is the price rate of change over 20 days; “MAWVAD” is calculated as the product of the difference between the closing price and the opening price, divided by the range of the highest and lowest prices, all multiplied by the volume, accumulated over six days; “Variance20” is the variance of the stock’s 20-day annualized returns.
+Most of these factors are compared based on a window of 20 timestamps, as the quantformer factor was also trained on 20 timestamps.
+In comparison with these factors, the quantformer factor (blue line) demonstrated a significantly better performance in terms of returns than the benchmark (orange line) and other factor strategies, expressing the improvement of the quantformer factor over traditional price-volume factors.
 5.2 Half-life period test
@@
 For instance, in the first strategy, the trade will start in January 2020; in the second, it will start in February 2020; in the second, it will start in March 2020.
-For the period from January 2020 to February 2022, the strategy achieved a total return of 65.29% and an annual return of 17.87%.
-The excess return over this period was 66.39%, indicating a strong performance above a benchmark or risk-free rate.
+For the period from January 2020 to February 2022, the strategy achieved a total return (TotalR) of 65.29% and an annual return of 17.87%.
+The total excess return (TER) over this period was 66.39%, indicating a strong performance above a benchmark or risk-free rate.
 The Sharpe Ratio for this period was 0.863, which is a respectable figure that suggests a favorable risk-adjusted return.
 Additionally, the strategy generated an alpha of 0.164, demonstrating its ability to outperform the expected return on a risk-adjusted basis.
-Table 3: Result of the same monthly strategy with different start month
-Strategy Total return Annual return Excess return Sharpe ratio Alpha
-2020/01-2022/02 65.29% 17.87% 66.39% 0.863 0.164
-2020/02-2022/03 67.52% 18.21% 65.58% 0.940 0.161
-2020/03-2022/04 56.41% 15.63% 52.96% 0.739 0.135
+Table 3: Result of the same strategy with different starting month
+Strategy TotalR AR TER SR Alpha CSI300
+2020/01-2022/02 65.29% 17.87% 66.39% 0.863 0.164 -0.66%
+2020/02-2022/03 67.52% 18.21% 65.58% 0.940 0.161 1.17%
+2020/03-2022/04 56.41% 15.63% 52.96% 0.739 0.135 2.26%
 In the subsequent period from February 2020 to March 2023, the strategy yielded a slightly higher total return of 67.52% and an annual return of 18.21%.
@@
 The alpha reduced to 0.135, which, while lower than the other periods, still indicates that the strategy managed to provide returns above those predicted by its risk profile.
-5.3 Discussion
 Comparing the results between monthly, weekly, and daily strategies provides further insights.
@@
 This indicates weekly data strikes a balance between filtering noise and responding swiftly to emerging patterns.
-The daily frequency strategies perform lower annual returns compared to others.
-This highlights the challenge of making profitable trades at higher frequencies - noise and volatility make reliable signals more difficult.
+The daily frequency strategies perform lower annual returns and higher risk compared to others This highlights the challenge of making profitable trades at higher frequencies - noise and volatility make reliable signals more difficult.
 The distribution of returns also shows increased dispersion at the daily level.
@@
 There is likely an optimal point between 1 week to 1 month where matches market regularly.
+5.3 Training under different scales
+Understanding data volume and data frequency is critical to designing infrastructure for processing data and model stability [55].
+To test the ability of stock selection of the model, three other factors ”QF_10%”, ”QF_5%”, ”QF_1%” are trained and tested (with ”QF_Month_1%, named as ”QF_20%”).
+The four groups are all trained under $d_{f}$ = 3, which include the first, middle, and tail 20%, 10%, 5%, and 1% of stocks in ranking, respectively.
+The backtest period and other settings are the same with 4.
+Table 4: Result of factors under different scales
+Strategy AR AER $\sigma_{p}$ WR SR Alpha VaR
+QF_2017.35% 19.43% 0.162 57.8% 0.915 0.162 2.81%
+QF_1013.12% 17.73 0.159 61.8% 0.574 0.128 2.023%
+QF_512.59% 16.02% 0.136 61% 0.63 0.117 2.015%
+QF_124.71% 35.74% 0.214 53.3% 0.967 0.249 3.048%
+CSI300 1.77% \ 0.197 \ -0.015 \ 3.19%
+Table 4 shows the result of the backtest.
+QF_10strategy recorded an AR of 13.12% and an AER of 17.73%, with a slightly lower portfolio volatility of 0.159.
+This strategy had the highest WR at 61.8%.
+QF_5strategy delivered 12.59% AR and 16.02% AW, with the lowest portfolio volatility at 0.136.
+It had the lowest VaR at 2.015%.
+QF_1strategy outperformed the others, achieving the highest AR of 24.71% and the highest AER of 35.74%, though it had the highest volatility at 0.214.
+These four factors all perform better than CSI300 based on their return and risk.
+The result shows that the model takes the ability to select stocks under different training set and requirements.
 6 Conclusion
-This study proposed an approach to exploit transformer models for quantitative stock prediction and trading.
+This study proposed a new model, quantformer for quantitative stock prediction and trading.
 We addressed the need for handling numerical input data rather than text and adapting the model for forecasting tasks rather than sequence-to-sequence problems common in NLP.
@@
 There is further scope to enhance the framework, for instance by incorporating additional signals like news and fundamentals as auxiliary inputs.
-The self-attention mechanism could likely be improved to encode greater temporal context.
-Or use transformer-based models such as GPT-4 [45] and Claude3 [46] to establish the strategy as these models are fine-tuned and may perform better in quantitative finance tasks.
-Overall, our work illustrates the viability of transformers for financial data modeling, provides a flexible framework to model the market from multiple perspectives of financial time series and develops profitable trading strategies that consider different variables.
+The self-attention mechanism could likely be improved to encode greater temporal context, or use transformer-based models such as GPT-4 [56] and Claude3 [57] to establish the strategy as these models are fine-tuned and may perform better for quantitative finance tasks.
+Overall, our work illustrates the viability of quantformer for financial data modeling, provides a flexible framework to model the market from multiple perspectives of financial time series and develops profitable trading strategies that consider different variables.
 With further research to solve the above limitations, we believe such ML-based quantitative methods hold rich potential.
 The framework provides a flexible foundation for better understanding markets and developing profitable trading strategies.
+The implementation code of quantformer is available at https://github.com/zhangmordred/QuantFormer.
 7 Acknowledgements and declaration
+This work was funded by the Natural Science Foundation of China (12271047); Guangdong Provincial Key Laboratory of Interdisciplinary Research and Application for Data Science, BNU-HKBU United International College (2022B1212010006); UIC research grant (R0400001-22; UICR0400008-21; R72021114); Guangdong College Enhancement and Innovation Program (2021ZDZX1046).
 The authors have no competing interests to declare that are relevant to the content of this article.
-Appendix A Detailed backtest results between 100 factors and the transformer factor
+Appendix A Detailed backtest results between 100 factors and the quantformer factor
 All the data of factors comes from JoinQuant [^4] https://www.joinquant.com/, which is one of the most common quantitative finance platforms.
-Detailed descriptions of the following factors are at JoinQuant Factors.
-Seven metrics are used to evaluate the performance of each factor:
-  • [label=•]
+Detailed descriptions of the following factors are available at JoinQuant Factors.
+Ten metrics are used to evaluate the performance of each factor:
   • Annual Return (AR) and Annual Excess Return (AER) are two metrics to reflect the return of the portfolio directly.
@@
   • Win Rate (WR) is the metric to evaluate the fraction between the action that makes a profit and all the actions.
-  • Sharpe Ratio (SR) is a measure of risk-adjusted return that can describe the additional earnings an investor receives per unit of increase, which is shown in Eq.(8).
-  • The Alpha measures the portfolio in generating returns above the market benchmark.
-A positive alpha indicates that the investment portfolio has achieved a higher return than its benchmark after risk adjustment.
+  • Sharpe Ratio (SR) is a measure of risk-adjusted return that describes the additional earnings an investor receives for each standard deviation unit increase, which is shown in equation (9).
+  • The Alpha measures the ability of a portfolio to generate returns above the market benchmark.
+A positive alpha indicates that the investment portfolio has achieved a higher return than its benchmark after risk adjustment, which is shown in equation (10).
   • Beta is the correlation between the portfolio and systemic risk, reflecting the sensitivity of the strategy to the change in the market.
-  • Max Drawdown (MD) means the potential worst-case scenario or the most extreme possible loss.
-  • Portfolio volatility ($\sigma_{p}$) is used to measure the risk of a strategy, where is positively correlated to the risk of a strategy.
-Factor AR (%) AER (%) WR (%) SR Alpha Beta MD (%) $\sigma_{p}$
-ARBR -1.47 0.32 54.7 -0.251 -0.002 0.914 34.46 0.221
-AR -5.87 -4.37 53.7 -0.406 -0.048 0.981 47.84 0.260
-ATR14 -2.12 -0.37 62.7 -0.250 0.006 1.160 46.55 0.250
-ATR6 -3.94 -2.31 61.2 -0.328 -0.014 1.175 50.11 0.253
-BBIC 1.17 3.14 50.9 -0.109 0.033 1.038 32.64 0.258
-BR -6.91 -5.47 53.6 -0.467 -0.064 0.934 51.72 0.254
-CCI10 -18.65 -17.99 43.3 -1.402 -0.261 0.812 72.04 0.220
-CCI15 -17.70 -16.98 44.2 -1.293 -0.240 0.832 71.43 0.223
-CCI20 -15.52 -14.66 46.3 -1.051 -0.196 0.868 69.12 0.235
-CR20 -4.27 -2.67 54.8 -0.329 -0.032 0.927 49.92 0.263
-DAVOL10 -7.44 -6.04 54.1 -0.504 -0.071 0.920 45.45 0.248
-DAVOL20 -10.95 -9.78 54.4 -0.702 -0.119 0.919 48.96 0.246
-DAVOL5 -9.54 -8.28 51.8 -0.612 -0.097 0.950 46.41 0.250
-EMA5 2.77 4.85 48.3 -0.047 0.049 1.041 30.70 0.259
-EMAC10 0.27 2.17 50.6 -0.144 0.024 1.037 31.42 0.258
-EMAC12 1.58 3.57 50.6 -0.093 0.037 1.032 31.76 0.258
-EMAC20 6.01 8.29 54.1 0.069 0.079 1.044 32.43 0.260
-EMAC26 3.83 5.97 54.1 -0.008 0.059 1.046 34.87 0.261
-EMAC120 1.74 3.75 48.3 -0.088 0.034 0.960 35.81 0.254
-Kurtosis20 -2.96 -1.27 52.5 -0.395 -0.027 0.772 27.71 0.182
-Kurtosis60 -0.28 1.59 53.9 -0.240 -0.001 0.716 22.07 0.179
-Kurtosis120 -4.12 -2.51 49.0 -0.468 -0.043 0.708 28.28 0.182
-MAC5 0.54 2.46 46.9 -0.134 0.026 1.035 30.03 0.257
-MAC10 -1.20 0.61 49.6 -0.206 0.008 1.039 34.15 0.256
-MAC20 2.47 4.52 53.5 -0.059 0.046 1.043 34.58 0.257
-MAC60 11.10 13.72 53.8 0.243 0.120 0.987 28.61 0.254
-MAC120 3.11 5.21 47.7 -0.035 0.049 0.987 35.40 0.259
-MACDC -11.12 -9.97 52.4 -0.674 -0.117 1.003 46.48 0.261
-MASS -10.55 -9.35 51.0 -0.673 -0.115 0.886 48.34 0.249
-MAWVAD 7.29 9.67 55.9 0.119 0.082 0.891 36.54 0.248
-MFI14 -6.21 -4.73 51.5 -0.456 -0.056 0.918 45.65 0.241
-PLRC6 -14.83 -13.92 47.8 -0.905 -0.178 0.960 66.24 0.259
-PLRC12 -7.28 -5.87 53.3 -0.452 -0.064 1.012 52.30 0.273
-PLRC24 -5.77 -4.26 58.1 -0.361 -0.044 1.034 53.28 0.289
-PSY -9.11 -7.83 50.6 -0.660 -0.096 0.880 49.79 0.223
-Price1M -13.55 -12.55 53.0 -0.760 -0.156 0.977 64.85 0.280
-Price3M -6.69 -5.25 58.4 -0.380 -0.056 1.023 59.14 0.305
-Price1Y 1.55 3.54 66.8 -0.075 0.041 1.116 60.25 0.326
-ROC6 -16.13 -15.30 47.4 -0.982 -0.200 0.978 68.14 0.262
-ROC20 -11.63 -10.51 54.7 -0.628 -0.123 1.014 60.28 0.291
-ROC60 -6.86 -5.43 59.6 -0.382 -0.056 1.062 62.16 0.309
-ROC120 2.36 4.41 66.9 -0.051 0.049 1.117 56.36 0.319
-Skewness20 -5.91 -4.41 50.8 -0.538 -0.058 0.820 40.65 0.197
-Skewness60 -6.47 -5.01 51.9 -0.595 -0.068 0.765 38.02 0.190
-Skewness120 -4.37 -2.77 51.1 -0.459 0.044 0.743 36.01 0.192
-TRIX5 -10.72 -9.54 52.5 -0.607 -0.110 1.009 56.07 0.280
-TRIX10 -3.51 -1.85 58.4 -0.268 -0.019 1.008 52.92 0.292
-TVMA20 -1.92 -0.16 64.9 -0.209 0.014 1.268 44.66 0.288
-TVMA6 -0.75 1.09 62.4 -0.166 0.026 1.259 39.46 0.288
-TVSTD20 -1.83 -0.06 58.5 -0.214 0.012 1.218 35.89 0.277
-TVSTD6 -3.77 -2.13 53.4 -0.297 -0.010 1.202 46.57 0.272
-VDEA -3.05 -1.37 57.6 -0.305 -0.023 0.857 37.01 0.239
-VDIFF -7.93 -6.57 0.2 -0.533 -0.079 0.895 39.79 0.247
-VEMA5 -0.41 1.46 57.9 -0.186 0.006 0.850 37.31 0.237
-VEMA10 0.95 2.91 58.9 -0.128 0.019 0.844 35.51 0.236
-VEMA12 1.01 2.96 59.3 -0.126 0.020 0.840 35.17 0.235
-VEMA26 2.41 4.46 59.3 -0.068 0.033 0.838 31.03 0.231
-VMACD -1.03 0.79 50.5 -0.216 0.001 0.875 33.96 0.235
-VOL5 -8.12 -6.76 55.9 -0.452 -0.066 1.155 53.86 0.296
-VOL10 -7.07 -5.64 58.3 -0.404 -0.053 1.145 50.65 0.298
-VOL20 -5.43 -3.89 58.6 -0.337 -0.032 1.156 50.19 0.298
-VOL60 0.42 2.34 58.1 -0.125 0.032 1.145 38.35 0.286
-VOL120 -1.88 -0.12 53.0 -0.213 0.007 1.136 42.09 0.281
-VOL240 -5.51 -3.98 47.0 -0.382 -0.038 1.075 43.53 0.265
-VOSC -6.35 -4.88 51.2 -0.453 -0.056 0.944 40.40 0.246
-VROC6 -16.30 -15.49 42.7 -1.179 -0.208 0.897 69.03 0.221
-VROC12 -11.47 -10.34 46.7 -0.793 -0.127 0.915 51.62 0.228
-VR -8.43 -7.10 52.8 -0.566 -0.083 0.932 48.57 0.244
-VSTD10 -0.77 1.07 56.3 -0.205 0.001 0.843 37.75 0.234
-VSTD20 -1.60 0.19 53.7 -0.245 -0.007 0.844 37.31 0.232
-Variance20 -1.94 -0.18 59.9 -0.201 0.005 1.124 46.38 0.302
-Variance60 9.96 12.51 65.2 0.168 0.122 1.178 30.63 0.313
-Variance120 5.40 7.65 60.8 0.040 0.083 1.193 39.54 0.312
-Volume1M -15.54 -14.68 20.7 -0.867 -0.189 0.991 63.68 0.285
-WVAD -0.71 1.13 52.3 -0.199 0.002 0.836 42.38 0.238
-arron_down_25
--7.35 -5.94 48.3 -0.593 -0.072 0.880 42.63 0.209
-arron_up_25 -15.51 -14.65 46.4 -1.049 -0.195 0.870 66.73 0.235
-bear_power -13.31 -12.30 50.8 -0.800 -0.154 0.934 59.58 0.261
-beta 10.30 12.87 62.4 0.168 0.134 1.344 35.96 0.328
-boll-up 2.19 4.23 54.5 -0.068 0.044 1.062 33.09 0.265
-book_to_price 1.87 3.86 58.5 -0.117 0.031 0.889 26.73 0.181
-bull_power -14.81 -13.91 51.1 -0.844 -0.176 0.984 65.99 0.278
-earnings_yield -2.41 -0.68 52.6 -0.348 -0.029 0.627 25.87 0.188
-growth 2.90 4.99 59.9 -0.045 0.052 1.068 37.35 0.244
-leverage 2.05 4.07 53.1 -0.089 0.023 0.719 22.57 0.217
-liquidity -2.71 -1.00 56.7 -0.243 -0.002 1.137 41.28 0.284
-momentum 10.93 13.54 65.8 0.195 0.129 1.159 43.19 0.310
-money_flow_20 -1.72 0.05 65.5 -0.202 0.016 1.269 44.10 0.288
-price_no_fq 6.83 9.17 76.1 0.095 0.092 1.132 44.86 0.267
-pull-down -2.28 -0.54 47.6 -0.317 -0.015 0.831 34.46 0.203
-residual_volatility
-4.13 6.30 62.3 0.003 0.062 1.048 41.34 0.269
-sharpe_ratio_20
--11.88 -10.78 54.0 -0.663 -0.129 0.979 57.66 0.282
-sharpe_ratio_60
--7.42 -6.02 58.8 -0.417 -0.065 1.029 65.19 0.300
-sharpe_ratio_120
--0.32 1.55 65.7 -0.143 0.019 1.062 60.52 0.303
-size -6.36 -4.89 51.8 -0.571 -0.058 0.906 36.00 0.196
-turnover_volatility
--10.80 -9.62 51.8 -0.624 -0.108 1.075 57.05 0.274
-1day_VPT -16.45 -15.65 44.8 -1.062 -0.211 0.904 70.43 0.248
-1day_VPT_6
--6.36 -4.89 53.0 -0.416 -0.053 0.995 47.06 0.269
-1day_VPT_12
--0.07 1.82 57.9 -0.146 0.020 1.033 46.51 0.278
-25week_close
--11.51 -10.39 51.0 -0.859 -0.133 0.817 47.79 0.211
-Trans_Month_1
-17.35 19.43 57.8 0.915 0.162 0.588 18.35 0.161
-Average -3.78 -2.15 53.71 -0.36 -0.03 0.98 44.88 0.26
-CSI300 1.77 \ \ \ \ \ \ 0.197
+For the daily return of the strategy, $D_{p}$, and the daily return of the benchmark $D_{m}$, the equation of beta is:
+    $Beta=\beta_{p}=\frac{Cov(D_{p},D_{m})}{Var(D_{m})}$
+    where $Cov(D_{p},D_{m})$ is the correlation of between $D_{p}$ and $D_{m}$, $Var(D_{m})$ is the variance of daily return of benchmark.
+  • Max Drawdown (MD) means the potential worst-case scenario or the most extreme possible loss from the previous peak.
+For the trough value of a portfolio $V_{trough}$ and the peak value of the portfolio $V_{peak}$, the MDD is calculated in the following way:
+    $MDD=\frac{V_{peak}-V_{trough}}{V_{peak}}$
+  • Portfolio volatility ($\sigma_{p}$) is the standard deviation of the portfolio returns.
+$\sigma_{p}$ avoids the problem of negative deviations canceling with positive deviations and also penalizes larger deviations from the mean.
+It provides a kind of weighted average deviation in which large deviations carry more weight [58].
+  • Sortino Ratio (STN) differentiates harmful volatility from total overall volatility by using the standard deviation of negative portfolio returns.
+This metric measures the performance of the investment relative to the downward risk.
+[…]
+Factor AR (%) AER (%) WR (%) SR Alpha Beta MD (%) $\sigma_{p}$ VaR (%) STN
+ARBR -1.47 0.32 54.7 -0.251 -0.002 0.914 34.46 0.221 3.74 -0.375
+AR -5.87 -4.37 53.7 -0.406 -0.048 0.981 47.84 0.260 3.94 0.612
+ATR14 -2.12 -0.37 62.7 -0.250 0.006 1.160 46.55 0.250 4.51 -0.35
+ATR6 -3.94 -2.31 61.2 -0.328 -0.014 1.175 50.11 0.253 4.48 -0.455
+BBIC 1.17 3.14 50.9 -0.109 0.033 1.038 32.64 0.258 4.16 -0.155
+BR -6.91 -5.47 53.6 -0.467 -0.064 0.934 51.72 0.254 3.80 -0.718
+CCI10 -18.65 -17.99 43.3 -1.402 -0.261 0.812 72.04 0.220 2.79 -2.026
+CCI15 -17.70 -16.98 44.2 -1.293 -0.240 0.832 71.43 0.223 2.90 -1.866
+CCI20 -15.52 -14.66 46.3 -1.051 -0.196 0.868 69.12 0.235 3.28 -1.519
+CR20 -4.27 -2.67 54.8 -0.329 -0.032 0.927 49.92 0.263 4.26 -0.511
+DAVOL10 -7.44 -6.04 54.1 -0.504 -0.071 0.920 45.45 0.248 3.38 -0.812
+DAVOL20 -10.95 -9.78 54.4 -0.702 -0.119 0.919 48.96 0.246 2.86 -1.079
+DAVOL5 -9.54 -8.28 51.8 -0.612 -0.097 0.950 46.41 0.250 3.17 -1.019
+EMA5 2.77 4.85 48.3 -0.047 0.049 1.041 30.70 0.259 4.66 -0.068
+EMAC10 0.27 2.17 50.6 -0.144 0.024 1.037 31.42 0.258 4.19 -0.203
+EMAC12 1.58 3.57 50.6 -0.093 0.037 1.032 31.76 0.258 4.17 -0.131
+EMAC20 6.01 8.29 54.1 0.069 0.079 1.044 32.43 0.260 4.32 0.102
+EMAC26 3.83 5.97 54.1 -0.008 0.059 1.046 34.87 0.261 4.29 -0.012
+EMAC120 1.74 3.75 48.3 -0.088 0.034 0.960 35.81 0.254 3.77 -0.145
+Kurtosis20 -2.96 -1.27 52.5 -0.395 -0.027 0.772 27.71 0.182 2.74 -0.6
+Kurtosis60 -0.28 1.59 53.9 -0.240 -0.001 0.716 22.07 0.179 2.64 -0.367
+Kurtosis120 -4.12 -2.51 49.0 -0.468 -0.043 0.708 28.28 0.182 2.44 -0.684
+MAC5 0.54 2.46 46.9 -0.134 0.026 1.035 30.03 0.257 4.29 -0.195
+MAC10 -1.20 0.61 49.6 -0.206 0.008 1.039 34.15 0.256 4.19 -0.286
+MAC20 2.47 4.52 53.5 -0.059 0.046 1.043 34.58 0.257 4.09 -0.083
+MAC60 11.10 13.72 53.8 0.243 0.120 0.987 28.61 0.254 4.34 0.399
+MAC120 3.11 5.21 47.7 -0.035 0.049 0.987 35.40 0.259 3.99 -0.058
+MACDC -11.12 -9.97 52.4 -0.674 -0.117 1.003 46.48 0.261 3.54 -1.046
+MASS -10.55 -9.35 51.0 -0.673 -0.115 0.886 48.34 0.249 6.19 -1.032
+MAWVAD 7.29 9.67 55.9 0.119 0.082 0.891 36.54 0.248 4.19 0.184
+MFI14 -6.21 -4.73 51.5 -0.456 -0.056 0.918 45.65 0.241 3.61 -0.699
+PLRC6 -14.83 -13.92 47.8 -0.905 -0.178 0.960 66.24 0.259 3.85 -1.368
+PLRC12 -7.28 -5.87 53.3 -0.452 -0.064 1.012 52.30 0.273 4.20 -0.682
+PLRC24 -5.77 -4.26 58.1 -0.361 -0.044 1.034 53.28 0.289 4.88 -0.549
+PSY -9.11 -7.83 50.6 -0.660 -0.096 0.880 49.79 0.223 3.36 -0.992
+Price1M -13.55 -12.55 53.0 -0.760 -0.156 0.977 64.85 0.280 4.21 -1.145
+Price3M -6.69 -5.25 58.4 -0.380 -0.056 1.023 59.14 0.305 5.44 -0.555
+Price1Y 1.55 3.54 66.8 -0.075 0.041 1.116 60.25 0.326 7.28 -0.105
+ROC6 -16.13 -15.30 47.4 -0.982 -0.200 0.978 68.14 0.262 3.56 -1.741
+ROC20 -11.63 -10.51 54.7 -0.628 -0.123 1.014 60.28 0.291 4.51 -0.971
+ROC60 -6.86 -5.43 59.6 -0.382 -0.056 1.062 62.16 0.309 5.28 -0.529
+ROC120 2.36 4.41 66.9 -0.051 0.049 1.117 56.36 0.319 6.79 -0.071
+Skewness20 -5.91 -4.41 50.8 -0.538 -0.058 0.820 40.65 0.197 3.04 0.839
+Skewness60 -6.47 -5.01 51.9 -0.595 -0.068 0.765 38.02 0.190 2.58 -0.917
+Skewness120 -4.37 -2.77 51.1 -0.459 0.044 0.743 36.01 0.192 2.69 -0.684
+TRIX5 -10.72 -9.54 52.5 -0.607 -0.110 1.009 56.07 0.280 4.16 -0.923
+TRIX10 -3.51 -1.85 58.4 -0.268 -0.019 1.008 52.92 0.292 5.14 -0.4
+TVMA20 -1.92 -0.16 64.9 -0.209 0.014 1.268 44.66 0.288 4.63 -0.31
+TVMA6 -0.75 1.09 62.4 -0.166 0.026 1.259 39.46 0.288 4.76 -0.246
+TVSTD20 -1.83 -0.06 58.5 -0.214 0.012 1.218 35.89 0.277 4.03 -0.318
+TVSTD6 -3.77 -2.13 53.4 -0.297 -0.010 1.202 46.57 0.272 4.40 -0.449
+VDEA -3.05 -1.37 57.6 -0.305 -0.023 0.857 37.01 0.239 3.06 -0.484
+VDIFF -7.93 -6.57 0.2 -0.533 -0.079 0.895 39.79 0.247 3.08 -0.842
+VEMA5 -0.41 1.46 57.9 -0.186 0.006 0.850 37.31 0.237 3.33 -0.277
+VEMA10 0.95 2.91 58.9 -0.128 0.019 0.844 35.51 0.236 3.36 -0.19
+VEMA12 1.01 2.96 59.3 -0.126 0.020 0.840 35.17 0.235 3.35 -0.186
+VEMA26 2.41 4.46 59.3 -0.068 0.033 0.838 31.03 0.231 3.37 -0.1
+VMACD -1.03 0.79 50.5 -0.216 0.001 0.875 33.96 0.235 3.38 -0.345
+VOL5 -8.12 -6.76 55.9 -0.452 -0.066 1.155 53.86 0.296 4.48 -0.656
+VOL10 -7.07 -5.64 58.3 -0.404 -0.053 1.145 50.65 0.298 4.35 -0.59
+VOL20 -5.43 -3.89 58.6 -0.337 -0.032 1.156 50.19 0.298 4.33 -0.481
+VOL60 0.42 2.34 58.1 -0.125 0.032 1.145 38.35 0.286 4.71 -0.175
+VOL120 -1.88 -0.12 53.0 -0.213 0.007 1.136 42.09 0.281 4.64 -0.3
+VOL240 -5.51 -3.98 47.0 -0.382 -0.038 1.075 43.53 0.265 3.99 -0.541
+VOSC -6.35 -4.88 51.2 -0.453 -0.056 0.944 40.40 0.246 3.44 -0.739
+VROC6 -16.30 -15.49 42.7 -1.179 -0.208 0.897 69.03 0.221 2.91 -1.778
+VROC12 -11.47 -10.34 46.7 -0.793 -0.127 0.915 51.62 0.228 2.74 -1.241
+VR -8.43 -7.10 52.8 -0.566 -0.083 0.932 48.57 0.244 3.72 -0.834
+VSTD10 -0.77 1.07 56.3 -0.205 0.001 0.843 37.75 0.234 3.19 -0.309
+VSTD20 -1.60 0.19 53.7 -0.245 -0.007 0.844 37.31 0.232 3.01 -0.366
+Variance20 -1.94 -0.18 59.9 -0.201 0.005 1.124 46.38 0.302 4.84 -0.301
+Variance60 9.96 12.51 65.2 0.168 0.122 1.178 30.63 0.313 5.91 0.241
+Variance120 5.40 7.65 60.8 0.040 0.083 1.193 39.54 0.312 5.93 0.058
+Volume1M -15.54 -14.68 20.7 -0.867 -0.189 0.991 63.68 0.285 3.61 -1.339
+WVAD -0.71 1.13 52.3 -0.199 0.002 0.836 42.38 0.238 3.67 -0.317
+arron_down_25 -7.35 -5.94 48.3 -0.593 -0.072 0.880 42.63 0.209 2.84 -0.826
+arron_up_25 -15.51 -14.65 46.4 -1.049 -0.195 0.870 66.73 0.235 3.19 -1.604
+bear_power -13.31 -12.30 50.8 -0.800 -0.154 0.934 59.58 0.261 3.61 -1.182
+beta 10.30 12.87 62.4 0.168 0.134 1.344 35.96 0.328 6.76 0.252
+book_to_price 1.87 3.86 58.5 -0.117 0.031 0.889 26.73 0.181 2.91 -0.162
+bull_power -14.81 -13.91 51.1 -0.844 -0.176 0.984 65.99 0.278 3.83 -1.285
+earnings_yield -2.41 -0.68 52.6 -0.348 -0.029 0.627 25.87 0.188 2.55 -0.55
+growth 2.90 4.99 59.9 -0.045 0.052 1.068 37.35 0.244 4.50 -0.067
+leverage 2.05 4.07 53.1 -0.089 0.023 0.719 22.57 0.217 3.22 -0.144
+liquidity -2.71 -1.00 56.7 -0.243 -0.002 1.137 41.28 0.284 4.56 -0.35
+momentum 10.93 13.54 65.8 0.195 0.129 1.159 43.19 0.310 6.82 0.271
+money_flow_20 -1.72 0.05 65.5 -0.202 0.016 1.269 44.10 0.288 4.63 -0.298
+price_no_fq 6.83 9.17 76.1 0.095 0.092 1.132 44.86 0.267 6.20 0.131
+pull-up 2.19 4.23 54.5 -0.068 0.044 1.062 33.09 0.265 3.92 -0.093
+pull-down -2.28 -0.54 47.6 -0.317 -0.015 0.831 34.46 0.203 3.87 -0.439
+residual_volatility 4.13 6.30 62.3 0.003 0.062 1.048 41.34 0.269 5.25 0.004
+sharpe_ratio_20 -11.88 -10.78 54.0 -0.663 -0.129 0.979 57.66 0.282 4.09 -1.019
+sharpe_ratio_60 -7.42 -6.02 58.8 -0.417 -0.065 1.029 65.19 0.300 5.17 -0.583
+sharpe_ratio_120 -0.32 1.55 65.7 -0.143 0.019 1.062 60.52 0.303 6.10 -0.2
+size -6.36 -4.89 51.8 -0.571 -0.058 0.906 36.00 0.196 2.64 -0.82
+turnover_volatility -10.80 -9.62 51.8 -0.624 -0.108 1.075 57.05 0.274 3.70 -0.905
+1day_VPT -16.45 -15.65 44.8 -1.062 -0.211 0.904 70.43 0.248 3.09 -1.553
+1day_VPT_6 -6.36 -4.89 53.0 -0.416 -0.053 0.995 47.06 0.269 3.81 -0.82
+1day_VPT_12 -0.07 1.82 57.9 -0.146 0.020 1.033 46.51 0.278 4.39 -0.22
+25week_close -11.51 -10.39 51.0 -0.859 -0.133 0.817 47.79 0.211 2.44 -0.684
+QF_Month_1 17.35 19.43 57.8 0.915 0.162 0.588 18.35 0.161 2.81 0.946
+QF_Month_2 9.91 13.53 49.3 0.289 0.102 0.684 26.77 0.205 3.61 0.372
+QF_Month_3 7.37 10.4 51.6 0.246 0.064 0.483 17.33 0.137 2.3 0.309
+QF_Week_1 -0.83 1.41 46.4 -0.236 -0.003 0.726 29.31 0.204 3.05 -0.291
+QF_Week_2 7.49 10.55 49.2 0.16 0.085 0.792 29.08 0.219 3.73 0.206
+QF_Week_3 12.3 12.4 54.4 0.372 0.116 0.793 29.86 0.223 3.77 0.48
+QF_Day_1 7.89 11.4 43.1 0.181 0.09 0.834 29.15 0.196 3.92 0.245
+QF_Day_2 10.23 10.68 44.4 0.279 0.097 0.819 30.02 0.223 3.91 0.361
+QF_Day_3 9.81 10.03 44.6 0.284 0.098 0.801 30.1 0.223 1.02 0.367
+QF_1013.12 17.73 61.8 0.754 0.128 0.594 18.32 0.159 2.02 0.714
+QF_512.59 16.02 61 0.63 0.117 0.504 13.3 0.136 2.01 0.796
+QF_124.71 35.34 53.3 0.967 0.249 0.667 23.28 0.214 3.35 1.299
+Average -2.92 -1.25 53.28 -0.32 -0.02 0.96 43.61 0.25 3.99 -0.467
+CSI300 1.77 \ \ 0.009 \ \ \ 0.197 3.19 \
+Appendix B Comparison between different quantformer strategies
+(a) Factor performance
+(b) Comparison between models
+Figure 6: Comparison of factor performance and model performance
+In the provided scatter plot 6(a), the $QF\_Month\_1$ factor exhibits notable performance in terms of annual return.
+Compared to other factors, $QF\_Month\_1$ offers better returns.
+Regarding risk, as measured by the $Variance\_Covariance\_VaR\_99$, the $QF\_Month\_1$ factor’s data points are positioned in a moderately high range, indicating a balanced risk profile.
+The size and color intensity of the $QF\_Month\_1$ points, reflecting higher Alpha values, suggest a robust excess return or relative performance.
+Combining these observations, the $QF\_Month\_1$ factor likely provides a compelling investment opportunity by delivering robust returns while maintaining a moderate risk level, an attractive proposition for strategies aiming to optimize the trade-off between risk and return.
+In scatter plot 6(b), most of the factors based on quantformer perform better compared with the other 100 factors.
+The performance of quantformer factors which are trained under different scales shows ability in making profit and risk management.

### v3

@@
 Fully capturing various market variables, including long-term information, as well as essential signals that may lead to profit remains a difficult task for learning algorithms.
-In order to tackle this challenge, this paper introduces quantformer, an enhanced neural network architecture based on transformers, to build investment factors.
+In order to tackle this challenge, this paper introduces quantformer, an enhanced neural network architecture based on transformer, to build investment factors.
 By transfer learning from sentiment analysis, quantformer not only exploits its original inherent advantages in capturing long-range dependencies and modeling complex data relationships, but is also able to solve tasks with numerical inputs and accurately forecast future returns over a given period.
-This work collects more than 5,000,000 rolling data of 4,601 stocks in the Chinese capital market from 2010 to 2019.
-The results of this study demonstrated the model’s superior performance in predicting stock trends compared with other 100 factor-based quantitative strategies.
-Notably, the model’s innovative use of transformer-liked model to establish factors, in conjunction with market sentiment information, has been shown to enhance the accuracy of trading signals significantly, thereby offering promising implications for the future of quantitative trading strategies.
-The implementation details and code is available on https://github.com/QuantFormer.
+This work collects more than 5,000,000 rolling data of 4,601 stocks in the Chinese capital market from 2010 to 2023.
+The results of this study demonstrate the model’s superior performance in predicting stock trends compared with other 100-factor-based quantitative strategies.
+Notably, the model’s innovative use of transformer-like model to establish factors, in conjunction with market sentiment information, has been shown to enhance the accuracy of trading signals significantly, thereby offering promising implications for the future of quantitative trading strategies.
+The implementation details and code is available on Github.
 keywords:
-quantformer, transformer, quantitative finance, stock selection, market sentiment
+quantformer, transformer, neural networks, quantitative finance, stock selection, portfolio optimization, market sentiment
+journal: arXiv
+[…]
+[…]
 1 Introduction
@@
 However, stocks are influenced by a large number of factors, which constitute a complex system and make it difficult for people to make a profit.
-The assessment of a stock’s evolving trend is inherently challenging due to the highly volatile and interconnected nature of the market, which sets it apart from typical time series modeling [1].
+The assessment of a stock’s evolving trend is inherently challenging due to the highly volatile and interconnected nature of the market, which sets it apart from typical time series modeling (Wang et al., 2022b).
 As a result, many strategies and tools have been built, in parallel with the development of capital markets, and quantitative strategies have been playing an important role among them.
-Figure 1: Quantitative trading architecture with factors
-Some traditional quantitative tools, such as the Markowitz portfolio theory [2] and the Capital Asset Pricing Model (CAPM) [3], focus mainly on static fundamental analysis.
+Some traditional quantitative tools, such as the Markowitz portfolio theory (Markowitz, 1952) and the Capital Asset Pricing Model (CAPM) (Sharpe, 1964), focus mainly on static fundamental analysis.
 In other words, these strategies aim to make a profit by simple calculation and analysis.
@@
 Within these methods, factor-based strategies have attracted much attention.
-In 1993, [4] introduced their Fama-French Three Factor Model (FF3), which has become an influential model in quantitative trading.
-In 2015, Fama and French revised their model with a Five-Factor Asset Pricing Model (FF5) [5].
+In 1993, Fama and French (1993) introduced the Fama-French Three Factor Model (FF3), which has become an influential model in quantitative trading.
+In 2015, Fama and French revised their model with a Five-Factor Asset Pricing Model (FF5) (Fama and French, 2015).
 Besides this classical theory, numerous trading strategies have been published for decades.
-[6] built a optimal portfolio to catch future investment opportunities (FIO) by multi-factor models.
+For example, Shi et al.
+(2024) built an optimal portfolio to catch Future Investment Opportunities (FIO) by multi-factor models.
+Figure 1: Quantitative trading architecture with factors
 Quantitative trading with factors typically follows two primary approaches, which are shown in Figure 1.
@@
 Traders can execute corresponding actions when they receive the signals from factors, allowing for a dynamic response to market fluctuations based on factor insights.
-In recent years, Machine Learning (ML) has become an instrumental tool for enhancing trading algorithms and decision-making processes.
-Its core principle, enabling systems to learn from and make decisions based on data, lends itself particularly well to the vast and dynamic landscapes of stock markets.
-[7] exploits machine learning algorithms based on manual indicators, but the underlying random walk hypothesis may hamper the task of understanding inherently non-stationary series.
-With the rise of different architectures, [8] finds some novel approaches to catch the sentiment of the market in trading data, which opened a new field in quantitative trading.
+In recent years, Machine Learning (ML) has became an instrumental tool in trading algorithms and decision-making processes.
+ML methods allow systems to learn from and make decisions based on data, and lend themselves particularly well to the vast and dynamic landscapes of stock markets.
+For example, Nayak et al.
+(2015) exploited machine learning algorithms based on manual indicators, but the underlying random walk hypothesis may hamper the task of understanding inherently non-stationary series.
+With the rise of different architectures, Feng et al.
+(2019) introduced Relational Stock Ranking (RSR) to catch the sentiment of the market in trading data.
+Figure 2: Example of stock prediction
 Although there exist several previous experiments with ML factors that attempted to fetch market sentiment in the quantitative finance field, this research still faces two difficulties.
 Firstly, in the field of sentiment analysis, which is a branch of Natural Language Processing (NLP, a field of computer science that aims to understand, interpret, and generate human language), models are used to convert words in text to word vectors through word embeddings to serve as inputs.
-However, financial datasets contain categorical data instead of words such as industry types, as well as quantitative data such as price fluctuation, turnover rate, and financial indicators.
-If the input comprises only categorical data, the time series can be treated as a sentence [9].
+However, instead of words, financial datasets contain both categorical data, such as industry types, as well as quantitative data, such as price fluctuation, turnover rate, and financial indicators.
+If the input comprises only categorical data, the time series can be treated as a sentence (Gorishniy et al., 2022).
 In most cases, the input will involve numerical data, which cannot be transformed via word embeddings.
 Secondly, most NLP tasks can be transformed into sequence-to-sequence (seq2seq) problems, such as in machine translation, dialogue systems, and speech recognition.
-As an example, the transformer architecture is based on the seq2seq architecture [10].
+As an example, the transformer architecture is based on the seq2seq architecture (Vaswani et al., 2017).
 To utilize existing outputs, decoders in transformer sequentially output samples and use masking operations to handle input sequences during training.
 However, in stock prediction, where the aim is often to accurately forecast future returns over a period, the transformer model is rarely used for such tasks.
-To address these problems, we propose quantformer, which is a modified transformer architecture adapted to quantitative data, and used as an investment factor.
+To address these problems, we propose quantformer, a modified transformer architecture adapted to quantitative data, and use it as an investment factor.
 Quantformer is able to input numerical data directly, which refers to a method similar to sentiment analysis.
+Our contributions lie in following aspects:
+  • We propose quantformer, whose structure adapts to rolling stock-related time series data as inputs without positioning module and word embedding.
+The new structure with linear embedding shows better fitting to the numerical type inputs.
+  • According to our experiments, quantformer-based factors perform better in back-tests compared with other traditional factor strategies under the same time period with different trading frequencies.
+These experiments indicate a potential direction for applying transformer-based models to quantitative financial tasks.
 The paper is structured as follows.
-Section 2 discussed the previous quantitative financial works mainly based on machine learning.
+Section 2 discusses previous quantitative financial works based on machine learning.
 Sections 3 and 4 introduce quantformers.
 A factor based on quantformer will be trained and back-tested.
-For the practical backtest, we collected data from more than 4,600 stocks in the past 14 years (from 2010 to 2023) from the data-collect platform.
-To comprehensively test the ability of the factor, we divided the data by different frequencies (Section 4.2) and trained under different training scales (Section 5.3).
-Finally, the result of the back-tests including the comparison between the quantformer-factor and other 100 factors as well as the insights gained from such comparative analysis will be discussed in Sections 4 and 5.
-Figure 2: Example of stock prediction
+For the practical backtest, we collected data from more than 4,600 stocks in the past 14 years (from 2010 to 2023).
+To comprehensively test the ability of the factor, we divide the data by different frequencies (Subsection 4.2) and trained under different training scales (Subsection 5.2).
+Finally, the results of the back-tests including the comparison between the quantformer-factor and other 100 factors as well as the insights gained from such comparative analysis are discussed in Sections 4 and 5.
 2 Preliminaries
-This section briefly introduces the related works about stock prediction with market sentiment and the development of quantitative financial trading methods, especially based on Machine Learning (ML) methods.
+This section briefly introduces the related works about stock prediction with market sentiment and the development of quantitative financial trading methods, especially those based on Machine Learning (ML) methods.
 2.1 Stock data and market sentiment
-It is known that, in some cases, stock data can be deemed to influence the future trend of a stock price [11].
-[12, 13] have discussed the correlation between investor sentiment and stock market metrics such as turnover rate, indicating that investor feelings, both individual and institutional, significantly influence market behavior, affecting aspects such as volatility and returns.
+It is known that, in some cases, stock data can be deemed to influence the future trend of a stock price (Asness, 1995).
+Chen et al.
+(2020), P.H and Rishad (2020) have discussed the correlation between investor sentiment and stock market metrics such as turnover rate, indicating that investor feelings, both individual and institutional, significantly influence market behavior, affecting aspects such as volatility and returns.
 This relationship is not uniform but varies with changes in investor sentiment.
 It also highlights that investor psychology, especially during periods of high market stress can drastically impact market behavior and investor decisions.
-This link between sentiment and market performance is further affirmed by studies focusing on factors such as the turnover ratio, which are seen as reflective of market liquidity and investor behavior [14].
-[15] built a new factor based on sentiment analysis with the average of trading signals from technical trading strategies to benchmark stocks of the S&P 500 index and DJIA.
+This link between sentiment and market performance is further affirmed by studies focusing on factors such as the turnover ratio, which are seen as reflective of market liquidity and investor behavior (Naseem et al., 2021).
+Ding et al.
+(2023) built a new factor based on sentiment analysis with the average of trading signals from technical trading strategies to benchmark stocks of the S&P 500 index and DJIA.
 The sentiment factor shows correlation with stock returns.
@@
 At the same time, the finance sector, particularly quantitative trading, has started to be aware of the potential of deep learning models for predicting stock movements, portfolio optimization, and risks.
-2.2.1 Support Vector Machines (SVM)
-SVM has been used in quantitative trading, as it can manage high-dimensional spaces and intricacies found in datasets.
-[16] used SVMs to forecast financial time series in 2015, with a focus on the South Korean stock market, exhibiting the robustness of SVM against traditional statistical methods and some other computational techniques, setting a precedent for its adoption in financial forecasting.
-Similarly, [17] showcased SVMs’ prowess in anticipating the movement direction of Japanese stock markets, especially when determining bullish or bearish market sentiments.
-[18] assessed SVMs’ adaptability and precision in financial tasks such as portfolio and quantitative trading.
+2.2.1 Support Vector Machines (SVMs)
+SVMs have been used in quantitative trading, as they can manage high-dimensional spaces and intricacies found in datasets.
+Kim (2003) used SVMs to forecast financial time series in 2003, with a focus on the South Korean stock market, exhibiting the robustness of SVM against traditional statistical methods and some other computational techniques.
+Similarly, Huang et al.
+(2005) showcased SVMs’ prowess in anticipating the movement direction of Japanese stock markets, especially when determining bullish or bearish market sentiments.
+Cavalcante et al.
+(2016) assessed SVMs’ adaptability and precision in financial tasks such as portfolio and quantitative trading.
 However, SVMs can also show some disadvantages in quantitative trading.
-In [19], SVM was underscored for the accuracy of overfitting.
+In Huck (2009)’s work, SVM was underscored for the accuracy of overfitting.
 When applied to financial datasets, the model’s propensity to fit too closely to the training data might lead to reduced generalization capabilities, making the strategy ineffective on unseen data.
-In high-frequency trading, though the model displayed competency in specific scenarios, researchers pinpointed potential limitations when grappling with certain volatile market dynamics, suggesting the need for adaptations [20].
+In high-frequency trading, though the model displayed competency in specific scenarios, researchers pinpointed potential limitations when grappling with certain volatile market dynamics, suggesting the need for adaptations (Kercheval and Zhang, 2015).
 2.2.2 Long Short-Term Memory (LSTM)
-LSTM is a neural network architecture designed to capture long-term dependencies in sequential data [21].
-LSTM neural networks can accentuate and effectively model volatile financial markets, showcasing significant improvements over traditional time series models.
-[22] combined LSTM with linear Kalman filter and set experiments on different volatility stocks to build strategies.
-[23] deployed LSTM with stacked autoencoders to forecast financial time series.
-[24] tested the efficacy of LSTM for price predictions they observed LSTMs’ prowess in noisy data and their superiority in prediction accuracy over other conventional models.
-[25] combined LSTM with other architectures and found that these can enhance the feature extraction process, thereby improving prediction accuracy.
-For predicting stock prices, [26] harnessed the power of LSTM to unearth multi-frequency trading patterns.
+LSTM is a neural network architecture designed to capture long-term dependencies in sequential data (Hochreiter and Schmidhuber, 1997).
+LSTM networks can accentuate and effectively model volatile financial markets, showcasing significant improvements over traditional time series models.
+Orsel and Yamada (2022) combined LSTM with linear Kalman filter and set experiments on different volatility stocks to build strategies.
+Bao et al.
+(2017) deployed LSTM with stacked autoencoders to forecast financial time series.
+Fischer and Krauss (2018a) tested the efficacy of LSTM for price predictions.
+They observed LSTMs’ prowess in noisy data and their superiority in prediction accuracy over other conventional models.
+Sezer and Ozbayoglu (2018) combined LSTM with other architectures and found that these can enhance the feature extraction process, thereby improving prediction accuracy.
+For predicting stock prices, Zhang et al.
+(2017) harnessed the power of LSTM to unearth multi-frequency trading patterns.
 Their method underlined the adaptability of LSTMs in modeling complex situations inherent in stock prices.
 2.2.3 Gated Recurrent Units (GRUs)
-GRUs [27] are another neural network architecture which has emerged as a method in the quest for accurate financial forecasting.
+GRUs (Chung et al., 2014) are another neural network architecture which has emerged as a noteworthy method in the quest for accurate financial forecasting.
 GRUs can capture long-term dependencies in time series data and have seen increasing popularity in the world of quantitative trading.
-[23] introduced a comprehensive deep-learning framework for financial time series, which incorporated GRUs alongside LSTM and stacked autoencoders.
+Bao et al.
+(2017) introduced a comprehensive deep-learning framework for financial time series, which incorporated GRUs alongside LSTM and stacked autoencoders.
 Their model, when applied to diverse datasets, performs better than traditional time series models, accentuating the adaptability and robustness of GRUs.
-[28] centralized on LSTM networks, provided comparative insights into GRUs.
+Fischer and Krauss (2018b) centralized on LSTM networks, provided comparative insights into GRUs.
 However, in their analysis, LSTMs showed a slight edge in prediction accuracy, indicating that while GRUs are powerful, selecting between them and LSTMs may boil down to specific use cases and computational constraints.
-Beyond that, [29] presented a model fusing GRUs with Convolutional Neural Networks (CNN) [30].
+Beyond that, Pak et al.
+(2018) presented a model fusing GRUs with Convolutional Neural Networks (CNN) O’Shea and Nash (2015).
 The strategy to combine CNNs and GRUs also exhibits advantages for stock price prediction.
-[31] introduced the Attention-GRU model which combined attention and GRU to establish a new factor based on CVaR [32, 33] portfolio.
-The Attention-GRU model fitted market return, stocks return from 28 Dow Jones Industrial Average index (DJIA) stocks, which achieved better performance than other models in 8 metrics such as annual return, standard deviation and information ratio.
-2.3 Transformer in quantitative trading
-Transformer [10], which was introduced in 2017, has received considerable attention in the NLP field.
+Sun et al.
+(2024) introduced the Attention-GRU model which combined attention and GRU to establish a new factor based on CVaR portfolio (Zhu and Fukushima, 2009, Ban et al., 2018).
+The Attention-GRU model fitted market return, stocks return from 28 Dow Jones Industrial Average index (DJIA) stocks, which achieved better performance than other models in eight metrics such as annual return, standard deviation and information ratio.
+2.3 Transformer in time series prediction
+Transformer (Vaswani et al., 2017), which was introduced in 2017, has received considerable attention in the NLP field.
 According to its special encoder and decoder stacks with self-attention blocks, the transformer shows its advantage in stability, speed, and long-term memory compared with other traditional models.
-The transformer-based models such as GPT [34] and BERT [35] are the basis of the specific field of application.
-In the application field, transformer-based models such as ChatGPT [36], FinBERT [37, 38] and BloomberGPT [39] show advantages and ability in solving different types of problems in different fields.
-[40] utilized transformer to to predict the stock market index (including CSI 300, S&P 500, Hang Seng Index, and Nikkei 225) and concluded that the model could better catch the rules of stock market dynamics.
+Transformer-based models such as GPT (Radford et al., 2019) and BERT (Devlin et al., 2019) have been influential in the NLP field.
+More recent transformer-based models include ChatGPT (Brown et al., 2020), FinBERT (Araci, 2019, Yang et al., 2020) and BloombergGPT (Wu et al., 2023).
 Similar to other neural sequence transduction models, transformer also has an encoder-decoder structure.
-The model takes as input a sequence of context which could be natural language sentences and generates output with texts depending on the input, similar to a translation according to the input or a predicted sentence.
-Transformer using stacks which are called self-attention and point-wise, which connect layers in the architecture for encoder and decoder.
+The model takes as input a sequence of context, which could be natural language sentences, and generates output with texts depending on the input, similar to a translation according to the input or a predicted sentence.
+Transformer consists of stacks of layers, which include self-attention and point-wise operations.
+These layers are used to connect the encoder and decoder in the architecture.
 For these encoder and decoder layers, the self-attention method plays an important role in matrix calculation.
 The linear and softmax functions are used in the dot-product attention module.
-In the vanilla transformer model, multi-head attention layers are used and the result from each layer will be outputted after going through an additional normalization layer.
-The above-mentioned steps cover most of the progress in the vanilla transformer.
-In the quantitative trading domain, some previous researchers tried to test the potential of transformer.
-[41] firstly tried to exploit transformer on trading sequences to classify stock price movements.
+In the vanilla transformer model, multi-head attention layers are used and the result from each layer is outputted after going through an additional normalization layer.
+The above-mentioned steps cover most of the progress made with the vanilla transformer.
+Table 1 shows a set of advanced transformer models, with a focus on time series prediction tasks, which are tasks of interest in quantitative finance.
+The table highlights the modifications made to the original transformer model.
+These modifications can be categorized into innovations in attention mechanisms, encoding and decoding structure, positioning module, and mixture model.
+Some research proposed enhancements of the attention module.
+For example, Adversarial Sparse Transformer (AST) (Wu et al., 2020) introduced sparse attention using the $\alpha$-Entmax function within adversarial training.
+Autoformer (Wu et al., 2021b) replaced traditional self-attention with an auto-correlation mechanism that utilizes periodic patterns in frequency-domain time series analysis.
+Similarly, Informer (Zhou et al., 2021) suggested ProbSparse attention that selects only the most relevant queries and introduces generative-style decoding to handle long-term dependencies.
+Gated Transformer Networks (GTN) (Liu et al., 2021) used separate attention for channel-wise and temporal-wise interactions for multivariate time series.
+Patch Time Series Transformer (PatchTST) (Nie et al., 2022) divided time series into patches to model local dependencies, and Quatformer (Chen et al., 2022) attempted to use rotation-based attention to capture periodic patterns in time series.
+AirFormer (Liang et al., 2023) introduced dartboard spatial attention and causal temporal attention in air quality prediction.
+Rough transformer (Moreno-Pino et al., 2024) facilitated multi-view signature attention and continuous-time path encoding for financial market modeling for complex financial time series prediction.
+In another direction, some works tried to make changes to the encoding or decoding of transformer.
+FEDformer (Zhou et al., 2022) splitted time series data into trend and seasonal components with frequency-enhanced modules.
+To improve computational efficiency, some models, including Multivariate Transformer (Zerveas et al., 2021) gave up using a decoder and focused solely on an encoder for feature extraction for forecasting tasks.
+As a kind of revision sight, some research has proposed to replace or extend the positioning modules to better accommodate time series data.
+For instance, HFformer (Barez et al., 2023) gave up positioning modules, and uses activation functions with a lightweight linear decoder to revise the model.
+Noting that temporal order is inherently embedded in time series data, iTransformer (Liu et al., 2024) introduced a dimension-inverted structure and removes positioning.
+Transformer Hawkes Process (THP) (Zuo et al., 2020) put time intervals into the positioning module directly.
+FX-Spot Transformer (Fischer et al., 2024) used the Time to Vector (Time2Vec) function to handle irregular time intervals.
+To overcome the limitations of standalone transformer models, an approach is to integrate them with other frameworks.
+ConvLSTM (Kim et al., 2024) merged convolutional neural networks (CNNs) and LSTM layers with transformer moduled to simultaneously model spatial and temporal patterns.
+ProTran (Tang and Matteson, 2021) integrated state-space models (SSMs) with transformers to address non-Markovian dynamics in time series data, providing robust long-term forecasts.
+These hybrid approaches improved the versatility and domain adaptability of transformer-based architectures.
+Table 1: Advanced transformer-based architectures for time series data
+Reference Model Method
+Wu et al. (2020) Adversarial Sparse Transformer (AST) Sparse attention with adversarial training
+Liang et al. (2023) AirFormer Dartboard spatial self-attention and causal temporal self-attention for different training stage
+Wang et al. (2022b) ALSP-TF Local interaction and global interaction between layers
+Wu et al. (2021b) Autoformer Auto-correlation instead of attention
+Wang (2023) BiLSTM-MTRAN-TCN Combined BiLSTM, Transformer and Temporary Revolution Network (TCN)
+Kim et al. (2024) ConvLSTM Combined LSTM with transformer
+Zeng et al. (2023a) LTSF-Linear DLinear and NLinear, MLP-based with transformer
+So et al. (2019) Evolved Transformer Neural Architecture Search (NAS) in transformer
+Zhou et al. (2022) FEDformer Timestamp encoding with frequency enhanced attention
+Liu et al. (2021) Gated Transformer Networks (GTN) Two-tower transformer with channel-wise & step-wise
+Ding et al. (2020) Gaussian transformer Multi-scale Gaussian prior, orthogonal regularization and trading gap splitter for self-attention
+Barez et al. (2023) HFformer Linear decoder and spiking activation
+Zhou et al. (2021) Informer ProbSparse self-attention, self-attention distilling operation and generative style decoder
+Liu et al. (2024) iTransformer Dimension inversion and deleted positioning
+Li et al. (2019) LogSparse Transformer Causal convolution attention and query-key matching
+Eisenach et al. (2020) MQ-Transformer Positional encoding from event indicators and decoder-encoder attention for context-alignment
+Ramos-Pérez et al. (2021) Multi-Transformer Average multi-head and bagging attention and positioning based on time stamps
+Zerveas et al. (2021) Multivariate transformer Deleted decoder and applied learnable positioning module for multivariate time series
+Liu et al. (2022b) Non-stationary Transformer Series standardization and De-stationary attention
+Nie et al. (2022) Patch Time Series Transformer (PatchTST) Divided time series into patches, applied channel-independence and instance normalization
+Tang and Matteson (2021) ProTran Modeled non-Markovian dynamics by State Space Models (SSMs) to
+Liu et al. (2022a) Pyraformer Pyramidal Attention Module (PAM) and intra-scale neighboring connections model
+Chen et al. (2022) Quatformer Learning-to-rotate Attention (LRA) with global memory and trend normalization
+Kitaev et al. (2020) Reformer Locality-sensitive hashing attention and reversible residual layers
+Moreno-Pino et al. (2024) Rough Transformer Multi-view signature attention and continuous-time paths encoding transformation
+Shabani et al. (2022) Scaleformer Multi-scale framework with various models (such as FEDformer, Autoformer)
+Lin et al. (2021) State Space Decomposition Neural Network (SSDNet) Transformer with SSMs
+Ma et al. (2025) Stockformer Dual-frequency spatiotemporal encoder for data in different frequency with graph embedding and fusion attention
+Chowdhury et al. (2022) TARNet Attention score for timestamps masking
+Olorunnimbe and Viktor (2024) Temporal Fusion Transformer (TFT) LSTM learning positional encoding
+Fischer et al. (2024) Transformer for FX-Spot Time2Vec positioning module
+Zuo et al. (2020) Transformer Hawkes process (THP) Positional encoding by translating time intervals
+Cirstea et al. (2022) Triformer Triangular,variable-specific patch attention
+2.4 Transformer in quantitative trading
+In the quantitative trading domain, a few research articles have investigated the potential of transformers.
+Wang et al.
+(2022a) utilized transformer to predict stock market indices (including CSI 300, S&P 500, Hang Seng Index, and Nikkei 225) and concluded that the model can better catch the rules of stock market dynamics.
+Ding et al.
+(2020) exploited transformer on trading sequences to classify stock price movements.
 Besides using a traditional transformer, other innovative methods of transformer are worth mentioning.
-[42] devoted to optimizing the efficiency of time complexity and memory usage of transformer on extremely long time series by informer.
-[43] takes the advantage of CNNs and transformers to model short-term and long-term dependencies in financial time series.
-They showed the advantage of the proposed approach on intraday stock price prediction of S&P 500 constituents.
-[44] proposed a new ConvLSTM model by combining convolutional transformer and LSTM.
-The model was trained by stocks from S&P 500 from 2004-2021 and performs better than other models such as VaR and ARIMA.
-Similar to previous work, [1] innovated the structure of a transformer, in the form of Adaptive Long-Short Pattern Transformer (ALSP-TF).
-Their model is structurally innovated for hierarchical representation and interaction of stock price series at different context scales.
-With the help of a learnable function, they make the self-attention aware of the weighted time intervals between patterns, to adaptively adjust their dependencies beyond similarity matching.
+Zhou et al.
+(2021) optimized the efficiency of time complexity and memory usage of transformer on extremely long time series by informer.
+Zeng et al.
+(2023b) combined the advantages of CNNs and transformers to model short-term and long-term dependencies in financial time series.
+They illustrate the merits of this approach on intraday stock price prediction of S&P 500 constituents.
+ConvLSTM (Kim et al., 2024) has been trained on stocks from the S&P 500 between 2004 and 2021, and performs better than other models such as VAR and ARIMA.
+Similar to previous work, Wang et al.
+(2022b) introduced Adaptive Long-Short Pattern Transformer (ALSP-TF).
+Their model is structurally designed for hierarchical representation and interaction of stock price series at different context scales.
+With the help of a learnable function, they make self-attention aware of the weighted time intervals between patterns, to adaptively adjust their dependencies beyond similarity matching.
 In the end, they obtained more than 10% of annual return on average.
-However, transformer has some disadvantages.
-[45] mentioned that the global self-attention module focuses on point-wise token similarities without contextual insights.
+However, the transformer model can also have some disadvantages in practice.
+Xu et al.
+(2021) mentioned that the global self-attention module focuses on point-wise token similarities without contextual insights.
 As fluctuations of stocks are conditioned on composite signals over manifold periods, lacking pattern-wise interaction hinders the adequate discrimination of stock tendency and is susceptible to noise points.
-On the other hand, [46] claimed that the basic query-key matching paradigm is position agnostic.
+On the other hand, Wu et al.
+(2021a) claimed that the basic query-key matching paradigm is position agnostic.
 Although position embedding is inserted into the sequential inputs, it may not be optimal because of the inability to reveal precise distances.
 3 Methodology
-This section introduces the framework of the work and the steps of the establishment of the model, including the data processing, quantformer construction and prediction.
+Several researchers have tried to improve transformers for time series and quantitative financial applications and proposed improved models (see time series papers in Table 1), but there are still some aspects that have not been considered.
+For data in time series type, such as stock prices or daily returns, it is reasonable to assume that each data in the series is assigned a “position” automatically, which is different from sentences.
+Based on this observation, the positional encoding module on the vanilla model may not be a necessary part.
+Furthermore, market data are related to the market trend.
+For example, the return of the stock may include the market sentiment and the turnover rate reflects investor sentiment and trading enthusiasm.
+Therefore, integrating such market-related features as input can enhance the model’s ability to capture underlying market dynamics and improve the accuracy of future trend predictions.
+To implement these ideas, we propose quantformer, an enhanced transformer-based model for quantitative financial trading, and assess its performance on real-world datasets.
+This section introduces the framework of the work and the steps of the establishment of the model, including data processing, quantformer construction and prediction.
 3.1 Framework overview
@@
 (i) Data Initialization to align the time series of stocks into a regular matrix;
-(ii) Embedding the training dataset and training them by a quantformer with word-embedding layer replaced by linear layer and with no Mask(opt.) layer;
+(ii) Unlike the transformer, embedding the training dataset and training them by a quantformer with word-embedding layer replaced by linear layer and with no Mask(opt.) layer;
 (iii) Lastly, using the trained model to predict the possible trend of stocks.
@@
 3.2 Problem formulation
-Inputs
-Given the candidate stock set $S^{t}$ for $N$ stocks on the trading timestamp $t$:
-$S^{t}=\{s_{1}^{t},s_{2}^{t}\ldots s_{N}^{t}\}$
-For a stock $s_{k}^{t}$ in the stock set $S^{t}$, where $k\in\{1,\ldots,N\}$, consider the two-dimensional historical 1-time step sequence (such as one month):
-${\chi}^{t}_{k}=[x_{k}^{t-20},x_{k}^{t-19}\ldots x_{k}^{t}]\in\mathbb{R}^{20\times 2}$
-Each $x_{k}^{t}$ contains two features.
-The first one is the accumulated daily profit rate during the time step, where $p_{i}$ and $p_{i-1}$ are the close price of the $i^{\mathrm{th}}$ and ${(i-1)}^{\mathrm{th}}$ day, respectively.
-$r^{t}=\sum r_{i}=\sum\frac{p_{i}-p_{i-1}}{p_{i-1}},\quad\sigma^{t}=\sum\sigma_{i}$ (1)
-The other feature is the accumulated daily turnover rate $\sigma^{t}$ during the time step and the calculation method of $\sigma^{t}$ is shown above. So each $\chi^{t}_{k}$ can be represented as:
-$\chi^{t}_{k}=\left[\begin{matrix}r^{t-20}&r^{t-19}&\ldots&r^{t}\\ \sigma^{t-20}&\sigma^{t-19}&\ldots&\sigma^{t}\\ \end{matrix}\right]\in\mathbb{R}^{20\times 2}$
-Outputs
-For a stock set $S^{t}$ on the trading timestamp $t$, each stock $s^{t}_{i}$ has a profit rate $r_{i}^{t+1}$ on the trading timestamp $t+1$, where the calculation method is the same as for $r^{t}$ in the sequence of inputs shown in equation (1).
-Then the value in the list of next-time stamp’s profit for $N$ stocks in the timestamps $t+1$ can be represented as:
-$P^{t+1}=\{r_{1}^{t+1},r_{2}^{t+1}\ldots r_{N}^{t+1}\}$
-This list is then sorted and partitioned into $q$ equal parts, where it is set to 3 or more.
-For example, set $q=5$; for the stocks $s_{k}$ whose profit values ranked in the bottom $20\%$, middle $20\%$, and top $20\%$ respectively.
-And the dimension of the outputs $d_{f}$ is the number of selected quantiles in training.
-The value of $d_{f}$ should be larger than 3 and not larger than $q$.
-Figure 4: Training outputs remarked between different $d_{f}$
-Both the dimensions of training and predicted outputs are the same, which is the $d_{f}$ set at the beginning.
-For example, set $q=5$; for the stocks $s_{k}$ whose profit values ranked in the bottom $20\%$, middle $20\%$, and top $20\%$ respectively, and the corresponding output of $d_{f}=3$ is recorded, respectively, as
-$y^{t}_{k}=[1,0,0]^{T},\quad y^{t}_{k}=[0,1,0]^{T},\quad y^{t}_{k}=[0,0,1]^{T},\quad y^{t}_{k}=[0,0,1]^{T}$
-The dimension of the output also can be 5, and in this case, the outputs of $d_{f}=5$ are:
-$y^{t}_{k}=[1,0,0,0,0]^{T},\quad y^{t}_{k}=[0,0,1,0,0]^{T},\quad y^{t}_{k}=[0,0,0,0,1]^{T}$
-In this example, when $d_{f}=3$, the output dimension will be 3 while the dimension will be 5 as $d_{f}$ is set as 5.
-3.3 Embedding initialization
+Training input
+Consider a candidate stock set $S^{t}$ containing $N$ stocks on the trading timestamp $t$:
+$S^{t}=\{s^{t}_{n}\}^{N}_{n=1}$
+For a stock $s_{n}^{t}$ in the stock set $S^{t}$, where $n\in\{1,\ldots,N\}$, consider the two-dimensional historical 1-time step (such as one month) input feature matrix $\mathcal{X}^{t}_{n}\in\mathbb{R}^{20\times 2}$, consisting of 20 continuous timestamps:
+${\mathcal{X}}^{t}_{n}=\{x^{t-m}_{n}|m=19,18,\dots,0\}$ (1)
+Each row vector $x_{n}^{t-m}$ for $m=0,\ldots,19$ contains two features.
+The first one is the accumulated daily profit rate $r^{t-m}_{n}$ during the time step, where $p^{t-m}_{n}$ is the close price of the last day at the timestamp tmt-m and $p^{t-m-1}_{n}$ is the close price of the last day at the timestamp tm1t-m-1.
+$r^{t-m}_{n}=\frac{p^{t-m}_{n}-p^{t-m-1}_{n}}{p^{t-m-1}_{n}},\quad v^{t-m}_{n}=\sum v^{t-m,i}_{n}$ (2)
+The second feature is the accumulated daily turnover rate $v^{t-m}_{n}$ during the time step tmt-m for stock $n$.
+$v^{t-m,i}_{n}$ is the daily turnover rate and $i$ is the day index in the time period.
+To sum up, each $\mathcal{X}^{t}_{n}$ can be represented as:
+$\mathcal{X}^{t}_{n}=\left[\begin{matrix}x^{t-19}_{n}&x^{t-18}_{n}&\ldots&x^{t}_{n}\\ \end{matrix}\right]^{T}\in\mathbb{R}^{20\times 2},x^{t-m}_{n}=[r^{t-m}_{n},v^{t-m}_{n}]$ (3)
 Normalization
-The next step is to normalize the input data for each time step by zero-mean, unit-variance normalization (Z-score), the normalization equation is shown in equation (2).
-$\tilde{x}^{t}_{k,i}$ represents the normalized value at $i^{\mathrm{th}}$ day of stock $k$ at time $t$.
-$\mathbb{E}[x^{t}_{i}]$ and $\mathrm{std}[x^{t}_{i}]$ are the mean and standard deviation of the variables $x^{t}_{i}$, respectively.
-$\tilde{x}^{t}_{k,i}=\frac{x^{t}_{k,i}-\mathbb{E}[x^{t}_{i}]}{\mathrm{std}[x^{t}_{i}]}$ (2)
-In this way, the input sequences are normalized with zero mean and unit variance, which aims to reduce the influence from outlying time points and allow different features to be comparable with each other [47].
+The next step is to normalize the input data for each time step by zero-mean, unit-variance normalization (Z-score).
+The normalization equation is shown in equation (4).
+For the row vector $x^{t}_{n}$ (equation (1)), $\mathbb{E}[x^{t}]$ and $\mathrm{std}[x^{t}]$ represent the mean and standard deviation of all the row vectors $x^{t}$ at time $t$, respectively, and the resulting values $\tilde{x}^{t}_{n}$ represent the normalized values of the row data $x^{t}_{n}$.
+$\tilde{x}^{t}_{n}=\frac{x^{t}_{n}-\mathbb{E}[x^{t}]}{\mathrm{std}[x^{t}]}$ (4)
+So, the normalized two-dimensional historical 1-time step input feature matrix can be represented in this way:
+${\tilde{\mathcal{X}}}^{t}_{n}=\{\tilde{x}^{t-m}_{n}|m=19,18,\dots,0\}$ (5)
+In this way, the input sequences are normalized with zero mean and unit variance, which aims to reduce the influence from outlying time points and allow different features to be comparable with each other (Klambauer et al., 2017).
+Target labels
+For a stock set $S^{t}$ on the trading timestamp $t$, each stock $s^{t}_{n}$ has a profit rate $r_{n}^{t+1}$ on the trading timestamp $t+1$, where the calculation method is the same as for $r^{t}_{n}$ in the sequence of inputs shown in equation (2).
+The set $\mathcal{G}^{t+1}$ contains the next-time stamp’s profit for $N$ stocks in the timestamps $t+1$.
+It is defined as $\mathcal{G}^{t+1}=\{r^{t+1}\}^{N}_{n=1}$.
+Each $r^{t+1}_{n}$ in the list $\mathcal{G}^{t+1}$ is then ranked by $\Psi(r^{t+1}_{n})$, which is the empirical quantile CDF for $r^{t}_{n}$:
+$\Psi(r^{t+1}_{n})=\frac{1}{n}\sum^{n}_{i=1}\mathbbm{1}_{\{r^{t+1}_{i}\leq r^{t+1}_{n}\}}$ (6)
+For the target label for each feature $s^{t}_{n}$ in $\mathcal{G}^{t+1}$, we have a one-hot vector $y^{t}_{n}\in\mathbb{R}^{\varrho}$ based on $\Psi(r^{t+1}_{n})$.
+$y^{t}_{n}$ is partitioned into $\varrho$ equal bins, where $\varrho$ is set to 3 or more.
+Each bin corresponds to a $\varphi\times 100\%$ of the stocks in the set, where $\varphi\varrho\leq 1$.
+And the boundary term $\xi$ is used to ensure non-overlapping intervals.
+Then $y^{t}_{n}$ is represented as:
+$y^{t}_{n}=[\mathbbm{1}_{\{(i-1)(\varrho+\xi)\leq\Psi(r^{t+1}_{n})<i\varrho\ +(i-1)\xi\}}|i=1,\dots,\varrho]\in\mathbb{R}^{\varrho},\quad\xi=\frac{1-\varphi\varrho}{\varrho-1}$
+For the empirical quantile CDF $\Psi(r^{t+1}_{n})$ of the stock in the range $i\varphi,(i+1)\xi)$ for $i=1,\ldots,\varrho$, $y^{t}_{n}$ is represented as $\mathbf{0}_{\varrho}$.
+For example, suppose $\varrho=3$, $\varphi=0.2$; for the stocks $s^{t}_{n}$ whose profit values $r^{t}_{n}$ ranked in the top, middle, bottom $\varphi\times 100\%=20\%$ respectively, as ($\xi=0.2$ in this case):
+$y_{n}^{t}=\begin{cases}[1,0,0]^{T}&\text{if }\Psi(r_{n}^{t+1})\in[0,0.2)\\ [0,1,0]^{T}&\text{if }\Psi(r_{n}^{t+1})\in[0.4,0.6)\\ [0,0,1]^{T}&\text{if }\Psi(r_{n}^{t+1})\in[0.8,1.0)\\ [0,0,0]^{T}&\text{otherwise}\end{cases}$
+The stocks in the intermediate ranges ($\Psi(r^{t+1}_{n})\in[0.2,0.4),[0.6,0.8)$) are marked as $0,0,0]^{T}$.
+Alternatively, when $\varrho=5$ and $\varphi=0.2$, the stocks are divided into five equally sized quantiles, with no null labels ($0,0,0,0,0]^{T}$), where each part contains $20\%$ of the stock in $S^{t}$:
+$y^{t}_{n}\in\{[1,0,0,0,0]^{T},[0,1,0,0,0]^{T},\ldots,[0,0,0,0,1]^{T}\}$
 3.3 Quantformer encoder
-Then a quantformer encoder structure will be built in this subsection.
+Then, a quantformer encoder structure is designed as described in this subsection.
 The representation hierarchy consists of $L$ blocks of multi-head self-attention layers.
-Taking initialized stock embedding sequences $\textbf{P}\in\mathbb{R}^{N\times 20\times 2}$ as input, the canonical self-attention in [10] can perform information exchange between every time points for each stock $s_{i}$.
+Taking initialized stock embedding sequences $\mathcal{X}=\{\tilde{\mathcal{X}}^{(t)}_{n}\}^{N}_{n=1}\in\mathbb{R}^{N\times 20\times 2}$ as inputs, the canonical self-attention layers introduced in Vaswani et al.
+(2017) can perform information exchange between every time points for each stock $s^{t}_{n}$.
 To process both categorical and numerical data, the word embedding layer is replaced by a standard linear layer, utilizing linear transformations to substitute the process of word embedding.
-On the other side, in stock prediction, we aim to accurately forecast the return for a future period, thus the model’s output is generally a single value representing the probability of price increase or decrease.
+The linear embedding is shown in equation 7, where $\mathbf{W}_{E}\in\mathbb{R}^{2\times d}$ is the trainable weight matrix and $\theta_{E}\in\mathbb{R}^{d}$ is the bias and $d$ is the dimension of the hidden feature space.
+The resulting sequence $\mathcal{X}^{\prime}_{i}$ is
+$\mathcal{X}^{\prime}_{i}=\mathcal{X}_{i}\mathbf{W}_{E}+\theta_{E}$ (7)
+In stock prediction, we aim to accurately forecast the return for a future period, thus the model’s output is generally a single value representing the probability of price increase or decrease.
 Therefore, the decoder is simplified by removing the autoregressive prediction mechanism (in the decoder) and the masking operations.
-More specifically, each $\textbf{P}_{i}\in\mathbb{R}^{20\times 2}$ is transformed into query, key, and value metrics, respectively:
-$\textbf{Q}_{i,h}=\textbf{P}_{i}\textbf{W}^{Q}_{h},\quad\textbf{K}_{i,h}=\textbf{P}_{i}\textbf{W}^{K}_{h},\quad\textbf{V}_{i,h}=\textbf{P}_{i}\textbf{W}^{V}_{h}$ (3)
-where $h=1,\ldots,H$ is the head index, as well as trainable weights $\textbf{W}^{Q}_{h}$, $\textbf{W}^{Q}_{h}$, $\textbf{W}^{Q}_{h}$ $\in\mathbb{R}^{2\times d_{f}}$.
-Then, the final layer is represented by the concatenation of all attention heads.
-$\textbf{F}_{i}=\mathrm{Multihead}(\textbf{Q}_{i,h},\textbf{K}_{i,h},\textbf{V}_{i,h})=\big{(}||^{h=H}_{h=1}\mathrm{Attention}(\textbf{Q}_{i,h},\textbf{K}_{i,h},\textbf{V}_{i,h})\big{)}\textbf{W}^{O}$ (4)
-where ||||| | represents the concatenation operator and $d_{f}$ is the dimension of the projected feature space.
-After the attention computation is completed, the multi-head attention output is divided into different ”heads”.
-These outputs from the heads need to be recombined back into the original input dimension.
-$\textbf{W}^{O}$ $\in\mathbb{R}^{Hd_{f}\times 2}$ in equation 5 plays the role of combining these separate heads back into the original dimension.
-The attention head is shown below:
-$\mathrm{Attention}(\textbf{Q}_{i,h},\textbf{K}_{i,h},\textbf{V}_{i,h})=\mathrm{softmax}(\textbf{Q}_{i,h},\textbf{K}_{i,h}^{T}/\sqrt{d_{f}})\textbf{V}_{i,h}$ (5)
-In the end, all input sequences are represented in the form of:
-$\textbf{F}=[\textbf{F}_{1};\textbf{F}_{2};\ldots\textbf{F}_{N}]\in\mathbb{R}^{N\times 20\times Hd_{f}}$
-which are fed into feed-forward layers. From Section 3.2, the embedding output is denoted as:
-$\textbf{Y}^{t}=\{y^{t}_{1},y^{t}_{2},\ldots y^{t}_{N}\}\in\mathbb{R}^{N\times 1\times d_{f}}$
+For the attention head $h=1,\ldots,H$, the query, key, and value metrics for each $\mathcal{X}_{i}\in\mathbb{R}^{20\times 2}$ are computed using separate learned linear projections:
+$\mathbf{Q}_{i,h}=\mathcal{X}^{\prime}_{i}\mathbf{W}^{Q}_{h},\quad\mathbf{K}_{i,h}=\mathcal{X}^{\prime}_{i}\mathbf{W}^{K}_{h},\quad\mathbf{V}_{i,h}=\mathcal{X}^{\prime}_{i}\mathbf{W}^{V}_{h}$ (8)
+where the trainable weights $\mathbf{W}^{Q}_{h}$, $\mathbf{W}^{Q}_{h}$, $\mathbf{W}^{Q}_{h}$ $\in\mathbb{R}^{d\times\varrho}$ for the $h\mathrm{th}$ head.
+Then the attention parameter for each head is computed using scaled dot-produced attention:
+$\mathrm{Attention}(\mathbf{Q}_{i,h},\mathbf{K}_{i,h},\mathbf{V}_{i,h})=\mathrm{softmax}(\mathbf{Q}_{i,h}\mathbf{K}_{i,h}^{T}/\sqrt{d})\mathbf{V}_{i,h}$ (9)
+The outputs from all heads are concatenated and sent back to the original dimension using $\mathbf{W}^{O}\in\mathbb{R}^{\varrho H\times d}$, and the final multi-head attention output $\mathbf{F}_{i}$:
+$\mathbf{F}_{i}=\mathrm{Multihead}(\mathbf{Q}_{i,h},\mathbf{K}_{i,h},\mathbf{V}_{i,h})=\big{(}||^{h=H}_{h=1}\mathrm{Attention}(\mathbf{Q}_{i,h},\mathbf{K}_{i,h},\mathbf{V}_{i,h})\big{)}\mathbf{W}^{O}$ (10)
+where $||$ represents the concatenation operator. After applying attention to all inputs, all the output from attention modules are represented in the form of:
+$\mathbf{F}=[\mathbf{F}_{1};\mathbf{F}_{2};\ldots\mathbf{F}_{N}]\in\mathbb{R}^{N\times 20\times\varrho H}$
+which are fed into feed-forward layers.
 3.4 Output process
-Eventually, the output is represented as:
-$\textbf{Z}^{t}=\{z^{t}_{1},z^{t}_{2},\ldots z^{t}_{N}\}\in\mathbb{R}^{N\times 1\times d_{f}}$ (6)
-The output layer uses the softmax activation function, and for each $z^{t}_{i}$ can be written as:
-$z^{t}_{i}=[z^{t}_{i,1},\ldots z^{t}_{i,m},\ldots z^{t}_{i,d_{f}}]\in\mathbb{R}^{1\times d_{f}}$
-($m=1,2,\ldots,n$ is the index number between $1$ and $n$). The output probability can be represented as:
-$Y^{t}_{o}=[y^{t}_{o,1},\ldots y^{t}_{o,m},\ldots y^{t}_{o,d_{f}}]^{T}=\left[\frac{e^{z_{1}}}{\sum_{i=1}^{n}e^{z_{i}}},\dots\frac{e^{z_{m}}}{\sum_{i=1}^{n}e^{z_{i}}},\ldots\frac{e^{z_{n}}}{\sum_{i=1}e^{z_{d_{f}}}}\right]$ (7)
-In equation (7), the $y^{t}_{o,i}$ take values between 00 and $1$ and sum to $1$, which means that they can be interpreted as the probability of the stock’s performance in the slice $m$ in the next timestamp
+After processing through the multi-head self-attention and feed-forward layers, with the encoder output $F_{i}$, the output layer can be represented as:
+$\mathbf{Z}^{t}=\mathrm{Softmax}(\mathbf{F}_{i}\mathbf{W}_{Z}+\theta_{Z})=\{z^{t}_{1},z^{t}_{2},\ldots z^{t}_{N}\}\in\mathbb{R}^{\varrho}$ (11)
+where $\mathbf{W}_{Z}\in\mathbb{R}^{\varrho\times d}$ and $\theta_{Z}\in\mathbb{R}^{d}$. With the softmax function, the predicted probability of the output can be represented as:
+$\hat{Y}^{t}_{n}=\{\hat{y}^{t}_{n,i}\}^{\varrho}_{i=1}=\left\{\frac{\exp(z^{t}_{i})}{\sum^{\varrho}_{i=1}\exp(z^{t}_{i})}\right\}^{\varrho}_{i=1}\in\mathbb{R}^{\varrho}$ (12)
+In equation (12), the $\hat{Y}^{t}_{n}$ take values between 0 and $1$ and sum to $1$, which means that they can be interpreted as the probability distribution of the stock’s performance.
 3.5 Prediction
-The Mean Squared Error Loss (MSELoss) will be used to quantify the loss.
-MSE Loss is a widely used metric for assessing the discrepancy between a model’s predictions and the actual values.
-It is defined as the average of the squares of the differences between the predicted value and the actual prices:
-$\mathrm{MSELoss}=\frac{1}{N}\sum^{N}_{i=1}(y_{i}-\hat{y}_{i})^{2}$ (8)
-In equation (8), $y_{i}$ is the actual value of the $i^{\mathrm{th}}$ observation, $\hat{y}_{i}$ is the corresponding model prediction, and $N$ is the total number of input samples.
-During each epoch, the MSE Loss is computed to guide the optimization process, to minimize this loss over successive iterations.
-A lower MSE indicates a closer alignment between the model’s predictions and the actual stock prices, signifying an improvement in the model’s learning and its ability to generalize from the training data.
+The Mean Squared Error Loss (MSELoss) is used to quantify the loss. It is defined as the average of the squares of the differences between the predicted value and the actual prices:
+[…]
+where $||\cdot||^{2}_{2}$ denotes the squared Euclidean distance between predicted and target class probability vectors.
 4 Experiments
-Based on the methodology in Section 3, in this section, the set of experiments is introduced, including the data resource, implementation details, trading strategy and metrics.
+In this section, we test the ability of the quantformer architecture, introduced in the previous section, for stock price prediction.
+We are going to detail the set of experiments, including the data resource, implementation details, trading strategy and metrics.
 4.1 Dataset
-The training data of the quantformer comes from the Chinese exchange market.
-The dataset is collected from the Shanghai Stock Exchange (SHSE) and Shenzhen Stock Exchange (SZSE), which contains 4601 stocks that are listed or had been listed from January 2010 to May 2023.
-The data comes from AKShare [^1] https://akshare.akfamily.xyz/ and Tushare [^2] https://www.tushare.pro/, which are quantitative finance terminals.
+Our training data contains 4601 stocks listed either on the Shanghai Stock Exchange (SHSE) or on the Shenzhen Stock Exchange (SZSE).
+The time period ranges from January 2010 to May 2023.
+The data has been obtained from AKShare [^1] https://akshare.akfamily.xyz/ and Tushare [^2] https://www.tushare.pro/, which are quantitative finance terminals.
 The training period is from January 2010 to December 2019 and the testing period starts from January 2020.
-Closing price adjustments are used for the stock training price [48].
-Closing price adjustments are essential in stock market analysis, particularly when analyzing historical data for long-term trends and patterns.
-This price takes into account factors such as dividends, stock splits, and other corporate actions that can affect a stock’s price over time [49].
-By adjusting for these events, the adjusted closing price provides a more accurate picture of a stock’s value and performance, which is more appropriate for financial backtesting.
-4.2 Timestamp
-Data frequency concerns the number of data points within a specific unit of time, which may reflect different market characteristics [50].
-To better test the performance of the factor, three frequencies are considered: monthly, weekly, and daily stock data.
-In the first experiments, a timestamp was set as one month.
-So for a stock $s_{i}$ from the stock set $S^{t}$, it details a feature sequence in which each item in the sequence contains the accumulated profit and accumulated turnover rate in a month.
-Both the accumulated profit and accumulated turnover rate are the sum of the trading day’s data in that month.
-In some cases, if one stock was first listed on the market or resumed trading in one month, the available data may not cover whole trading days, the data of the stock from that from would still be recorded.
-However, if there is a whole month during which the stock did not have any trading (perhaps due to stopped trading or not being listed in that month), this month’s data will be recorded as “NaN”.
-In the sequence $\chi^{t}_{k}$, the sequence will not be used to train the model when there are missing values in the sequence.
-Within the first experiments, three sub-experiments were set.
+Closing price adjustments, such as dividends, stock splits, and other corporate actions that can affect a stock’s price, are applied to the stock training prices.
+These adjustments are essential for stock price analysis, especially on long time period, as they provide a more accurate picture of a stock’s value and performance over time, which is more appropriate for financial backtesting.
+(Diamond, 1971, Wei et al., 2022).
+4.2 Time frequency
+Data frequency refers to the number of data points within a specific unit of time, which may reflect different market characteristics (De Prado, 2018).
+We consider three frequencies: monthly, weekly, and daily stock data.
+In the first experiments, the time frequency is set to one month.
+For a stock $s^{t}_{n}$ from the stock set $S^{t}$, each item in the feature sequence contains the accumulated profit and accumulated turnover rate in the month (equation (3)).
+Sometimes, the available data may not cover whole trading days, for example, because a stock was first listed on the market or resumed trading after an interruption.
+In such a situation, we would still record the stock data for this day.
+However, if there is a whole month during which the stock did not have any trading (perhaps due to stopped trading or not being listed in that month), this month’s data will be recorded as “NaN”, and the sequence $\mathcal{X}^{t}_{n}$ will not be used to train the model.
+Within the first experiment, three sub-experiments are set.
 The sub-experiments share the same inputs.
-For output data, outputs with different dimensions and different lengths but the same dimension are used.
-$n=5$ and $n=7$ are used in this experiments.
-In the group of $n=5$, the first one deleted the data with output such as $[0,0,0]^{T}$, and the second one used the origin dataset (zero-output).
-Outliers of both accumulated profit and accumulated turnover rate will not be removed from the dataset these situations may happen in the future and they are expected to be predicted, though these incidents rarely happen.
+The values $\varrho=3$ and $\varrho=5$ are used in this experiment.
+The value of $\varphi$ is fixed for all the experiments to 0.2.
+In the groups of $\varrho=5$, stocks are divided into five parts and marked respectively.
+Then, in the groups of $\varrho=3$, the top $20\%$, middle $20\%$ and bottom $20\%$ of the stocks are labeled with one-hot vectors $[1,0,0],[0,1,0]$ and $[0,0,1]$, respectively.
+Stocks in the intermediate quantile (i.e, $20\%-40\%$ and $60\%-80\%$) are marked as $[0,0,0]$ (as the boundary term $\xi=0.2$ here, and $\varphi=0.2$ as well).
+In the first group of $\varrho=3$, the null labels ($0,0,0]^{T}$) are excluded from training, whereas the second group of $\varrho=3$ includes them.
+Outliers of both accumulated profit and accumulated turnover rate are not removed from the dataset.
+These situations may happen in the future and they are expected to be predicted, though these incidents rarely happen.
 In the second and third experiments, the inputs are in weekly and daily frequency, respectively.
 The sub-experiments under different groups are similar to the first one.
-All the accumulated number of parameters, the number of trained sections (such as 100 months from 2010 to 2019), and dimension are shown in Table 1.
-Table 1: Detailed information of experiments
-Strategy Frequency Output Dim Training samples Section zero-output
+All the accumulated number of parameters, the number of trained sections (such as 100 months from 2010 to 2019), and output dimensions are shown in Table 2.
+Table 2: Detailed information about experiments
+Strategy Frequency Lable dim($\varrho$) Training samples Section Null-label
 Month_1 Monthly 3  85,490 100 w/o
 Month_2 Monthly 3 142,409 100 w/
-Month_3 Monthly 5 142,409 100 w/
+Month_3 Monthly 5 142,409 100 w/o
 Week_1 Weekly 3 455,157 466 w/o
 Week_2 Weekly 3 758,300 466 w/
-Week_3 Weekly 5 758,300 466 w/
+Week_3 Weekly 5 758,300 466 w/o
 Day_1 Daily 3 3,586,435 2,420 w/o
 Day_2 Daily 3 5,140,279 2,420 w/
-Day_3 Daily 5 5,140,279 2,420 w/
+Day_3 Daily 5 5,140,279 2,420 w/o
 4.3 Implementation details
-The training process and model are implemented with PyTorch [51].
-The parameters of the model are optimized by grid search, which is reliable in low dimensional spaces and simple to implement [52].
-Here the input dimension is 2, and the dimension of the hidden feature space $d_{f}$ is 16.
+The model and its training process are implemented with PyTorch (Paszke et al., 2019).
+The hyperparameters of the model are optimized by grid search, as it is simple to implement and reliable in low-dimensional spaces (Bergstra and Bengio, 2012).
+Here the input dimension is 2, and the dimension of the hidden feature space $d$ is 16.
 The number of multi-head attention modules is 16, and the number of layers of encoder and decoder is 6.
-The model was trained on an NVIDIA GeForce RTX 2080 GPU and NVIDIA A100 Tensor Core GPU using the Adam optimizer [53] for 50 epochs, the learning rate is 0.001, and the batch size is 64.
+The model was trained on an NVIDIA GeForce RTX 2070 GPU and NVIDIA A100 Tensor Core GPU using the Adam optimizer (Kingma and Ba, 2015) for 50 epochs.
+The learning rate is 0.001, and the batch size is 64.
 4.4 Trading strategy
-A trading strategy is used for the model and Algorithm 1 shows the pseudocode of the strategy.
-Before the first trade date of the timestamp $t$, all sequence $\chi^{t}_{i,k}$ from the stock set $S^{t}$ will be put in the model and obtain the list of outputs.
-Then, the stocks will be ranked according to the first element of the output and the first $\frac{1}{q}\%$ stocks will be added to the stock pool.
-If the stock already was in the stock pool on the last timestamp, it will be held; if the stock is in the predicted pool but not in the previous pool, it will be bought in with the same proportion of the whole account.
-Stocks that are not in the predicted pool will be sold out.
+Algorithm 1 shows the pseudocode of the trading strategy.
+Before the first trade date of the timestamp $t$, all sequences $\mathcal{X}^{t}_{n}$ from the stock set $S^{t}$ are put in the model and the list of outputs $\hat{Y}^{t}$ is required.
+The portfolio value at time $t$, denoted as $P^{t}$ is updated based on the previous period’s weights and returns as follows:
+$P^{t}=P^{t-1}(\sum^{N}_{n=1}w^{t-1}_{n}(1+r^{t-1}_{n}))$ (14)
+where $w^{t}_{n}$ is the weight of stock $s^{t}_{n}$, $r^{t}_{n}$ is the return of the stock, and $\sum^{N}_{n=1}w^{t}_{n}=1$.
+To determine the weight of each stock, a trading strategy $\Phi\in\mathbb{R}^{\varrho}$ and the decision factor $\mathbf{b}$ is used, where $\mathbf{b}$ determines how many quantile groups are selected, satisfying $1\leq\mathbf{b}<\varrho$.
+The strategy does not allow short selling (as the Chinese stock market does not allow short selling), so $\Phi$ only contains 0 or $1$.
+[…]
+For example:
+  • If $\varrho=3$, $\mathbf{b}=1$, then $\boldsymbol{\Phi}$ could be one of: $[1,0,0]$, $[0,1,0]$, or $[0,0,1]$.
+  • If $\mathbf{b}=2$, then possible values of $\boldsymbol{\Phi}$ include: $[1,1,0]$, $[1,0,1]$, or $[0,1,1]$.
+Recall the predicted output $\hat{y}^{t}_{n}=\{\hat{y}^{t}_{n,i}\}^{\varrho}_{i=1}$, we sort $S^{t}$ by $\Psi(\hat{y}^{t}_{n,1})$, which is the empirical quantile CDF for $\hat{y}^{t}_{n,1}$.
+Similar to the equation 6, $\Psi(\hat{y}^{t}_{n,1})$ is computed in this way:
+$\Psi(\hat{y}^{t}_{n,1})=\frac{1}{n}\sum^{n}_{i=1}\mathbbm{1}_{\{\hat{y}^{t}_{i,1}\leq\hat{y}^{t}_{n,1}\}}$ (15)
+The sorted predicted vector $\tilde{y}^{t}_{n}$ for stock $s^{t}_{n}$ at time t is shown below.
+Here the stocks are ranked into $\varrho$ parts and each part contains $\varphi\times 100\%$ of the stocks, and the parameters $\varrho$ and $\varphi$ satisfy $\varphi\varrho\leq 1$:
+$\tilde{y}^{t}_{n}=[\mathbbm{1}_{\{(i-1)(\varrho+\xi)\leq\Psi(\hat{y}^{t}_{n,1})<i\varrho\ +(i-1)\xi)}|i=1,\dots,\varrho]\in\mathbb{R}^{\varrho},\quad\xi=\frac{1-\varphi\varrho}{\varrho-1}$
+Then the weight for each selected stock is computed using equation 16 shown below, and all the chosen stocks are equal-weighted in this way:
+$w^{t}_{n}=\frac{1}{\mathbf{b}\cdot\varrho}(\tilde{y}^{t-1}_{n}\Phi^{T}\tilde{y}^{t}_{n})^{T}\cdot\mathbf{1}$ (16)
 The same method is run repeatedly during the subsequent periods.
 The backtest starts from January 2020, in other words, the result of the sequences from May 2018 to December 2019 will be used as the first stock pool to trade.
+The exchange fee in the Chinese stock market is 0.00384% (Shanghai stock exchange, 2023; Shenzhen stock exchange, 2023), and the brokerage commission is capped at 0.3% of the transaction amount (China securities regulatory commission, 2002).
+To keep our backtest results conservative, we set the transaction fee in the backtest to 0.3%.
 Algorithm 1 Trading strategy
-1:The sequence $x^{t}_{i,k}$ for each stock $s_{i}$ in the set $S^{t}$ and the amount of cash.
-2:run_monthly(trade, monthday=1, time=‘open’)
-3:function obtain_stockpool($x^{t}_{i,k}$):
-4:     stockpool $=$ sort(model($x^{t}_{k}$), key=lambda x : x[1], reverse=True)
-5:     return stockpool
+1:The feature sequence $\chi^{t}_{n}$ for each stock $s^{t}_{n}\in S^{t}$, initial portfolio cash $P^{0}$, trading strategy $\Phi$, strategy decision factor $\mathbf{b}$.
+2:run_frequency(trade, frequency, startday=1, time=‘open’)
+3:function Predict($\chi^{t}_{n}$): return $\hat{y}^{t}_{n}=\text{quantformer}(\chi^{t}_{n})$
+4:end function
+5:function SortLabel($\hat{y}^{t}_{n,1}$): return $\tilde{y}^{t}_{n}=\left[\mathbbm{1}\!\left\{\frac{i-1}{\varrho}\leq\Psi(\hat{y}^{t}_{n,1})<\frac{i}{\varrho}\right\}\right]_{i=1}^{\varrho}$
 6:end function
-7:function trade
-8:     stockpool $=$ obtain_stockpool($x^{t}_{i,k}$)
-9:     for stock in previous_stockpool do
-10:         if stock not in stockpool then
-11:              order(stock, 0) $\triangleright$ sellout
-12:         end if
-13:     end for
-14:     for stock in stockpool do
-15:         if stock not in previous_stockpool then
-16:              order(stock, amount / len(stockpool)) $\triangleright$ buy
-17:         end if
-18:     end for
-19:end function
+7:function ComputeWeight($\tilde{y}^{t-1}_{n}$, $\tilde{y}^{t}_{n}$, $\Phi$):
+8:  numerator = $\tilde{y}^{t-1}_{n}\cdot\Phi^{\top}\cdot\tilde{y}^{t}_{n}$
+9:  denominator $\sum_{k=1}^{N}\tilde{y}^{t-1}_{k}\cdot\Phi^{\top}\cdot\tilde{y}^{t}_{k}$
+10:  return $w^{t}_{n}=\frac{1}{\mathbf{b}\cdot\varrho}\cdot\frac{\text{numerator}}{\text{denominator}}$
+11:end function
+12:function Trade($w^{t}_{n}$, $P^{t}$):
+13:  for $s^{t}_{n}$ in $S^{t}$ do order($s^{t}_{n}$, $w^{t}_{n}\cdot P^{t}$)
+14:  end for
+15:end function
 4.5 Metrics
-Formally, the Sharpe Ratio (SR) and the $\alpha$ ratio will be used to test the performance of the strategy.
-The SR [3] is a measure of risk-adjusted return that describes the additional earnings an investor receives for each standard deviation unit increase, which is shown in equation (9), where $R_{p}$ is the return of the portfolio and $R_{f}$ is the risk-free rate is the rate of London Interbank Offered Rate (LIBOR) which calculated averaged from estimates submitted by banks in London.
-The risk-free rate in the backtest will be calculated under the average LIBOR rates during the test period.
+The Sharpe Ratio (SR) (Sharpe, 1964) and the $\alpha$ rate will be used to test the performance of the strategy.
+The SR is a measure of risk-adjusted return that describes the additional earnings an investor receives for each standard deviation unit increase (equation (17)), where $R_{p}$ is the return of the portfolio and $R_{f}$ is the risk-free rate, chosen to be the London Interbank Offered Rate (LIBOR) which averages rate estimates submitted by banks in London [^3] Although LIBOR officially came to an end on September 30th, 2024, the time period covered by our backtesting dataset is prior to this date.
+https://www.bankofengland.co.uk/news/2024/october/the-end-of-libor.
+The risk-free rate in the backtest will be calculated as the average LIBOR rate during the test period.
 $\mathrm{SR}=\frac{\mathbb{E}[R_{p}]-R_{f}}{\mathrm{std}[R_{p}]}$ (17)
 Alpha represents the excess return of an investment portfolio relative to its benchmark.
-It reflects the stock selection skills of an investment.
+It aims to measure the stock selection skills of an investment strategy.
 A positive alpha indicates that the investment portfolio has achieved a higher return than its benchmark after risk adjustment.
 Specifically, alpha is the excess return of the actual portfolio return over its expected theoretical return.
-Equation (10) shows the calculation method of the alpha rate, where $E(R_{m})$ is the market return and $\beta$ is the correlation between the portfolio and systemic risk.
-$Alpha=(R_{p}-R_{f})-\beta(E[R_{m}]-R_{f})$ (10)
-Besides the SR ratio and alpha rate, the annual return (AR) of the stock, the annual excess return (AER), the average turnover rate (TR) of the portfolio, and the win rate (WR) will also shown.
-The excess returns are returns achieved above and beyond the return of a proxy and the CSI 300 index is used as the basis when calculating the excess returns.
-The turnover rate is the average percentage of the portfolio adjusted daily in terms of market.
-The win rate is the percentage of trade days where the portfolio encounters a positive return overall trade days.
-$VaR=\mu+\sigma N^{-1}(X)$ (11)
+Equation (18) shows the calculation method of the alpha rate, where $E(R_{m})$ is the market return and $\beta$ is the correlation between the portfolio and the benchmark’s systemic risk.
+$\alpha=(R_{p}-R_{f})-\beta(\mathbb{E}[R_{m}]-R_{f})$ (18)
+Besides the SR ratio and alpha rate, the annual return (AR) of the stock, the annual excess return (AER), the win rate (WR) and the average turnover rate (TR) of the portfolio will also be shown.
+The excess returns are returns achieved above and beyond the return of a proxy, chosen to be the CSI 300 index.
+The win rate is the percentage of trading days during which the portfolio strategy generates a positive return.
+The turnover rate of the portfolio is the proportion of portfolio weight changes between two consecutive rebalancing periods.
+The definition of TR at each time $t$ is:
+$\mathrm{TR}^{t}=\frac{1}{2}\sum^{N}_{k=1}|w^{t}_{n}-w^{t-1}_{n}|$ (19)
+where $w^{t}_{n}$ and $w^{t-1}_{n}$ denote the portfolio weight of the stock $s_{n}$ in the portfolio at time $t$ and t1t-1, respectively.
+Value at risk (VaR) is a method to summarize the total risk in a portfolio (Hull, 2012).
+$\mathrm{VaR}_{\alpha}=\inf\{x:P(\mathcal{L}>x)\leq\alpha\}$ (20)
 […]
@@
 Building upon the analysis of individual strategies, this section presents an evaluation and discussion of the experimental outcomes.
-The discussion focuses on the comparative effectiveness of different timestamps (monthly, weekly, and daily) and different training scales.
-The results underscore the nuanced relationship between trading frequency and market performance, offering valuable insights for the development of robust quantitative trading strategies.
+The discussion focuses on the comparative effectiveness of different time frequencies (monthly, weekly, and daily) and different training scales.
 5.1 Overall performance
-The result of the experiment is shown in Table 2.
+Our results are shown in Table 3.
 For the monthly group, Month_1 shows an annual return of 17.35% and an annual excess return of 19.43%, indicating a robust performance over that period.
 Conversely, Month_3 shows a diminished annual return and annual excess return of 7.37% and 9.91% respectively.
-The turnover rate is notably high in the second month at 51.69%.
+The is notably high in the second month at 51.69%.
 The win rate ranges from 43.1% to 57.3%, with the first month showing the highest win rate.
-The Sharpe Ratio and Alpha, vary across the strategies, with some periods showing negative values, suggesting underperformance relative to the risk taken, while others show positive values.
-In the comparison of 99% VaR, the strategies in monthly frequency perform better than others, where Month_1 and Month_3 are lower than 3%, which means the portfolio maximum loss 2.81% and 2.3% of return a day in 99% confidence.
+The Sharpe Ratio and Alpha vary across the strategies, with some periods showing negative values, suggesting underperformance relative to the risk taken, while others show positive values.
+In the comparison of 99% VaR, the strategies in monthly frequency perform better than others, where Month_1 and Month_3 are lower than 3%.
+This means that, with 99% confidence, the portfolio maximum daily losses are 2.81% and 2.3% respectively.
 Table 3: Result of experiments
@@
 CSI300 1.77% \ \ \ -0.015 \ 3.19%
-In the weekly strategy results, Week_3 showed a remarkable annual return of 12.3% and an annual excess return of 12.73%, while Week_1 displayed a negative annual return.
-For the daily strategies, Day_2 had a return of 10.23% and an annual excess return of 10.94%.
-The turnover rate across the daily strategies showed lesser variation than the monthly or weekly, indicating a more consistent trading frequency.
-The win rate for the daily strategies remained fairly stable, hovering around the mid-40% mark.
+In the weekly strategy results, Week_3 shows an annual return of 12.3% and an annual excess return of 12.73%, while Week_1 displays a negative annual return.
+For the daily strategies, Day_2 has a return of 10.23% and an annual excess return of 10.94%.
+The turnover rate across the daily strategies shows lesser variation than the monthly or weekly, indicating a more consistent trading frequency.
+The win rate for the daily strategies remain fairly stable, hovering around the mid-40% mark.
 The Sharpe Ratio and Alpha for weekly and daily strategies show a mix of positive and negative values, reflecting the fluctuating nature of shorter-term trading efficacy.
-To demonstrate the advantages of the quantformer factor over traditional factors, 100 price-volume type factors from JoinQuant  [^3] https://www.joinquant.com/ (one of China’s largest financial quantitative platforms) were selected for backtesting under the same trading strategy (Algorithm 1).
-Price-volume factors were used because these are calculated based on stocks’ prices, volumes, and turnover rates, which are similar to the training data used by quantformer factors (also used stock turnover rates and price data).
-The detailed performance of factors and benchmark is shown in Appendix A.
-On average, these factors achieved an annual return of -3.78%, an average excess return of -2.15%, and a Sharpe ratio of -0.36, where the benchmark return is -1.77%.
-The average maximum drawdown of the strategy was 44.88%, which means the potential worst-case scenario or the most extreme possible loss.
-The average volatility of the strategy ($\sigma_{p}$) was 0.26, with the benchmark at 0.197.
+To demonstrate the advantages of the quantformer factor over traditional factors, 100 price-volume type factors from JoinQuant  [^4] https://www.joinquant.com/ (one of China’s largest quantitative finance platforms) are selected for backtesting under the same trading strategy (Algorithm 1).
+Price-volume factors are used because these are calculated based on stocks’ prices, volumes, and turnover rates, which are similar to the training data used by quantformer factors.
+The detailed performance of factors and benchmark is shown in A.
+On average, these factors achieves an annual return of -3.78%, an average excess return of -2.15%, and a Sharpe ratio of -0.36, where the benchmark return is -1.77%.
+The average maximum drawdown of the strategy is 44.88%, and the average volatility of the strategy ($\sigma_{p}$) is 0.26, with the benchmark at 0.197.
 Figure 4: Backtest results based on different factors
-Among these factor-based strategies, the quantformer factor performed well.
-Except for the win rate, which was slightly lower than a few strategies, it excelled in terms of annual return, annual excess return, $\sigma_{p}$, Sharpe Ratio and sortino ratio ranking the best among the 101 factors.
-The 99% VaR of quantformer factor is in the quantile of the first 10% among all factors.
+Among these factor-based strategies, the quantformer factor performs well.
+Except for the win rate, which is slightly lower than a few strategies, it performs well in terms of annual return, annual excess return, $\sigma_{p}$, Sharpe ratio and Sortino ratio, ranking the best among the 101 factors.
+The 99% VaR of the quantformer factor is among the best 10% factors.
 Figure 4 illustrates the return curves of the quantformer (QF_Month_1) alongside the benchmark and some traditional factors.
@@
 Most of these factors are compared based on a window of 20 timestamps, as the quantformer factor was also trained on 20 timestamps.
-In comparison with these factors, the quantformer factor (blue line) demonstrated a significantly better performance in terms of returns than the benchmark (orange line) and other factor strategies, expressing the improvement of the quantformer factor over traditional price-volume factors.
-5.2 Half-life period test
-To test the half-life period of the factor, another three portfolios based on $Month\_3$ were conducted.
-The first one is the stock pool used in the next month; the second is using it in the month after the month; the last one is using it in the third month after the last feature of inputs.
-For instance, in the first strategy, the trade will start in January 2020; in the second, it will start in February 2020; in the second, it will start in March 2020.
-For the period from January 2020 to February 2022, the strategy achieved a total return (TotalR) of 65.29% and an annual return of 17.87%.
-The total excess return (TER) over this period was 66.39%, indicating a strong performance above a benchmark or risk-free rate.
-The Sharpe Ratio for this period was 0.863, which is a respectable figure that suggests a favorable risk-adjusted return.
-Additionally, the strategy generated an alpha of 0.164, demonstrating its ability to outperform the expected return on a risk-adjusted basis.
-Table 3: Result of the same strategy with different starting month
-Strategy TotalR AR TER SR Alpha CSI300
-2020/01-2022/02 65.29% 17.87% 66.39% 0.863 0.164 -0.66%
-2020/02-2022/03 67.52% 18.21% 65.58% 0.940 0.161 1.17%
-2020/03-2022/04 56.41% 15.63% 52.96% 0.739 0.135 2.26%
-In the subsequent period from February 2020 to March 2023, the strategy yielded a slightly higher total return of 67.52% and an annual return of 18.21%.
-The excess return was 65.58%, with an improved Sharpe Ratio of 0.940, reflecting a better risk-adjusted return compared to the previous period.
-The alpha for this interval was 0.161, indicating the strategy’s consistent performance in delivering additional returns.
-Lastly, Table 3 shows the result of the backtest.
-From March 2020 to April 2023, the strategy’s total return decreased to 56.41%, with an annual return of 15.63% and an excess return of 52.96%.
-The Sharpe Ratio also declined to 0.739, but it still represents a relatively efficient investment compared to the volatility of returns.
-The alpha reduced to 0.135, which, while lower than the other periods, still indicates that the strategy managed to provide returns above those predicted by its risk profile.
-Comparing the results between monthly, weekly, and daily strategies provides further insights.
-The monthly strategies can deliver the highest returns, suggesting that longer-term signals better capture short-term trends.
-The weekly strategies underperform monthly ones but outperform the daily frequency overall.
-This indicates weekly data strikes a balance between filtering noise and responding swiftly to emerging patterns.
-The daily frequency strategies perform lower annual returns and higher risk compared to others This highlights the challenge of making profitable trades at higher frequencies - noise and volatility make reliable signals more difficult.
-The distribution of returns also shows increased dispersion at the daily level.
-However, daily strategies achieve more consistent turnover rates around 5-7%.
-This allows for portfolio adjustments while avoiding excessive trading costs.
-In contrast, monthly strategies see a turnover rate above 50% in some cases.
-There is likely an optimal point between 1 week to 1 month where matches market regularly.
+In comparison with these factors, the quantformer factor (blue line) demonstrates a significantly better performance in terms of returns than the benchmark (orange line) and other factor strategies.
+Compared to other factors by the Sharpe ratio, the quantformer (0.915) still performs better than others (the highest Sharpe ratio among these factors is 0.243).
+Furthermore, the quantformer also shows a lower downside risk in terms of risk assessment as it achieves a VaR of 2.81, compared to the highest VaR of 2.86 among other factors.
+These indicate the improvement of the quantformer factor over traditional price-volume factors.
 5.2 Training under different scales
-Understanding data volume and data frequency is critical to designing infrastructure for processing data and model stability [55].
-To test the ability of stock selection of the model, three other factors ”QF_10%”, ”QF_5%”, ”QF_1%” are trained and tested (with ”QF_Month_1%, named as ”QF_20%”).
-The four groups are all trained under $d_{f}$ = 3, which include the first, middle, and tail 20%, 10%, 5%, and 1% of stocks in ranking, respectively.
-The backtest period and other settings are the same with 4.
+To test the stock selection capabilities of the model, three other factors “QF_10%”, “QF_5%”, “QF_1%” are trained and tested (with “QF_Month_1%”, named as “QF_20%”).
+The four groups are all trained under $\varrho$ = 3, which include the first, middle, and tail 20%, 10%, 5%, and 1% of stocks in ranking, respectively.
+The backtest period and other settings are the same as in Section 4.
 Table 4: Result of factors under different scales
 Strategy AR AER $\sigma_{p}$ WR SR Alpha VaR
-QF_2017.35% 19.43% 0.162 57.8% 0.915 0.162 2.81%
-QF_1013.12% 17.73 0.159 61.8% 0.574 0.128 2.023%
-QF_512.59% 16.02% 0.136 61% 0.63 0.117 2.015%
-QF_124.71% 35.74% 0.214 53.3% 0.967 0.249 3.048%
+QF_20% 17.35% 19.43% 0.162 57.8% 0.915 0.162 2.81%
+QF_10% 13.12% 17.73 0.159 61.8% 0.574 0.128 2.023%
+QF_5% 12.59% 16.02% 0.136 61% 0.63 0.117 2.015%
+QF_1% 24.71% 35.74% 0.214 53.3% 0.967 0.249 3.048%
 CSI300 1.77% \ 0.197 \ -0.015 \ 3.19%
 Table 4 shows the result of the backtest.
-QF_10strategy recorded an AR of 13.12% and an AER of 17.73%, with a slightly lower portfolio volatility of 0.159.
-This strategy had the highest WR at 61.8%.
-QF_5strategy delivered 12.59% AR and 16.02% AW, with the lowest portfolio volatility at 0.136.
-It had the lowest VaR at 2.015%.
-QF_1strategy outperformed the others, achieving the highest AR of 24.71% and the highest AER of 35.74%, though it had the highest volatility at 0.214.
+QF_10% strategy records an AR of 13.12% and an AER of 17.73%, with a slightly lower portfolio volatility of 0.159.
+This strategy has the highest WR at 61.8%.
+QF_5% strategy delivered 12.59% AR and 16.02% AW, with the lowest portfolio volatility at 0.136.
+It has the lowest VaR at 2.015%.
+QF_1% strategy outperformed the others, achieving the highest AR of 24.71% and the highest AER of 35.74%, though it takes the highest volatility at 0.214.
 These four factors all perform better than CSI 300 based on their return and risk.
-The result shows that the model takes the ability to select stocks under different training set and requirements.
+These results suggest that the quantformer model’s performance seems robust to the choice of different training sets and requirements.
 6 Conclusion
-This study proposed a new model, quantformer for quantitative stock prediction and trading.
-We addressed the need for handling numerical input data rather than text and adapting the model for forecasting tasks rather than sequence-to-sequence problems common in NLP.
-To enable direct processing of numerical time series data, we replaced the word embedding layer with a standard linear layer and removed the output masking operations.
-We also simplified the decoder to produce a probability distribution over future price movements rather than autoregressively generating token sequences.
-Eventually, we tested the factor in the reality back test of the financial market and compared the factor with other 100 price-volume factors.
+This study proposes a new neural network architecture, quantformer, inspired by the transformer architecture, for quantitative stock prediction and trading.
+We address the need for handling numerical input data rather than text, and adapted the model for forecasting tasks rather than sequence-to-sequence problems common in NLP.
+To enable direct processing of numerical time series data, we replace the word embedding layer with a standard linear layer and removed the output masking operations.
+We also simplify the decoder to produce a probability distribution over future price movements rather than autoregressively generating token sequences.
+We also backtest the quantformer factor on financial market data and compare it with 100 other price-volume factors.
 Our experimental results demonstrate the promise of this approach.
-The model-based trading strategies were able to deliver substantial excess returns over the benchmark under a longer time period, with Sharpe Ratios indicating sound risk-adjusted performance.
+The quantformer-based trading strategies are able to deliver substantial excess returns over the benchmark under a longer time period, with Sharpe Ratios indicating sound risk-adjusted performance.
 The positive alphas affirm the strategy’s ability to outperform expected returns after accounting for risk factors.
-There is further scope to enhance the framework, for instance by incorporating additional signals like news and fundamentals as auxiliary inputs.
-The self-attention mechanism could likely be improved to encode greater temporal context, or use transformer-based models such as GPT-4 [56] and Claude3 [57] to establish the strategy as these models are fine-tuned and may perform better for quantitative finance tasks.
-Overall, our work illustrates the viability of quantformer for financial data modeling, provides a flexible framework to model the market from multiple perspectives of financial time series and develops profitable trading strategies that consider different variables.
-With further research to solve the above limitations, we believe such ML-based quantitative methods hold rich potential.
-The framework provides a flexible foundation for better understanding markets and developing profitable trading strategies.
+Overall, our work illustrates the viability of the quantformer architecture to handle financial time series, design profitable and robust trading strategies, and calls for further analysis and investigations of this proposed ML-based quantitative methods.
 The implementation code of quantformer is available at https://github.com/zhangmordred/QuantFormer.
 7 Acknowledgements and declaration
-This work was funded by the Natural Science Foundation of China (12271047); Guangdong Provincial Key Laboratory of Interdisciplinary Research and Application for Data Science, BNU-HKBU United International College (2022B1212010006); UIC research grant (R0400001-22; UICR0400008-21; R72021114); Guangdong College Enhancement and Innovation Program (2021ZDZX1046).
+This work was funded by the Natural Science Foundation of China (12271047); Guangdong Provincial/Zhuhai Key Laboratory of Interdisciplinary Research and Application for Data Science, Beijing Normal-Hong Kong Baptist University (2022B1212010006); BNBU research grants (UICR0400008-21; UICR0700041-22; R72021114); and the Guangdong College Enhancement and Innovation Program (2021ZDZX1046).
 The authors have no competing interests to declare that are relevant to the content of this article.
 Appendix A Detailed backtest results between 100 factors and the quantformer factor
-All the data of factors comes from JoinQuant [^4] https://www.joinquant.com/, which is one of the most common quantitative finance platforms.
-Detailed descriptions of the following factors are available at JoinQuant Factors.
+All the data of factors come from JoinQuant [^5] https://www.joinquant.com/, which is one of the most common quantitative finance platforms in China.
+Detailed descriptions of the following factors are available at https://www.joinquant.com/help/api/help#factor_values.
 Ten metrics are used to evaluate the performance of each factor:
   • Annual Return (AR) and Annual Excess Return (AER) are two metrics to reflect the return of the portfolio directly.
-The excess returns are returns achieved above and beyond the return of a proxy and the CSI300 index is used as the basis when calculating the excess returns.
-  • Win Rate (WR) is the metric to evaluate the fraction between the action that makes a profit and all the actions.
-  • Sharpe Ratio (SR) is a measure of risk-adjusted return that describes the additional earnings an investor receives for each standard deviation unit increase, which is shown in equation (9).
+The excess returns are returns achieved above and beyond the return of a proxy and the CSI 300 index is used as the basis when calculating excess returns.
+  • Win Rate (WR) is the metric defined as the fraction between the number of trading periods that generate a profit over the total number of trading periods.
+  • Sharpe Ratio (SR) is a measure of risk-adjusted return that describes the additional earnings an investor receives for each unit of the standard deviation of returns. It is shown in equation (17).
   • The Alpha measures the ability of a portfolio to generate returns above the market benchmark.
-A positive alpha indicates that the investment portfolio has achieved a higher return than its benchmark after risk adjustment, which is shown in equation (10).
-  • Beta is the correlation between the portfolio and systemic risk, reflecting the sensitivity of the strategy to the change in the market.
+A positive alpha indicates that the investment portfolio has achieved a higher return than its benchmark after risk adjustment.
+It is shown in equation (18).
+  • Beta is the correlation between the portfolio and systemic risk, reflecting the sensitivity of the strategy to changes in the market.
 For the daily return of the strategy, $D_{p}$, and the daily return of the benchmark $D_{m}$, the equation of beta is:
-    $Beta=\beta_{p}=\frac{Cov(D_{p},D_{m})}{Var(D_{m})}$
-    where $Cov(D_{p},D_{m})$ is the correlation of between $D_{p}$ and $D_{m}$, $Var(D_{m})$ is the variance of daily return of benchmark.
+    $\mathrm{Beta}=\beta_{p}=\frac{\mathbb{C}\mathrm{ov}(D_{p},D_{m})}{\mathbb{V}\hskip-0.84pt\mathrm{ar}(D_{m})}$
+    where $\mathbb{C}\mathrm{ov}(D_{p},D_{m})$ is the correlation between between $D_{p}$ and $D_{m}$, and $\mathbb{V}\hskip-0.84pt\mathrm{ar}(D_{m})$ is the variance of the daily returns of the benchmark.
   • Max Drawdown (MD) means the potential worst-case scenario or the most extreme possible loss from the previous peak.
-For the trough value of a portfolio $V_{trough}$ and the peak value of the portfolio $V_{peak}$, the MDD is calculated in the following way:
-    $MDD=\frac{V_{peak}-V_{trough}}{V_{peak}}$
-  • Portfolio volatility ($\sigma_{p}$) is the standard deviation of the portfolio returns.
-$\sigma_{p}$ avoids the problem of negative deviations canceling with positive deviations and also penalizes larger deviations from the mean.
-It provides a kind of weighted average deviation in which large deviations carry more weight [58].
+For the trough value of a portfolio $A_{\mathrm{trough}}$ and the previous peak value of the portfolio $A_{\mathrm{peak}}$, the MD is calculated in the following way:
+    $\mathrm{MD}=\frac{A_{\mathrm{peak}}-A_{\mathrm{trough}}}{A_{\mathrm{peak}}}$
+  • Portfolio volatility ($\sigma_{p}$) is the standard deviation of the portfolio returns. It provides a kind of weighted average deviation in which large deviations carry more weight (Bacon, 2023).
   • Sortino Ratio (STN) differentiates harmful volatility from total overall volatility by using the standard deviation of negative portfolio returns.
@@
 QF_Day_3 9.81 10.03 44.6 0.284 0.098 0.801 30.1 0.223 1.02 0.367
-QF_1013.12 17.73 61.8 0.754 0.128 0.594 18.32 0.159 2.02 0.714
-QF_512.59 16.02 61 0.63 0.117 0.504 13.3 0.136 2.01 0.796
-QF_124.71 35.34 53.3 0.967 0.249 0.667 23.28 0.214 3.35 1.299
+QF_10% 13.12 17.73 61.8 0.754 0.128 0.594 18.32 0.159 2.02 0.714
+QF_5% 12.59 16.02 61 0.63 0.117 0.504 13.3 0.136 2.01 0.796
+QF_1% 24.71 35.34 53.3 0.967 0.249 0.667 23.28 0.214 3.35 1.299
 Average -2.92 -1.25 53.28 -0.32 -0.02 0.96 43.61 0.25 3.99 -0.467
@@
 Appendix B Comparison between different quantformer strategies
-(a) Factor performance
-(b) Comparison between models
-Figure 6: Comparison of factor performance and model performance
-In the provided scatter plot 6(a), the $QF\_Month\_1$ factor exhibits notable performance in terms of annual return.
-Compared to other factors, $QF\_Month\_1$ offers better returns.
-Regarding risk, as measured by the $Variance\_Covariance\_VaR\_99$, the $QF\_Month\_1$ factor’s data points are positioned in a moderately high range, indicating a balanced risk profile.
-The size and color intensity of the $QF\_Month\_1$ points, reflecting higher Alpha values, suggest a robust excess return or relative performance.
-Combining these observations, the $QF\_Month\_1$ factor likely provides a compelling investment opportunity by delivering robust returns while maintaining a moderate risk level, an attractive proposition for strategies aiming to optimize the trade-off between risk and return.
-In scatter plot 6(b), most of the factors based on quantformer perform better compared with the other 100 factors.
-The performance of quantformer factors which are trained under different scales shows ability in making profit and risk management.
+Figure 5: Comparison between factor performance and quantformer performance
+In the provided scatterplot 5, the $\mathrm{QF\_Month\_1}$ factor exhibits notable performance in terms of annual return.
+Compared to other factors, $\mathrm{QF\_Month\_1}$ offers better returns.
+Regarding risk, as measured by the
+$\mathrm{Variance\_Covariance\_VaR\_99}$, the $\mathrm{QF\_Month\_1}$ factor’s data points are positioned in a moderately high range, indicating a balanced risk profile.
+The size and color intensity of the $\mathrm{QF\_Month\_1}$ points, reflecting higher Alpha values, suggest a robust excess return or relative performance.
+Combining these observations, the $\mathrm{QF\_Month\_1}$ factor likely provides a compelling investment opportunity by delivering robust returns while maintaining a moderate risk level, an attractive proposition for strategies aiming to optimize the trade-off between risk and return.
+Overall, most of the factors based on quantformer perform better compared with the other 100 factors.
+Appendix C Symbol table
+Symbol Description
+$A_{\mathrm{peak}},A_{\mathrm{trough}}$ peak value of the portfolio
+$\alpha$ alpha ratio
+$\mathbf{b}$ decision factor in the strategy
+$\theta$ bias in the model
+$\beta_{p}$ beta ratio
+$D_{p}$ daily return
+$h,H$ attention head
+$\chi^{t}_{n}$, $\tilde{\chi}^{t}_{n}$ input feature matrix, normalized sequence
+$\hat{Y}^{t}_{n}$ predicted output
+$\mathbb{E}[x]$ mean
+$\mathcal{G}^{t}$ set of the next timestamp profit
+$\mathrm{MD}$ max drawdown
+$t$ timestep
+$m$ the number of timesteps before $t$
+$\mathrm{MSELoss}$ MSE loss
+$n$ the index of the stock
+$p^{t}$ the close price of the stock
+$P$ amount value of portfolio
+$\Phi$ trading strategy
+$\Psi$ the empirical quantile CDF
+$R_{f}$, $R_{m}$, $R_{p}$ return of risk-free rate, market return, return of portfolio
+$r^{t}_{n}$ profit rate
+$s^{t}_{n}$ stock
+$S^{t}$ stock set
+$\mathrm{SR}$ Sharpe ratio
+$\mathrm{std}[x]$ standard deviation
+$\mathrm{TR}$ turnover rate of the portfolio
+$v^{t}_{n}$ turnover rate of the stock over the time period $t$
+$\mathrm{VaR}_{\alpha}$ value at risk at $\alpha$ level
+$w^{t}_{n}$ weight of portfolio
+$x^{t}_{n}$, $\tilde{x}^{t}_{n}$ row vector, normalized vector
+$y^{t}_{n}$, $\tilde{y}^{t}_{n}$ target label, sorted predicted output
+$\varrho$ number of bins in $y^{t}_{n}$
+$\varphi$ percentage of stocks for each $\varrho$
+$\xi$ middle percentage of $y^{t}_{n}$
+$\mathcal{L}$ loss of holding stock
+$\mathbf{F}$ multi-head attention output
+$\mathbf{K}$, $\mathbf{Q}$, $\mathbf{V}$ key, query, value in quantformer
+$\mathbf{W}$ trainable weight matrix
+$\mathbf{Z}$ output layer